A Complete Bibliography of ACM Transactions on Asian and Low-Resource Language

Information Processing (TALLIP)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 October 2025 Version 1.45

Title word cross-reference

3 [823, 810, 819]. 3 + 2 [470]. ^{2} [612]. K [30]. μ [172]. N [506, 713, 24, 26]. S^{3} [195].

-Forcing [172]. -Gram [26, 506, 713, 24]. -Means [30]. -NET [195].

15919 [188]. 19 [956]. 1DCNN [947]. 1DCNN-BiLSTM [947].

2 [708]. 2.0 [901]. 2.0-Based [901].

3 [317]. 3.0 [471].

5G [488].

65 [645].

Abdel [507]. Abilities [293]. Ability [476, 318]. Abstract [768, 764]. Abstract-Generation [768]. Abstractive [1012, 782, 889, 527, 772, 891, 143]. Abusive [884, 754]. Accelerating [720]. Accent [1001]. Accented [6]. Acceptability [991]. According [246]. Accounting [471]. Accounts [921]. Accuracy [973, 813]. Accurate [614, 1005, 631]. Accurately [240]. Acoustic [773, 446, 22, 335, 699]. Acquisition [953]. Across [1013, 771]. Act [95, 18]. Action [26]. Activation [799]. Active [400, 838, 420, 590]. Activity [581]. Actor [319]. Acyclic [934]. Adaptation [807, 406, 754, 77, 310, 718, 32]. Adapter [691]. Adapter-based [691]. Adaptive [994, 454, 242, 725, 876, 887, 543, 1015, 591, 805]. Additional [144]. Addressing [959, 982]. Advanced [851, 215]. Advancements [982]. Advances [139, 985, 530]. Advancing [436]. AdvAudio [977]. Adverbial [331]. Adversarial [419, 624, 893, 177, 900, 697, 595, 462, 508, 213, 660, 382, 543]. Advertising [730, 841, 810, 819]. AER [412]. Affecting [292]. Affective [953]. African [551, 754]. Age [859]. Age-Diverse [859]. Agency [859]. Agents [850, 157]. Agglutinative [808, 137, 902]. Aggregated [715]. Aggregation [995, 410, 705, 547]. Aggregation-based [547]. Agnostic [1001]. AI [913, 974, 850, 612]. AI-enabled [612]. AI-Enhanced [974]. AI-Generated [913]. AIC [538]. AIC-FCM [538]. Aid [324]. AIP [479]. Air [569]. AKER [963]. Alabib [645]. Alabib-65 [645]. Albanian [444]. AlgBERT [724]. Algerian [837, 684, 724, 645]. Algorithm [538, 393, 824, 945, 415, 288, 15, 545, 464, 769, 294, 443, 815, 939, 540, 286, 337]. Algorithms [851, 937, 820, 458, 850, 800, 369, 765, 468, 259]. Aligned [967]. Alignment [691, 538, 990, 29, 319, 650, 950, 832, 772, 798, 449, 82]. All-Words [152]. Allocation [395]. Alphabet [499]. Already [832]. Am [854, 490, 741]. Ambiguity [492, 625]. Amharic [677, 425, 965]. Among [570]. Amount [296]. Amplitude [412]. Analysis [30, 445, 730, 786, 127, 667, 915, 132, 605, 921, 106, 71, 771, 103, 574, 189, 959, 744, 670, 696, 566, 610, 881, 985, 363, 628, 745, 181, 200, 165, 686, 50, 446, 776, 712, 507, 326, 523, 724, 109, 956, 151, 330, 792, 746, 321, 288, 655, 896, 379, 336, 734, 423, 477, 817, 355, 567, 895, 224, 871, 467, 462, 490, 170, 450, 384, 800, 765, 372, 356, 842, 918, 542, 721, 826, 512, 386, 304, 414, 700, 412, 737, 554, 84, 316, 563, 604, 251, 752, 719, 975, 733, 307, 685, 633, 583, 56, 142, 460, 292, 813, 925, 263, 159]. Analysis [779, 874, 805]. Analytics [603, 269, 725, 619]. Analyzer [397, 758]. Analyzing [520, 865, 236, 528, 141]. Ancient [424, 1018, 166, 242, 944]. Animate [107]. Animation [72]. Annealing [540]. Annotate [111]. Annotated [45, 724, 255, 57, 152, 642]. Annotating [270]. Annotation [978, 14, 129, 101, 499, 138, 950, 546, 388, 416, 504, 183]. Answer [672, 958, 861]. Answerability [578]. Answering [672, 323, 525, 327, 185, 922, 755, 941, 958, 204, 238, 59, 919, 549, 580, 673]. Antitheses [356]. Any [183, 220]. App [915, 324]. Appearance [246]. Application [823, 775, 4, 97, 488, 814, 811]. Applications [496, 139, 771, 168, 877, 341, 604, 325]. Applied [740]. Applying [269]. Approach [966, 442, 240, 771, 615, 189, 48, 608, 959, 620, 472, 940, 155, 367, 570, 931, 378, 289, 1003, 365, 523, 713, 433, 956, 116, 728, 330, 290, 665, 4, 162, 334, 114, 950, 36, 951, 335, 965, 407, 923, 450, 86, 57, 681, 702, 76, 542, 897, 340, 386, 412, 285, 59, 639, 888, 635, 187, 908, 82, 767, 354, 953, 17, 163, 413, 886]. Approaches [573, 634, 227, 300, 656, 569, 1013, 256, 562, 2, 868, 421]. Approximating [593]. Arab [570, 507, 447]. Arabic [648, 398, 427, 182, 30, 496, 695, 667, 672, 73, 401, 1002, 924, 963, 968, 308, 915, 231, 81, 132, 1014, 560, 442, 117, 732, 261, 509, 226, 7, 418, 1006, 139, 240, 884, 71, 727, 704, 574, 189, 594, 1005, 959, 135, 153, 985, 281, 675, 31, 499, 378, 1003, 982, 365, 713, 381, 433, 260, 431, 728, 122, 1013, 552, 217, 246, 241, 238, 192, 661, 964, 99, 671, 89, 369, 765, 372, 562, 758, 681, 59, 27, 714, 834, 668, 639, 16, 95, 262, 752, 868, 565, 70]. Arabizi [192]. AraEventCoref [968]. AraSpot [834]. Architecture [781, 574, 576, 551, 721]. Architectures [746]. Argument [726, 790]. AROMA [73]. Arousal [423]. Array [46]. Art [561, 378, 320, 469, 811, 605, 379]. Article [439, 413]. Article-aware [413]. Article-Wise [439]. Articles [793, 1006, 628, 463]. Artificial [851, 254, 688, 863, 471, 79, 469]. Aryan [377, 415]. ASER [594]. Asian [603, 324, 617, 530, 848, 542, 926, 144, 789, 469, 540, 532, 207]. ASN [612]. Aspect [127, 667, 616, 956, 895, 414, 251, 752, 719, 307, 685, 934, 159, 788, 805]. Aspect-Based [127, 667, 159, 788, 616, 956, 895, 414, 752, 719, 307, 685, 805]. Aspect-Level [251, 934]. Aspects [141]. Assamese [377, 762, 936, 628, 432, 874, 448, 702, 373, 84]. Assessing [562, 701]. Assessment [725]. Assignments [910]. Assistance [755]. Assisted [688, 173]. Assisting [586]. Assistive [261, 509, 447]. Association [743]. Asynchronous [602]. Attack [333, 301]. Attention [445, 730, 722, 574, 323, 439, 498, 894, 155, 766, 616, 659, 712, 221, 389, 723, 322, 524, 486, 782, 933, 355, 567, 224, 407, 708, 527, 798, 359, 357, 329, 508, 869, 743, 928, 429, 243, 158, 979, 259, 729, 1018, 219]. Attention-Based [359, 766, 389, 708]. Attentive [189, 855, 295]. Attitude [296]. Attractions [821]. Attributes [234, 936]. Attribution [117, 531, 402, 76, 796]. Audio [538]. Audiobook [1010]. Audios [804]. Augmentation [781, 947, 884, 1008, 646, 353, 743, 803, 385, 559, 543]. Augmented [782, 1017, 876, 521, 728]. Augmenting [858]. Authentic [684]. Authentication [282]. Authorship [560, 117, 402, 76, 196, 392, 796]. Auto [230, 404, 631]. Auto-Completion [230]. Auto-Encoder [404]. Autoencoder [239]. Autoencoders [553, 172]. Automated [393, 285, 565, 376]. Automatic [903, 398, 973, 808, 709, 837, 536, 519, 101, 677, 724, 348, 397, 29, 114, 768, 609, 140, 965, 451, 405, 357, 1001, 173, 115, 679, 977, 953, 263]. Automatically [793, 81]. Automation [561]. Autoregressive [962, 752, 592]. Auxiliary [68, 179]. Aware [894, 484, 1016, 454, 399, 494, 723, 666, 900, 782, 75, 1004, 899, 598, 696, 483, 950, 435, 537, 557, 855, 354, 413]. Awareness [971]. AyaTEC [238].

Backbone [917]. Backpropagation [242]. Backward [902, 506]. Badminton [476]. Bagging [665, 939]. Baidu [230]. Balancing [197]. Bangla [85, 48, 456, 233, 455, 548, 97, 352, 563]. Barriers [989, 796]. Base [92, 204, 681, 460, 131]. Based [130, 445, 127, 667, 401, 1002, 898, 132, 978, 704, 615, 65, 759, 620, 223, 538, 984, 520, 894, 14, 155, 634, 905, 775, 367, 331, 473, 659, 510, 374, 499, 324, 289, 533, 982, 824, 148, 698, 391, 254, 713, 381, 843, 822, 433, 422, 838, 922, 491, 974, 643, 945, 415, 370, 120, 746, 122, 88, 288, 607, 621, 322, 777, 22, 4, 80, 896, 900, 738, 995, 850, 246, 328, 319, 470, 486, 545, 817, 768, 933, 53, 355, 531, 544, 273, 951, 224, 882, 335, 871, 778, 969, 238, 438, 407, 860, 124, 291, 126, 369, 128, 539, 901, 195, 942]. Based [542, 1000, 826, 412, 466, 741, 59, 359, 584, 297, 632, 740, 49, 201, 534, 115, 852, 214, 275, 329, 104, 919, 338, 68, 343, 82, 20, 816, 56, 108, 167, 218, 250, 409, 829, 818, 815, 939, 925, 421, 743, 803, 17, 183, 159, 814, 37, 156, 779, 844, 875, 840, 118, 259, 993, 788, 872, 810, 198, 337, 178, 943, 182, 691, 730, 308, 732, 674, 146, 1006, 525, 594, 658, 603, 716, 711, 610, 247, 940, 519, 748, 766, 616, 366, 165, 686, 729, 326, 190, 956, 389, 437, 792, 290, 333, 587, 348, 941, 397, 854, 618, 453, 726, 697, 650, 157, 895]. based [555, 996, 848, 255, 464, 579, 708, 501, 649, 842, 500, 312, 504, 678, 414, 405, 669, 806, 563, 606, 274, 886, 557, 752, 719, 920, 975, 307, 685, 929, 869, 805, 998, 613, 469, 590, 163, 699, 225, 305, 547]. Baseline [594, 988]. Baselines [644, 341]. Basic [812, 87]. Basis [498]. Bayesian [539, 627]. BayesKGR [627]. BC [894]. Be [693, 880]. Behavioral [775, 992, 791]. Behaviour [888]. Benchmark [308, 594, 622, 680, 981, 988, 550, 891, 970]. Benchmarking [390, 325]. Benchmarks [968]. Bengali [425, 857, 25, 865, 363, 388, 76]. BenLem [25]. BERIS [464]. BERT [781, 1002, 1006, 857, 659, 707, 713, 738, 828, 726, 650, 596, 299, 671, 714, 1007, 338, 869]. BERT-Based [1002, 338, 1006, 650, 869]. BERT-Inspired [857]. Betalogger [333]. Better [272, 825, 592]. between [314, 517, 661, 879, 259]. Beyond [951]. Bhojpuri [341]. Bi [797, 499, 1008, 737, 489]. Bi-directional [323, 489]. Bi-Gram [499]. Bi-lingual [737]. Bi-LSTM [797, 1008]. Bias [1017]. Bidirectional [513, 744, 713, 646, 629, 409, 830, 934, 547]. Bidirectionl [787]. Big [603, 290, 253, 318, 252, 999]. Bigram [8]. Bilingual [999, 289, 63, 148, 11, 86, 277, 57, 6, 173, 1004, 24, 108, 17, 28]. BiLSTM [947, 856, 281]. BiLSTM-CNN [856]. Binarization [800]. BIO [768]. BIO-Based [768]. Biomedical [757, 873]. Biometric [775]. Bisect [30]. Black [811]. Blended [473, 474]. Blending [470]. Blessing [637]. Block [390]. Blockchain [366]. Blockchain-based [366]. Bone [343]. Book [613]. Boosted [58, 897]. Boosting [984, 365, 52]. Bootstrapping [48].

Boundary [486, 782, 94]. Boundary-Aware [782]. Brahmi [287]. Branch [753]. Bravely [578]. Breaking [989, 455]. Bridge [895, 867]. Bringing [209]. British [459]. Broadcast [75, 205]. Broadened [5]. BSL [459]. BSML [547]. Building [457, 308, 81, 709, 331, 572, 552, 423, 238, 584, 549]. Built [802]. Burmese [41, 200, 165, 299, 118]. business [292]. BYANJON [352].

C2C [292]. Calculate [236]. Can [693, 741, 880]. Canonical [106]. Capsule [631]. CAPTCHAs [301]. Caption [936, 280, 772]. Captioning [407, 579, 708, 501, 1019, 987]. Captions [316]. Carrier [180]. Cascade [855]. Cascaded [265]. Case [127, 23, 197, 55, 408, 531, 948, 232, 373, 78, 235, 390, 376, 475]. Catalan [155]. Categorial [3]. Categories [66]. Categorization [226]. Categorize [233]. Category [589]. Causal [27, 593]. Causality [157]. Cause [83]. CCG [1]. CDT [74]. CDT-Styled [74]. Cells [208]. centric [809]. CESS [233]. CG [555]. CGAN [1000]. Chain [162]. Chain-to-chain [162]. Challenges [139, 135, 1003, 982, 584]. Changes [888]. Channel [153, 567, 407, 713, 333, 933, 329, 719]. Character [48, 622, 440, 1008, 734, 29, 453, 107, 340, 436, 16, 534, 91, 1010, 279, 993]. Character-based [453]. Characteristic [881, 475]. Characteristics [762]. Characters [424, 44, 650, 242, 735, 285, 97]. Chatbot [491, 248]. Chatbots [442]. ChatGPT [671, 765]. Check [35, 36, 591, 69]. Checker [460, 37]. Checking [433, 34, 887, 707]. Chengyu [338]. Chest [864]. Chief [207]. Children [235, 190]. China [472, 531, 443]. Chinese [35, 223, 538, 83, 23, 14, 155, 331, 332, 659, 541, 148, 1018, 221, 105, 38, 454, 100, 399, 643, 494, 967, 92, 625, 474, 720, 138, 74, 162, 34, 423, 524, 12, 29, 61, 67, 477, 726, 497, 650, 835, 768, 950, 644, 757, 1017, 36, 53, 166, 355, 567, 517, 954, 558, 591, 944, 996, 312, 436, 826, 735, 938, 338, 60, 485, 91, 167, 276, 283, 556, 585, 753, 887, 953, 143, 928, 248, 429, 613, 244, 710, 222, 385, 586, 660, 590, 286, 638, 873, 37, 163, 380, 663, 840, 718, 809, 993, 69, 647]. Chinese-centric [809]. Chinese-to-Spanish [14]. Chronological [622]. Chunk [33, 274]. Chunk-based [274]. Chunker [748]. Circumstance [292]. Citation [252]. Civility [897]. CLASENTI [109]. Class [109, 531, 455, 849]. Class-Specific [109]. Classes [7]. Classical [135, 621, 517, 320]. Classics [950]. Classification [211, 935, 915, 674, 146, 884, 234, 615, 282, 716, 620, 153, 775, 616, 50, 499, 212, 44, 391, 843, 221, 190, 361, 290, 322, 191, 22, 995, 119, 933, 355, 596, 242, 546, 79, 694, 671, 518, 369, 765, 147, 539, 294, 455, 386, 368, 297, 642, 888, 390, 635, 852, 920, 351, 493, 237, 409, 589, 953, 939, 743, 803, 375, 586, 934, 578, 844, 788, 886]. Classifier [742, 291]. Classifiers [372]. Classify [98]. Classifying [741]. Clean [173]. Cleaning [615, 85]. Cleansing [802]. Cloning [804]. Close [77]. Closed [549]. Closed-Domain [549]. Cloud [337]. Cloze [663]. Cloze-style [663]. Cluster [910]. Clustering [30, 525, 391, 128, 910, 337]. Clustering-Based [391, 525]. Clusters [451]. CNN [781, 326, 713, 330, 856, 738, 871, 287, 714, 329, 975]. CNN-Based [738, 871, 326, 975]. CNN-LSTM [330, 329]. Co [234, 134, 862]. Co-occurrence [134]. Co-training [234, 862]. Coarse [487, 158]. Coarse-to-Fine [487, 158]. Code [978, 210, 774, 601, 300, 529, 516, 991, 731, 778, 948, 986, 832, 700, 141, 1007, 521]. Code-Mixed [978, 516, 991, 778, 948, 832, 700, 1007, 210, 601]. Code-switched [300]. Code-Switching [731, 986, 141]. CodeFed [731]. CodeKGC [774]. Coding [374]. Coefficients [1009]. Cognates [497]. Cognition [271]. Cognitive [877, 842, 813]. Cognitive-Inspired [877]. Coherence [849, 266]. Coherent [19, 347]. Collaboration [979]. Collaborative [688, 334, 816, 163]. Collection [238, 662]. Collective [20]. College [533, 812]. Colleges [469]. Collocation [845]. Combination [14, 590, 225]. Combinatory [3]. Combine [164]. Combining [442, 226, 890, 756, 479, 866, 255, 193, 420, 819]. CoMix [862]. Command [834]. Comment [894, 359, 409, 683]. Commentary [39, 5]. Comments [304, 375]. Commerce [841, 992]. Common [650, 953]. Communication [569, 477]. Communities [113]. Community [419]. Compact [46, 540]. Comparable [171, 23]. Comparative [1002, 727, 913, 50, 569, 964, 664, 633, 142, 218, 699]. Comparison [523, 111, 661, 64]. Completion [230, 717]. Complex [681, 312, 19]. Complexity [959, 945]. Component [608]. Components [64]. Composing [526]. Composite [217, 467]. Composition [64]. Compound [526]. Comprehending [683]. Comprehension [963, 594, 430, 517, 463, 195, 466, 511, 829, 663]. Comprehensive [898, 139, 921, 388, 320, 563]. Compressed [864]. Compression [761]. Computation [840]. Computational [603, 751, 1003, 541, 530, 597]. Computer [331, 839, 688, 602]. Computing [620, 922, 259]. Concept [374]. ConceptNet [437]. Condition [239]. Condition-Transforming [239]. Conditional [893, 22, 404, 10]. Conditions [245]. Conducting [206]. Confidence [376]. Configurational [332]. Configuring [987]. Conflation [393]. Conflict [749]. Confronting [862]. Confusion [53, 613]. Connectivity [21]. Connectors [331]. Consensus [778]. Consensus-Based [778]. Consistency [417, 880, 910, 143]. consistent [591]. Constituency [680]. Constituent [274, 449, 26]. Constituent-to-Dependency [274]. Constrained [830]. Constraint [86, 743, 17]. Constraints [484]. Construct [57]. Constructing [101, 19]. Construction [774, 472, 247, 751, 724, 822, 687, 476, 470, 906, 356, 485]. Consumer [330, 296]. Contemporary [376]. Content [695, 722, 913, 742, 370, 315, 923, 539, 70, 972]. Content-Based [370, 539]. Contents [615, 581, 330]. Context [598, 216, 483, 616, 206, 980, 399, 900, 734, 54, 1017, 450, 537, 159, 18, 673, 974]. Context-Aware [399, 900, 598, 483, 537]. Context-Dependent [216, 450]. Context-focused [616]. Context-Free [54, 18]. Contextual [947, 894, 575, 454, 364, 930]. Contextual-Aware [454]. Contextualized [932, 712, 169]. Contextually [798]. Continual [1019]. Continuous [24, 1015]. Continuous-Space [24]. Contrast [445, 773, 803]. Contrastive [624, 643, 517, 951, 785, 885, 910, 780]. Contribution [675, 89]. Controlling [1010]. Conv [303]. Conv-LSTM [303]. Conv2D [736]. Conventional [595, 364]. Conversation [442, 723, 555, 929, 830, 703, 179]. Conversational [846, 850, 428, 178]. Conversations [520, 239, 345, 705, 750]. ConVerSum [951]. Convert [509]. Converting [145, 24]. Convolution [697, 579, 805]. Convolution-based [579]. Convolutional [226, 716, 153, 440, 221, 956, 607, 458, 715, 441, 147, 295, 354, 409, 928, 872, 827]. ConvS2S [186]. Cooperative [437]. Coping [918]. Copy [389]. Copying [96]. Core [826]. Coreference [968, 1014]. Corpora [81, 171, 23, 111, 307, 68, 270, 1]. Corpus [457, 371, 135, 984, 83, 492, 684, 331, 289, 383, 990, 724, 687, 184, 859, 290, 168, 906, 438, 255, 662, 923, 550, 57, 463, 356, 152, 669, 668, 961, 642, 173, 852, 975, 68, 764, 679, 485, 813, 421, 183, 844, 662]. Corpus-Based [331, 438, 68, 669]. Correcting [38]. Correction [231, 857, 216, 120, 900, 883, 302, 906, 802, 174, 556]. Correlates [512]. Correlation [106, 879]. Correlative [725]. Counterfactual [907]. Countermeasures [471]. Counterparts [77]. Course [473, 812, 822]. Coverage [127, 1]. Covering [139]. COVID [956]. COVID-19 [956]. Creation [688, 923]. Creole [161]. CRFs [787, 118]. Criteria [460]. Critic [319]. Critical [460, 64]. CRNN [582]. Cross [691, 127, 113, 984, 581, 227, 31, 982, 383, 712, 874, 13, 419, 831, 643, 967, 482, 895, 957, 951, 866, 858, 384, 832, 721, 514, 640, 740, 1001, 310, 1004, 493, 790, 339, 452, 543, 788, 863]. Cross-Accent [1001]. Cross-Domain [543, 788, 581, 419]. Cross-Language [113, 31, 13, 967]. Cross-Language-Script [832]. Cross-Lingual [984, 227, 712, 951, 866, 858, 740, 1004, 691, 383, 831, 643, 895, 957, 384, 721, 514, 640, 310, 493, 790, 339, 452, 863]. Cross-Linguality [127]. Cross-linguistic [874]. Cross-Platform [982]. Cross-Style [482]. Crossing [796]. Crosslingual [229]. Crowdsourced [662]. Crowdsourcing [438]. Cry [741]. Crying [741]. CSTD [662]. CT [769]. Cues [174, 683]. Cultural [443, 811]. Culture [477]. Curation [263]. Curriculum [470]. Curse [455]. Cursive [655]. Curve [343]. Customer [841]. Customization [728]. Cyberbullying [581, 897, 250]. cyberhate [863].

D [612, 823, 810, 819]. DA-STCM [933]. Dakje [235]. Dashboard [921]. Data [793, 781, 947, 884, 799, 349, 603, 905, 505, 529, 843, 792, 1008, 350, 646, 846, 738, 334, 1017, 531, 951, 871, 353, 467, 93, 193, 832, 253, 700, 443, 318, 565, 252, 556, 803, 385, 468, 559, 619, 886, 490]. Data-Augmented [1017]. Data-Scarce [951]. Data-to-Text [505]. Database [622, 40, 335, 262, 833]. Databases [964]. Datalog [707]. Dataset [672, 1002, 968, 1005, 999, 45, 859, 720, 651, 645, 657, 166, 889, 988, 952, 49, 891, 919, 1004, 580, 710]. Datasets [656, 1013, 644, 986, 443, 926]. DB [964]. DB-QM [964]. DDM4TST [1011]. De-Identification [949]. De-Training [186]. Deaf [447, 459]. DEAF-BSL [459]. Decision [542, 741, 325]. Decision-making [741]. Decisions [606].

Decoder [959, 430, 407, 579]. Decoders [902]. Decoding [513, 487, 38]. Decomposition [836, 28]. Decorrelation [969]. Decorrelation-Based [969]. Deep [398, 445, 73, 231, 799, 71, 598, 189, 360, 670, 456, 881, 634, 748, 363, 745, 327, 324, 541, 712, 440, 569, 185, 433, 370, 330, 746, 290, 168, 553, 256, 600, 856, 847, 846, 983, 298, 453, 817, 169, 596, 957, 558, 595, 467, 848, 599, 317, 249, 255, 280, 291, 769, 372, 842, 721, 340, 678, 405, 412, 466, 737, 364, 396, 98, 888, 534, 251, 908, 733, 421, 225, 198, 864, 459]. Deep-learning [255]. Deep-supervised [596]. DeepSVM [741]. Defect [351]. Degraded [456, 387]. Deixis [710]. Delexicalized [164]. Delivery [667]. Denigrate [359]. Denoising [553]. Dense [333, 501]. DenseNet [534]. Dependence [293]. Dependencies [310, 161]. Dependency [278, 65, 984, 227, 145, 206, 381, 105, 406, 136, 3, 482, 726, 756, 169, 957, 126, 312, 51, 504, 414, 274, 115, 1, 218, 925, 159, 18, 156]. Dependency-Based [218, 726]. Dependent [216, 450]. Depression [733]. Derived [113]. Derogatory [360]. Description [14, 1012, 361]. Descriptive [641]. Design [824, 812, 688, 710, 469, 286, 810, 819]. Designed [421]. Detect [529, 364]. Detected [376]. Detecting [921, 570, 369, 897, 916, 248]. Detection [496, 695, 231, 978, 573, 598, 722, 615, 904, 837, 913, 444, 742, 999, 601, 894, 581, 745, 367, 982, 575, 365, 588, 433, 1016, 626, 967, 260, 431, 728, 516, 997, 425, 333, 854, 257, 358, 600, 617, 618, 856, 846, 738, 850, 983, 1017, 731, 948, 923, 992, 384, 754, 832, 631, 918, 436, 495, 514, 623, 368, 359, 640, 632, 885, 926, 508, 733, 250, 276, 420, 763, 998, 602, 878, 912, 701, 863, 827, 886]. Detector [599]. Determination [395]. Devanagari [582, 601, 15, 301]. Developed [14]. Developing [327, 383, 415, 313, 550]. Development [622, 529, 397, 657, 923, 504, 84, 488, 286, 810, 976]. Developments [267]. Device [447]. Devices [356]. Dharma [187]. Diacritic [418, 281, 122]. Diacritic-Based [122]. Diacritics [140]. Diacritization [648, 398, 378]. Diagnosis [820]. Dialect [182, 132, 985, 724, 1013, 552, 140, 192, 914, 868]. Dialect-standard [552]. Dialectal [959, 982]. Dialects [81, 222, 385]. Dialog [347]. Dialogue [572, 893, 859, 404, 535, 345, 639, 971, 602, 339]. DialoguePFM [929]. Dialogues [454, 151]. Dictionaries [70]. Dictionary [903, 11, 544, 86, 277, 376, 17]. DIEU [703]. Differences [380]. Different [727, 141, 693, 490]. Differential [725, 337]. Diffractive [995]. Diffusion [1011, 872]. Digit [657]. Digital [536, 332, 897, 443, 814, 811]. Diglossia [972]. Dimensional [423]. Dimensionality [727]. Direct [951]. Directed [934]. directional [323, 489]. Dirichlet [395]. Disabilities [725]. Disambiguate [418]. Disambiguating [246]. Disambiguation [427, 947, 890, 620, 773, 675, 21, 437, 486, 232, 152, 396, 714, 112]. Discharge [949]. Discourse [146, 88, 138, 74, 67, 411, 681, 879, 59, 98, 710]. Discourse-Based [59]. Discovery [254]. Discrepancy [867]. Discrimination [222, 385]. Discriminator [424]. Discussion [350]. Disentangled [770, 1011]. Disfluency [420]. Disorders [570]. Dissimilarity [826]. Distant [247, 258, 784]. Distillation [510, 524, 761, 543, 1015]. Distilling [902]. Distinct [314]. Distinguishing [939]. Distortions [16]. Distributed [132, 922, 576, 612]. Distribution [589, 143]. Diverse [859, 239, 938]. Diversity [849]. Do [1009]. Doc2Vec [538]. Document [937, 62, 1018, 399, 417, 4, 55, 246, 119, 470, 387, 527, 466, 285, 297, 640, 910, 876, 110, 829, 939, 571, 502, 899, 946]. Document-Level [399, 417, 119, 110, 829, 571, 502, 876, 899]. Documentation [268]. Documents [30, 132, 7, 456, 865, 233, 348, 220, 416, 196, 352]. Domain [1002, 399, 807, 406, 591, 749, 564, 346, 754, 919, 549, 110, 887, 543, 718, 788, 178, 673, 581, 419, 435]. Domain- [435]. Domain-Adaptive [887, 591]. Domain-Aware [399]. Domain-Invariant [543]. Domain-specific [110]. Domains [771, 788]. Dongba [944]. Don’t [578]. Double [46, 180]. Double-Array [46]. Double-Carrier [180]. Down [416]. Downstream [378]. Driven [804, 555, 426, 222, 785]. Drug [744]. DSC [1008]. DSISA [756]. Dual [424, 625, 404, 933, 757, 719, 753, 28, 118, 805]. Dual-Branch [753]. Dual-channel [933, 719]. Dual-Layer [118]. Dual-View [404]. Duration [528]. Dynamic [204, 579, 495, 899, 703, 892]. DZ [684]. DZ-SMS [684].

E-Commerce [841, 992]. E-mail [831]. EADRE [899]. Early [596, 888]. Early-exiting [596]. Economic [485]. ed [802, 514]. editing [357]. Editor [630, 207]. Editor-in-Chief [207]. Editorial [39, 603, 5, 597, 207]. EDU [4]. EDU-Based [4]. Education [472, 474, 476, 470, 471, 542, 468]. Effect [431]. Effective [71, 533, 365, 67, 906, 531, 124, 154, 357]. Effectively [116]. Effectiveness [817, 546, 972]. Effects [865, 296, 814]. Efficient [487, 622, 529, 120, 962, 599, 407, 631, 24, 660, 194]. EFL [331]. Egyptian [1003]. EiAP [894]. EiAP-BC [894]. Electrical [917]. Electronic [374, 283, 590]. Elements [292, 413]. ElHalim [507]. Emails [668]. Embedding [182, 781, 947, 127, 932, 153, 100, 643, 755, 846, 558, 353, 297, 632, 108, 993, 180, 827]. Embedding-Based [297]. Embeddings [648, 605, 210, 229, 66, 526, 600, 169, 490, 700, 1007, 134, 198, 976]. EmoBank [423]. Emoji [566, 894, 600, 224]. Emoji-Based [224]. Emojis [879]. Emotion [598, 696, 610, 83, 212, 575, 516, 723, 458, 850, 428, 871, 464, 914, 653, 929, 767, 840, 703, 705, 750]. Emotion-aware [696]. Emotional [730, 893, 404, 1000, 830, 179]. EmotionNet [212]. Emotions [290]. Empathy [916]. Emphasis [434]. Empirical [472, 980, 372, 861, 258, 91]. Empowering [897]. enabled [612]. Encoder [781, 947, 959, 744, 893, 646, 404, 715, 407, 579, 631]. Encoder-CNN [781]. Encoder-Decoder [959, 407, 579]. Encoders [757]. Encoding [729, 896, 403, 410, 244]. Encyclopedia [275, 919]. End [808, 736, 96, 74, 878, 222]. End-to-End [808, 96, 74, 878, 222, 736]. Endangered [42]. Endeavour [880]. ended [962]. Ending [264]. Engine [442, 346]. Engineering [917, 468]. English [978, 240, 999, 601, 33, 529, 289, 533, 383, 980, 185, 687, 516, 197, 306, 676, 67, 948, 384, 57, 700, 141, 668, 783, 640, 740, 816, 142, 293, 814, 559, 325]. English-to-Mizo [306]. English-Urdu [384]. English/Arabic [668].

Enhance [400, 857, 681, 468, 452]. Enhanced [973, 959, 982, 990, 389, 974, 967, 676, 205, 539, 798, 719, 869, 1010, 638, 180, 626, 252, 158]. Enhancement [566, 686, 324, 651, 441, 438]. Enhancing [849, 417, 782, 835, 883, 848, 527, 783, 641, 26]. Enlightenment [475]. EnML [694]. Enough [483]. Enriched [71, 989, 801, 777, 54, 963]. Enriching [781, 364]. Enrichment [1017]. Ensemble [771, 959, 742, 665, 618, 749, 694, 671, 769, 737, 364, 920, 939]. Ensembles [897]. Ensembling [702]. Entities [102, 248, 590, 899]. Entity [966, 211, 498, 58, 66, 503, 990, 729, 838, 801, 494, 625, 168, 777, 651, 646, 111, 782, 757, 219, 479, 954, 558, 785, 79, 57, 312, 403, 187, 110, 283, 585, 660, 159, 873, 452]. Entity-Augmented [782]. Entity-Based [777]. Entity-driven [785]. Entrepreneurial [473, 817]. Entropy [918, 840]. Environment [331, 987, 337]. Epidemic [779, 474]. Epidemic-related [779]. Epileptic [745]. Equipment [351]. Equivalence [7]. Equivalents [951]. Era [474, 207]. ERC [705]. Error [231, 931, 900, 883, 302, 906, 591, 276, 556]. Error-consistent [591]. Errors [865, 38, 649]. Establishment [92, 292]. Estimates [341]. Estimation [793, 80, 160, 197]. Estimator [80, 465]. Etext [270]. Euphemisms [481]. Evaluating [981, 70, 854]. Evaluation [1002, 472, 581, 656, 838, 474, 476, 256, 29, 114, 531, 372, 952, 908, 213, 156]. Event [968, 659, 713, 967, 61, 726, 835, 436, 514, 876, 485, 763, 790, 429, 375, 710, 502, 899, 522]. Event-type [899]. Events [953]. Everyday [520]. Evolution [616, 337]. Evolutionary [815]. Evolutive [826]. eWOM [296]. Exhaustive [654]. Existing [460]. exiting [596]. Expanding [903, 90]. Expansion [793, 62, 373, 309, 115, 998, 673]. Experience [157]. Experience-based [157]. Experiment [70]. Experimental [176, 296, 992]. Experiments [988, 664]. Expert [766]. Explainable [712, 535, 979]. Explanation [683, 761]. Explicitly [175]. Exploitation [469]. Exploiting [90, 497, 681, 452]. Exploration [851, 357]. Exploratory [311]. Exploring [936, 462, 879, 740, 342, 868, 91, 972]. Expression [720, 2, 879]. Expression-Level [720]. Expressions [520]. Extended [898, 587, 986]. Extensible [398, 237]. Extensive [117]. Extra [33]. Extra-Chunk [33]. Extracted [499]. Extracting [171, 217, 27, 183]. Extraction [672, 1002, 7, 1006, 594, 223, 23, 905, 659, 990, 656, 92, 458, 15, 61, 328, 726, 835, 11, 219, 567, 124, 681, 528, 312, 752, 876, 20, 829, 790, 429, 571, 208, 502, 413, 899, 1015, 819]. Extraction-based [594]. Extractive [689, 75, 205, 291, 142, 354]. Extractor [454]. eXtreme [365]. Extremely [940]. Extremism [615].

Facebook [132, 311, 581]. Fact [707, 433]. Fact-checking [707]. Factoid [389]. Factorized [722]. Factors [223]. Faithfulness [782]. Fake [444, 367, 366, 433, 361, 358, 992, 631, 364, 368, 632, 885, 912]. fall [886]. Families [86, 277]. Family [510]. Farsi [52]. Fast [614, 797, 596, 46]. FastSpeech2 [1000]. FastText [713, 250]. Fault [905]. FCM [538]. Feasibility [396]. Feasible [129]. Feature [123, 670, 658, 566, 281, 766, 686, 990, 151, 458, 15, 524, 328, 567, 527, 294, 495, 632, 880, 910, 752, 878, 543, 705, 819]. Feature-Attention [524]. Feature-based [658]. Feature-rich [281]. Features [921, 48, 773, 148, 44, 440, 190, 22, 242, 992, 528, 395, 97, 47, 993]. Federated [954, 731]. FedINER [954]. Feed [499, 256, 554]. Feedback [194]. Few [996, 763, 627, 1015, 547]. Few-Shot [627, 1015, 996, 763, 547]. Field [917]. Fields [22, 10]. Fields-Based [22]. Filling [998]. Filter [960]. Filtered [184]. Filtering [523, 587, 334, 539, 554, 163]. Filtering-based [163]. Financial [720, 664, 606]. Find [478]. Finding [272]. Fine [487, 1013, 1011, 633, 718, 788, 158]. Fine-Grained [1013, 718, 1011]. Fine-Tuning [788, 633]. Fingerspelled [72]. Finite [868]. Finite-State [868]. Fintech [720]. Five [999]. Flexible [129]. Fluent [215]. Focal [711]. Focus [793, 5]. focused [616]. Folk [815]. Following [271]. Food [667]. Fooling [977]. Forcing [172]. Foreign [791]. Form [209, 85]. Formatted [938]. Formerly [897]. Forms [266]. Forward [499, 506]. Forward-backward [506]. Four [45, 490]. Four-Tier [45]. Fragment [23]. Frame [557]. Frame-semantic [557]. Framework [62, 614, 360, 989, 35, 601, 628, 366, 831, 185, 109, 459, 524, 302, 906, 75, 280, 579, 708, 501, 291, 196, 975, 460, 237, 887, 971, 998]. Free [85, 54, 18]. Free-form [85]. Frequency [412, 686]. Function [799, 498]. Functional [520, 681]. Fusing [885]. Fusion [566, 628, 990, 698, 1018, 621, 328, 883, 558, 882, 193, 769, 412, 329, 929, 276, 753, 878, 663, 619, 912, 805]. Future [985]. Fuzzy [445, 620, 391, 21, 437].

GA [521]. GA-SCS [521]. GAGPT [708]. GAGPT-2 [708]. Game [437]. GAN [424, 595]. Gated [430, 355, 409]. Gaussian [440, 577]. GCNIIT [818]. Geéz [931]. Gender [311, 775]. General [591, 795]. General-purpose [795]. Generalities [90]. Generalizable [1005]. Generalization [318]. Generalized [86, 47]. Generated [913, 330]. Generating [389, 53, 239]. Generation [849, 793, 924, 865, 505, 1012, 994, 484, 831, 422, 607, 666, 168, 404, 535, 768, 172, 770, 428, 555, 280, 988, 564, 861, 639, 6, 201, 938, 214, 908, 565, 971, 909, 347, 830, 638, 134, 179, 178, 943, 864, 791]. Generative [774, 525, 893, 697]. Generator [958]. Genesis [161]. Genetic [540]. Genitive [217]. geoGAT [322]. Geographic [322]. Geometric [708]. Gestures [693]. Ghost [1018]. Ginni [758]. GLCM [242]. Global [483, 625, 244]. Gloss [947, 757, 869]. Glossed [268]. Glove [509]. Glyphs [479]. Gnani [220]. Goals [19]. Google [802]. Gossips [683]. GPT [966, 708]. GPT-2 [708]. Gradient [365, 465, 749]. Grading [235]. Grained [1013, 718, 1011]. Gram [499, 26, 506, 713, 754, 24]. Grammar [704, 931, 217, 3, 302, 681, 214, 460]. Grammars [54, 18]. Grammatical [381, 900, 883, 906, 556]. Granularity [909, 502, 760, 663]. Graph [171, 774, 904, 890, 223, 932, 519, 327, 289, 994, 922, 956, 967, 21, 723, 322, 348, 900, 950, 611, 275, 910, 970, 354, 493, 108, 252, 928, 429, 375, 347, 934, 578, 780, 521, 627, 892, 872, 522, 912, 305, 943, 805, 707]. Graph-Augmented [521]. Graph-Based [223, 900, 108, 519, 348, 305]. Graphical [293]. Graphs [410, 18]. Great [269]. Greyscale [301]. Ground [352]. Grounded [858]. Grouping [510]. GS2F [912]. Guided [353, 909, 761, 892, 912, 191]. Guideline [200, 388]. Gujarati [759, 266, 44, 405]. Gurmukhi [506, 49, 133]. Gurumukhi [665]. Gwab [672].

Hadith [282, 654]. Hafez [507]. hand [693]. Handling [931, 268, 124, 450]. Handwriting [123, 848, 16]. Handwritten [401, 85, 48, 797, 45, 775, 622, 440, 1008, 665, 734, 242, 387, 147, 416, 285, 97, 352, 16, 49, 133, 390, 534, 8]. Happiness [236, 253]. Harnessing [918]. Hash [945]. Hashing [847]. Hate [884, 601, 982, 575, 1016, 431, 257, 856, 948, 923, 926, 1007]. HateCircle [575]. Hateful [695, 843]. Head [869]. Headlines [935]. Health [445, 776, 463]. Hearing [324]. Helps [1009]. Herb [928]. Heritage [99, 811]. Hesitation [779]. Heterogeneous [956, 540]. Heuristic [289, 443, 286, 325, 739]. Hi [987]. Hi-MetaCap [987]. Hidden [394, 22]. Hiding [977]. HiEnWrite [999]. Hierarchical [890, 610, 994, 284, 98]. High [394, 720, 150, 488]. High-order [394]. High-pole [488]. High-Resource [150]. High-Tech [720]. Higher [470, 542]. Hindi [935, 65, 742, 620, 999, 601, 212, 185, 326, 21, 120, 437, 425, 587, 348, 397, 854, 55, 379, 827, 846, 734, 706, 983, 948, 280, 407, 579, 708, 501, 1019, 987, 918, 2, 141, 359, 632, 740, 885, 1007, 635, 880, 891, 142]. Hindi-English [601]. HindiPersonalityNet [846]. HindiSumm [891]. HindiWordNet [131]. Hinge [218]. Hinglish [1001]. Historical [1005, 246, 285, 287]. Hoist [905]. Holy [238, 875]. Homographs [418]. Hong [972]. hop [922, 755, 836, 919]. Hospitals [820]. Hostile [742]. HTK [84]. Human [913, 14, 116, 688, 991, 16, 886, 565, 602]. Human-Computer [688, 602]. Human-readable [565]. Humour [827]. HumourHindiNet [827]. Hurt [854]. Hybrid [966, 781, 940, 14, 330, 552, 15, 600, 856, 847, 36, 965, 949, 126, 853, 681, 842, 495, 888, 733, 460, 283, 829, 69]. Hybridization [259]. Hydrangea [488]. Hyper [582]. Hypergraph [750, 705]. Hypernymy [425]. Hyperparameter [765].

I3rab [381]. Ibn [758]. Ibn-Ginni [758]. ICON [680]. Identification [935, 560, 722, 360, 492, 677, 569, 1018, 713, 1013, 441, 969, 202, 568, 986, 949, 853, 528, 107, 448, 362, 215, 196, 392, 1007, 753, 18, 739]. Identifying [141]. Idiom [677, 644, 663]. Igbo [87, 154]. IIITH [662]. IIITH-CSTD [662]. Image [823, 45, 936, 628, 40, 843, 595, 280, 407, 579, 708, 501, 1019, 987, 769, 611, 816, 815, 909, 811, 42]. Image-Text [628]. Images [837, 1018, 546, 441, 800, 769, 772, 316]. Imbalance [455]. Imbalanced [290]. IMDP [971]. Impact [648, 519, 980, 458, 241, 583]. Impacts [841]. Imperceptibly [116]. Implementation [536, 324, 46, 87, 810]. Implicates [492]. Implications [331, 1003]. Implicit [67, 411, 98, 971]. Importance [223, 174]. Important [103]. Improve [314, 229, 289]. Improved [771, 88, 657, 306, 768, 882, 758, 1000, 395, 806, 443, 8, 143, 337]. Improvement [670]. Improves [184]. Improving [691, 127, 106, 103, 65, 210, 266, 636, 646, 3, 54, 203, 697, 609, 895, 353, 948, 568, 121, 901, 754, 294, 51, 405, 16, 449, 917, 82, 144, 493, 293, 380, 692]. In-context [974]. Inclusive [133]. Incorporating [575, 100, 119, 650, 149]. Incorporation [972]. Incremental [763]. Index [236, 37]. Indexing [536, 376]. Indian [113, 766, 747, 480, 300, 980, 190, 271, 336, 600, 981, 467, 464, 860, 853, 174, 495, 514, 78, 669, 214, 245, 68]. Indic [377, 123, 176, 636, 794, 193, 961, 390]. Indic-ST [961]. Indigenous [600, 848, 597]. Indo [377, 415]. Indonesian [572, 422, 651, 680, 302, 906, 263]. Inducing [63, 1]. Induction [544, 93, 86, 277, 68, 17]. Industrial [954]. Industry [475]. Infant [741]. Inference [206, 333, 596, 611, 295, 28]. Inflectional [266]. Influence [471, 735, 1009]. Information [427, 937, 146, 171, 113, 804, 707, 31, 454, 720, 621, 131, 470, 877, 205, 882, 193, 52, 436, 740, 537, 604, 149, 855, 342, 917, 977, 909, 469, 586, 159, 814, 791, 840, 279, 692, 32, 872, 739, 833]. Informatization [816]. Informed [904]. Inheritance [540, 811]. injected [873]. Injection [78]. Inner [993]. Inner-character [993]. Inner-word [993]. Innovation [488]. Innovations [207]. Innovative [473]. Input [116, 559]. Inscription [242, 800]. Inspired [857, 877, 259, 863]. Instance [666]. Instance-Aware [666]. Institutes [542]. Instruction [289, 962]. Intangible [811]. Integrate [116]. Integrated [808, 23, 326, 354, 493]. Integrating [94, 1, 540, 630]. Intelligence [730, 851, 254, 688, 863, 471, 469]. Intelligence-based [469]. Intelligent [851, 615, 473, 839, 812, 157, 949, 450, 554, 534, 816, 814, 811, 791]. Intent [596, 602, 345, 998]. Intention [492, 841, 296, 791]. Inter [894, 33]. Inter- [33]. Inter-Attention [894]. Interaction [824, 447, 876, 703]. Interaction-Design [824]. Interactive [513, 430, 533, 723, 625, 251]. Interest [739]. Interlinear [268]. Interlinear-Glossed [268]. Intermodal [316]. Internet [820, 318]. Interpolated [197]. Interpolation [203]. Interpretable [619]. Intra [33]. Intra- [33]. Introduction [358, 34, 530, 877, 317]. Inutile [615]. Invariant [543]. Inverted [37]. Investigating [431, 1013, 676, 396, 790]. Investigation [821, 296]. Investigational [335]. Involving [460]. IPO [472]. IR [669]. Iraqi [182, 132]. Irony [730, 904]. Irony-based [730]. Irrelevant [798]. Isarn [187]. IsiXhosa [503]. ISO [188]. Isolated [732, 773, 495]. Issue [358, 34, 530, 877, 317, 249, 597]. Issues [478]. Iterative [228].

Japanese [520, 23, 33, 151, 497, 1]. Japanese-Chinese [497]. Japanese-to-English [33]. Jewel [802]. Joined [403]. Joint [137, 129, 453, 470, 219, 299, 754, 275, 252, 998, 602]. Journal [39, 5]. Judgements [991]. Judgment [439, 413]. Junior [531]. Justifies [712].

KaEnLandetector [978]. Kannada [903, 978, 220]. KannadaLex [833]. Karaka [397]. KArSL [262]. Kathora [635]. Kazakh [786]. KDTree [945]. KenSwQuAD [580]. Kernel [106, 618, 467]. Kernel-based [618]. Key [135, 621, 943, 720]. Key-Phrase [720]. Keyphrases [499]. Keypoints [732]. Keyword [7, 1006, 223, 190, 524, 770, 613]. Keyword-based [190, 613]. Keywords [736, 178]. KG [707]. KG-BERT [707]. Khasi [421]. Khmer [350, 545]. Kit [87]. Know [832, 578]. Knowledge [963, 774, 890, 603, 932, 327, 510, 254, 922, 801, 100, 626, 967, 728, 792, 746, 92, 131, 526, 995, 883, 204, 555, 760, 681, 51, 902, 449, 275, 410, 970, 925, 761, 873, 452, 641, 627, 892, 1015, 943, 780, 707]. Knowledge-augmented [728]. Knowledge-base [131]. Knowledge-based [792]. Knowledge-enhanced [626]. Knowledge-Enriched [801, 963]. Knowledge-injected [873]. Known [897]. Komala [635]. Kong [972]. Konkani [519, 446, 438]. Korean [492, 284, 96, 136, 191, 197, 690, 10, 126, 107, 649, 201]. Kurdish [130, 457].

Label [499, 545, 1004, 585, 998, 743, 586, 844, 862, 694]. Label-wise [585]. Labeled [793]. Labeling [482, 408, 199, 344, 629, 451, 678, 829, 28, 47]. Labelling [790]. Labels [520, 94]. LAMGCN [928]. Landmark [732]. Landscape [821]. Language [786, 695, 73, 371, 898, 732, 261, 509, 561, 978, 884, 400, 573, 762, 598, 103, 973, 722, 113, 774, 525, 135, 620, 492, 936, 634, 748, 66, 766, 331, 480, 572, 510, 839, 206, 31, 383, 393, 541, 677, 821, 729, 656, 432, 13, 506, 326, 523, 822, 105, 399, 820, 967, 260, 431, 728, 997, 120, 792, 746, 333, 587, 666, 191, 651, 645, 220, 336, 600, 515, 676, 946, 958, 298, 717, 408, 268, 75, 205, 845, 335, 848, 317, 249, 490, 965, 170, 199, 870, 384, 629, 527, 320, 451, 86, 277, 402, 46, 901, 87, 754, 434, 832]. Language [942, 918, 721, 72, 435, 295, 927, 495, 362, 735, 78, 623, 669, 359, 84, 215, 584, 642, 640, 836, 262, 604, 852, 214, 258, 8, 752, 784, 975, 733, 633, 68, 583, 549, 679, 213, 767, 789, 795, 869, 70, 24, 161, 489, 580, 222, 385, 814, 761, 791, 844, 780, 972, 32, 325, 647, 739, 532, 207, 701, 976, 342, 459]. Language-Based [331]. Language-Family-Based [510]. Language-Hindi [326]. Language-knowledge [780]. Languages [127, 314, 62, 799, 377, 808, 113, 608, 137, 176, 229, 164, 636, 794, 747, 367, 300, 980, 776, 698, 432, 190, 415, 425, 358, 738, 530, 877, 981, 551, 150, 467, 202, 860, 121, 853, 77, 174, 362, 741, 514, 737, 345, 926, 902, 597, 245, 68, 144, 1004, 540, 134, 279, 809, 207, 863]. Large [240, 922, 997, 946, 230, 680, 166, 204, 662, 550, 961, 307, 976]. Large-Scale [230, 680, 204, 961, 240, 946, 662]. Largest [255]. Late [628]. Latent [807, 395]. Lattice [38, 276]. LatVis [946]. Law [824, 822, 413]. Layer [186, 118, 807]. Layer-Wise [186]. Leader [617]. Leaf [242]. Learned [901]. Learner [331, 364]. Learning [903, 398, 730, 73, 401, 911, 231, 400, 799, 573, 709, 189, 759, 989, 670, 984, 960, 634, 748, 745, 331, 480, 473, 327, 570, 529, 541, 569, 822, 433, 422, 838, 491, 974, 102, 643, 807, 820, 120, 330, 746, 290, 288, 474, 476, 666, 168, 896, 256, 336, 600, 900, 846, 738, 983, 524, 67, 298, 319, 453, 470, 545, 576, 817, 157, 517, 531, 957, 954, 951, 785, 944, 150, 969, 467, 462, 848, 599, 317, 249, 694, 661, 280, 860, 1019, 987, 725, 313, 121, 291, 769, 518, 765, 372, 232, 681, 649, 842, 76, 918, 721, 340, 678, 405, 412, 396]. Learning [664, 916, 606, 885, 98, 303, 888, 886, 1001, 534, 449, 852, 275, 410, 908, 920, 733, 767, 460, 161, 420, 589, 939, 421, 803, 385, 660, 590, 47, 873, 468, 791, 844, 619, 993, 325, 627, 862, 198, 547, 459, 146, 255, 514, 639]. Learning-Based [759, 951, 534, 421, 748, 290, 848, 649, 842, 678]. left [693]. Legal [439, 958, 817, 981, 1004, 413]. Legitimacy [363]. Lemmatizer [25, 432]. Length [403, 273]. Less [127, 625]. Less-Resourced [127]. letter [266]. Level [960, 363, 44, 588, 399, 720, 417, 119, 486, 441, 340, 640, 251, 307, 110, 829, 930, 743, 571, 502, 279, 966, 274, 876, 934, 899]. Levels [368, 390]. Leveraging [457, 947, 973, 618, 950, 757, 944, 700, 98, 1007, 144, 237, 787]. Lexical [445, 240, 131, 417, 833]. Lexically [676]. Lexicon [793, 730, 103, 181, 63, 212, 415, 545, 263]. Lexicons [90]. LFWE [632]. Library [332, 286]. Light [418]. Limbu [927]. Limited [349]. Line [387, 390]. Linear [203, 355]. Lingual [984, 227, 712, 951, 866, 858, 740, 1004, 789, 691, 383, 831, 643, 863, 895, 957, 462, 384, 721, 514, 737, 640, 310, 493, 790, 339, 452]. Linguality [127]. Linguistic [658, 675, 1003, 190, 288, 526, 341, 481, 554, 27, 796, 632, 82, 953, 380, 619, 874]. Linguistic-Relationships-Based [82]. Linguistically [1016, 680, 426]. Linguistically-Aware [1016]. Linguistically-Motivated [680]. Linguistics [603, 541, 530, 597]. Link [240, 932, 373]. Linking [757]. Lipi [220]. List [37]. ListMLE [114]. Lists [669]. Literary [76]. Literature [655, 269, 654, 434, 235, 653, 685]. LLM [968]. LMF [70]. Loanword [202, 568, 986]. Local [473, 206, 625]. Localization [501]. Localization-based [501]. Localness [592]. Log [346]. Logic [20]. Logographic [279]. Logs [51]. Long [323, 241, 939, 244]. Looking [656]. Loss [711, 218, 590]. Loss-based [590]. Low [73, 674, 799, 377, 598, 973, 608, 525, 176, 658, 711, 229, 936, 940, 960, 66, 164, 367, 572, 736, 776, 698, 432, 44, 822, 801, 643, 945, 184, 425, 587, 358, 738, 315, 676, 125, 946, 524, 877, 858, 731, 551, 150, 353, 462, 426, 93, 965, 202, 760, 124, 870, 121, 518, 564, 451, 86, 277, 901, 77, 174, 362, 737, 623, 783, 359, 961, 461, 597, 955, 752, 784, 549, 144, 1004, 789, 580, 385, 814, 194, 134, 543, 809, 788, 547, 207, 863]. Low-Level [44]. Low-Power [524]. Low-Resource [674, 973, 608, 176, 936, 960, 66, 367, 698, 822, 801, 358, 738, 125, 877, 858, 731, 150, 353, 426, 93, 202, 124, 121, 518, 564, 451, 86, 77, 174, 359, 961, 597, 549, 144, 1004, 385, 194, 543, 809, 788, 207, 799, 598, 525, 658, 711, 940, 164, 572, 432, 643, 184, 425, 587, 315, 676, 946, 462, 760, 277, 362, 737, 783, 461, 752, 784, 580, 547]. Low-Resourced [901, 736, 776, 551, 965]. Loyalty [841]. LSTM [797, 713, 422, 491, 330, 1008, 303, 329, 276, 787]. LSTMs [126, 928]. Lue [440]. Lyrics [507, 883].

MABERT [659]. Machine [851, 401, 911, 963, 314, 146, 377, 762, 727, 759, 913, 620, 487, 483, 940, 14, 155, 430, 960, 636, 794, 745, 480, 33, 374, 931, 529, 148, 698, 13, 175, 687, 116, 624, 807, 820, 184, 177, 417, 552, 80, 690, 306, 676, 125, 54, 114, 203, 272, 756, 517, 845, 944, 778, 749, 426, 93, 228, 661, 760, 124, 765, 463, 681, 195, 867, 466, 511, 783, 584, 606, 461, 593, 902, 682, 173, 449, 852, 258, 784, 917, 60, 144, 460, 24, 108, 829, 630, 158, 468, 559, 156, 194, 186, 380, 844, 209, 9, 279, 809, 892, 225, 532]. Machine-assisted [173]. Machine-learning [146]. Magahi [341]. MahaEmoSen [696]. mail [831]. Maithili [341, 500]. Major [812]. Majoring [476]. Majors [476]. Making [325, 741]. MAKS [989]. Mal [889]. Malay [975]. Malayalam [1008, 941, 889, 527, 434, 528, 340, 395, 584]. Management [470, 345, 443, 286]. Managing [270]. Mandarin [198]. Mandelbrot [918]. Manipuri [773, 147]. Manual [101]. Manuscript [496]. Manuscripts [1005, 802]. Mapping [240, 974]. Maps [230]. Marathi [696, 515, 869]. Marginal [689]. Markers [675]. Market [606]. Marketing [470]. Markov [394, 20]. Markovian [56]. Mask [659]. Mask-Attention-Based [659]. Masked [836, 855]. Match [167]. Matching [400, 171, 614, 122, 195, 611, 343, 478]. Math [149, 825]. Mathematical [531]. Matrix [995]. Maximal [273]. Maximal-length [273]. Maximum [689]. Mayek [147]. Mayfly [534]. mBERT [464]. mBERT-based [464]. Me [311]. Meaning [764, 619]. Means [30]. Measure [604]. Measurement [100, 964]. Measurements [128]. Measures [519, 21]. Mechanism [574, 729, 712, 389, 322, 96, 486, 933, 219, 355, 329, 540, 243, 37, 194, 259, 413]. Mechanisms [658, 155, 928]. Media [695, 762, 360, 366, 841, 570, 529, 776, 236, 843, 838, 738, 1017, 770, 568, 986, 386, 888, 733, 143, 263, 811, 209, 739, 856]. Media/Messaging [366]. Medical [454, 820, 785, 949, 283, 590]. Medicine [970, 928]. Melody [607]. Meme [683]. Memes [843]. Memorizing [411]. Memory [323, 255, 251, 159, 208]. Memory-based [255]. Mental [445, 776, 478]. Mention [983]. Message [115]. Message-Response [115]. Messages [236]. Messaging [366]. Meta [711, 987, 514, 639, 1001, 739, 514]. Meta-ED [514]. Meta-Heuristic [739]. Meta-Learning [987, 1001, 514, 639]. MetaCap [987]. Metadial [639]. Metaphor [840]. Metaverse [818]. Method [582, 146, 171, 536, 751, 529, 824, 365, 116, 474, 657, 933, 53, 479, 531, 845, 770, 996, 387, 867, 606, 98, 682, 173, 318, 1004, 818, 815, 977, 613, 779, 863]. Methodology [289]. Methods [282, 656, 111, 255, 562, 373, 396, 642, 926, 461, 593, 213, 701]. Metric [156, 547]. Mexican [679]. MHG [705]. MHG-ERC [705]. Micro [292]. Micro-business [292]. Microblog [375, 779]. Microblogs [83]. Middle [531, 826]. MILPaC [981]. Mind [269]. Mind-section [269]. Mine [905]. Minimal [799, 202, 49]. Minimally [61]. Mining [73, 139, 654, 346, 468]. Misinformation [358]. Mising [874]. Mislabeling [862]. Mixed [978, 529, 516, 476, 991, 778, 948, 832, 700, 1007, 210, 601]. Mixing [948]. MixSong [938]. Mixture [577]. Mixup [1009]. MizBERT [828]. Mizo [512, 973, 687, 828, 306, 676]. MMHFND [885]. Mobile [820, 325]. MobileNet [287]. Modal [624, 817, 798, 736, 546, 842, 840]. Modalities [885]. Modality [775]. Mode [812, 822]. Model [398, 73, 732, 400, 189, 759, 323, 774, 525, 265, 349, 266, 472, 247, 894, 153, 281, 227, 616, 839, 499, 533, 677, 506, 838, 394, 399, 728, 370, 88, 625, 476, 284, 322, 1008, 552, 465, 600, 618, 738, 995, 828, 946, 12, 319, 717, 817, 768, 408, 219, 355, 845, 1011, 866, 802, 555, 299, 923, 671, 527, 46, 542, 295, 511, 741, 94, 297, 6, 303, 888, 1001, 534, 115, 557, 275, 329, 338, 318, 577, 876, 733, 351, 929, 56, 795, 64, 292, 283, 293, 252, 489, 977, 909, 743, 602, 803, 710, 761, 156, 259, 69, 647, 872, 26]. Model-enhanced [252]. Modeling [65, 440, 175, 333, 29, 67, 477, 75, 205, 241, 448, 16, 752, 767, 1010, 699, 875, 198]. Modelling [553, 493, 252]. Models [672, 1002, 1006, 139, 884, 71, 594, 572, 982, 698, 523, 997, 621, 854, 991, 676, 958, 11, 882, 944, 871, 490, 661, 870, 629, 320, 416, 500, 714, 836, 8, 633, 583, 24, 556, 972, 32, 179, 305]. Modern [182, 418, 166, 517, 868]. Modified [824, 617, 689, 242, 291]. Mongolian [349, 1000, 699, 452]. Monkey [617]. Monochrome [301]. Monolingual [29, 93, 384]. MOOC [925]. Mood [879]. Morpheme [698, 10, 201, 452]. Morpheme-Based [698]. Morphemes [126]. Morphological [176, 229, 165, 99, 758, 232, 52, 340, 902]. Morphologically [71, 245, 559]. Morphology [78, 56]. Motion [824]. Motivated [680, 593, 559]. Movie [334]. MPCNN [351]. MRI [769]. MRMI [804]. MRMI-TTS [804]. MSA [652]. Mulan [439]. Multi [966, 849, 127, 804, 153, 570, 499, 736, 990, 391, 1018, 713, 822, 922, 624, 807, 755, 4, 160, 856, 847, 946, 119, 697, 717, 817, 355, 546, 150, 462, 749, 426, 694, 760, 860, 769, 649, 842, 721, 435, 466, 798, 345, 642, 274, 836, 534, 149, 508, 919, 789, 869, 485, 585, 815, 998, 909, 743, 660, 159, 502, 844, 663, 840, 325, 705]. Multi-attention [1018]. Multi-Channel [153, 713]. Multi-class [849]. Multi-Classification [355]. Multi-Contextually [798]. Multi-Document [4, 466, 946]. Multi-Domain [807, 749, 435]. Multi-Entity [159]. Multi-Feature [990]. Multi-Granularity [502, 760, 663]. Multi-Hashing [847]. Multi-Head [869]. Multi-hop [922, 755, 836, 919]. Multi-hypergraph [705]. Multi-intent [345, 998]. Multi-Label [499, 743, 844, 694]. Multi-Level [119, 743, 966, 274]. Multi-Lingual [789, 462]. Multi-Linguality [127]. Multi-Modal [624, 817, 798, 736, 546, 842, 840]. Multi-Mode [822]. Multi-Objective [534, 325, 815]. Multi-Reference [804]. Multi-Round [150]. Multi-scale [697]. Multi-Sentence [149]. Multi-social [856]. Multi-stage [721]. Multi-Target [642, 508]. Multi-Task [426, 391, 160, 717, 462, 649, 585, 660]. Multi-Turn [502]. Multi-Type [485]. Multi-View [570, 769]. Multiclass [885]. Multidisciplinary [45]. MultiLayer [333]. Multilevel [754]. Multilingual [457, 62, 709, 113, 658, 711, 636, 794, 367, 510, 652, 300, 792, 651, 900, 738, 650, 11, 121, 518, 128, 754, 362, 554, 961, 955, 910, 342, 522]. Multimedia [812, 877, 816, 469, 468]. Multimodal [911, 360, 696, 628, 821, 980, 843, 859, 256, 328, 923, 772, 885, 784, 975, 683, 814, 912, 305]. Multiple [771, 439, 83, 1013, 61, 150, 345, 1]. Multiple-User [83]. Multiresolution [734]. Multitask [136, 753]. Multiview [67]. Multiword [63]. Multization [798]. Music [815]. Mutual [804, 221, 98, 840, 619]. Mutual-Attention [221]. Myanmar [783, 41, 200, 165].

N [612]. Named [211, 498, 58, 66, 503, 729, 838, 801, 494, 625, 168, 651, 646, 111, 479, 954, 558, 79, 57, 403, 187, 110, 283, 585, 660, 452]. Names [665, 217]. Naming [966]. National [236]. Native [948, 215]. Natural [839, 206, 820, 792, 746, 298, 317, 249, 295, 584, 852, 733, 565, 767, 789, 183, 844, 739, 207]. Naturalistic [595]. naturally [886]. Nature [259]. Nature-Inspired [259]. Nearest [945]. Negative [738, 296]. Neighbor [945]. Neighborhood [410]. Neighbors [756]. Nepali [481, 504, 678]. NER [781, 247, 996, 121, 955, 1004, 543]. Nested [403]. NET [195]. Nets [736]. Network [445, 799, 598, 722, 797, 716, 439, 610, 881, 153, 324, 289, 812, 1018, 185, 838, 956, 333, 723, 288, 607, 136, 458, 553, 706, 697, 717, 558, 242, 555, 441, 79, 291, 853, 147, 500, 295, 495, 737, 604, 251, 508, 855, 719, 354, 493, 409, 753, 429, 878, 375, 934, 382, 663, 872, 705, 750, 225, 864, 827, 805]. Network-based [697]. Networks [226, 615, 904, 603, 505, 499, 221, 893, 755, 617, 618, 246, 125, 576, 224, 595, 369, 195, 918, 359, 20, 928, 32, 198, 522]. NeuMorph [176]. Neural [226, 278, 234, 377, 598, 722, 594, 904, 797, 176, 658, 716, 487, 483, 155, 881, 153, 960, 636, 794, 505, 499, 931, 378, 324, 698, 185, 175, 221, 394, 807, 184, 333, 177, 88, 288, 607, 417, 458, 553, 552, 80, 160, 690, 465, 706, 246, 306, 676, 125, 203, 272, 576, 756, 558, 802, 242, 441, 79, 749, 299, 426, 228, 760, 344, 291, 147, 52, 500, 295, 495, 368, 783, 461, 593, 357, 682, 449, 149, 258, 577, 917, 213, 354, 493, 218, 409, 489, 375, 934, 630, 158, 559, 194, 380, 259, 279, 809, 32, 892, 750, 225, 198, 179, 864, 827]. Neurocomputer [786]. New-generation [791]. News [457, 793, 444, 363, 628, 367, 366, 254, 433, 361, 777, 256, 358, 75, 205, 463, 631, 304, 554, 364, 368, 664, 632, 885, 912]. Newspaper [935]. Newton [824]. NLP [966, 797, 472, 267, 378, 515, 470, 490, 949, 87, 542, 897, 235]. NLP-Based [542]. NMT [510, 980, 315, 497, 150, 353]. Noised [324]. Noisy [980, 862]. NomBank [482]. Non [747, 300, 215, 592, 209]. Non-Autoregressive [592]. Non-code-switched [300]. Non-Indian [747]. Non-Native [215]. Non-Standard [209]. Normalization [349, 873, 209]. North [690]. Noun [273]. Nouns [107]. NOVA [129]. Novel [1002, 442, 509, 1005, 894, 378, 523, 922, 116, 728, 625, 1008, 301, 981, 988, 992, 447, 107, 635, 908, 351, 767, 795, 647, 337]. Nucleusmax [647]. Number [72]. Nyushu [42].

Object [501, 987]. Objective [454, 534, 325, 815]. occurrence [134]. occurring [886]. OCR [456, 40, 220, 802, 964]. OCR-ed [802]. OCRing [496]. OdeBERT [596]. Odia [634, 748, 622]. Offensive [573, 260, 431, 728, 923, 362, 623]. Offline [622, 193, 753]. OK [612]. OK-ASN [612]. One [483, 596, 410]. One-Shot [410]. One-stage [596]. Online [85, 123, 670, 601, 618, 328, 296, 193, 842, 942, 97, 352, 49, 133, 8, 816]. Only [693]. Ontologies [621, 540]. Ontology [182, 614, 709, 603, 746, 606]. Ontology-Based [746, 603, 606]. OOV [12, 124]. Open [994, 92, 962, 564, 919, 270, 178, 673]. Open-Domain [564, 178, 673]. Open-ended [962]. Open-Vocabulary [994]. Opinion [73, 139, 617, 375]. Optimisation [255, 918]. Optimization [582, 617, 848, 765, 277, 1007, 534, 955, 816, 110, 815, 337]. Optimized [959, 1008, 888]. Optimizing [958, 114, 866]. Oracle [343]. Order [296, 178, 394]. Order-Sensitive [178]. Ordering [513, 33]. Orders [314]. Oriented [510, 491, 511, 602, 339]. Orthographic [1003]. OT [321]. OT-ET [321]. Other [377]. Outlier [701]. Outline [208]. Output [487, 143]. Overcome [649]. Overhead [813]. Overhead-accuracy [813]. Overseas [474]. Overview [911, 480, 633].

Page [390]. Pain [854]. Pair [894, 206, 383, 587, 676, 384, 640, 115]. Pairs [773, 321, 258, 784, 532]. Pairwise [50]. Palm [242]. PAMR [764]. Parallel [457, 371, 81, 171, 23, 63, 687, 668, 532, 184]. Parameter [582, 962]. Parameterization [435]. Parameters [553, 335]. Paraphrase [90, 588, 353]. Paraphrased [997]. Paraphrasing [308]. Parenthetical [20]. Parl [371]. Parser [278, 704, 164, 13, 105, 74, 504, 149]. Parsing [65, 984, 227, 422, 88, 406, 136, 271, 3, 54, 169, 957, 860, 126, 312, 51, 414, 274, 557, 167, 161, 293, 156, 26]. Parsing-based [312]. Part [799, 634, 129, 165, 96, 350, 453, 99, 154, 702, 500, 303, 201, 421, 317]. Part-of-Speech [634, 129, 96, 350, 453, 99, 154, 421, 799, 165, 702, 303, 201]. Particle [554, 1007]. Parts [845]. Pashto [787]. Path [332, 203, 571]. Paths [410]. Pattern [898, 472, 652, 476]. Patterns [725, 528, 27, 183]. Pause [124]. Pause-Based [124]. PBL [812]. PDF [536]. Pen [261]. People [447]. Perception [296]. Perfection [269]. Performance [850, 800, 121, 901, 500]. Perplexity [918]. Persian [103, 145, 181, 729, 656, 370, 637, 964, 199, 988, 313, 368, 919, 764]. Persian/Arabic [964]. PersianMHQA [919]. Person [360, 942]. Personal [243]. Personality [999, 570, 618, 846]. Personalized [230, 971, 979]. Perspective [139, 480, 821, 78, 813, 814, 809]. Perspectives [39, 5]. Phenomena [65]. Phishing [831, 668]. Phone [300]. Phoneme [930]. Phoneme-Level [930]. PhonemeVec [930]. Phonetic [188]. Phonetics [180]. Phrase [145, 93, 124, 94, 584, 64, 720]. Phrase-Based [584]. Phrase-Boundary [94]. Phrases [273]. Physical [476]. Physiological [745]. Picture [43]. Pipeline [771, 1005, 552]. Pivot [17]. Pivot-Based [17]. Pivoted [955]. Pivoting [155]. Place [665]. Plagiarism [997]. Plan [277]. Planning [823]. Platform [982, 810]. Platforms [366]. Plural [321]. Plus [384]. Poems [528]. Poetry [637, 635, 880, 938, 908, 638]. POI [230]. Pointer [136, 829]. Pointer-sequence [829]. Polarity [620, 181, 183]. pole [488]. Polish [321]. Polishing [644]. Polysemy [932]. Poor [490, 170, 402, 336]. POS [421, 210, 748, 284, 353, 299, 10, 527, 649, 52, 161, 421, 158, 112, 137, 265].

POS-Tagging [353]. Position [729, 896, 75, 882]. Position-Aware [75]. Positional [793, 323]. Post [360, 474, 357, 630, 478]. Post-editing [357]. Post-Editor [630]. Post-epidemic [474]. Posted [236]. Posterior [555]. Posting [894]. Posts [117]. Pottery [811]. Power [524, 351]. Practical [822, 70]. Pre [884, 400, 566, 748, 33, 553, 657, 676, 264, 717, 883, 408, 644, 944, 426, 870, 564, 867, 318, 556, 761, 780, 972, 788]. Pre-Ordering [33]. Pre-Trained [408, 884, 400, 566, 748, 553, 676, 644, 944, 870, 556, 761]. Pre-Training [426, 564, 318, 788, 657, 264, 717, 883, 867, 780, 972]. Prediction [240, 439, 932, 311, 621, 80, 197, 264, 373, 548, 888, 329, 855, 872, 522]. Predictions [487]. Predictive [991]. Predictor [80]. Preference [119]. Preferences [342]. Preliminary [512]. Preordering [13]. Preparation [352]. Preprocessing [431]. Presentation [603, 296]. Preservation [315]. Pretrained [866, 633, 795]. Pretraining [839, 591]. Prevention [817]. Price [621, 329]. Principles [217]. Printed [582, 44, 40, 220, 387, 416]. Prior [440, 100]. Privacy [604, 613]. Private [497]. Probabilistic [704, 35, 918, 18]. Probability [974]. Problem [12]. Problems [177, 149, 825]. Process [440, 241]. Processing [973, 985, 839, 820, 792, 746, 131, 515, 298, 877, 848, 317, 249, 174, 584, 597, 852, 733, 767, 789, 469, 814, 844, 42, 718, 207, 701]. Product [234, 670, 842, 819]. Progression [858]. Progression-Grounded [858]. Progressive [857, 543]. Project [87]. Projection [13, 504]. Projection-based [504]. Prominent [415]. Promotion [292]. Prompt [674, 994, 801, 626, 666, 896, 996, 920, 929, 998, 873]. Prompt-based [674, 929, 998]. Prompt-tuning [626, 996]. PromptCNER [996]. Prompting [858]. Pronoun [162, 163]. Pronouns [243]. Propagation [160, 545]. PropBank [482]. Property [704]. Proposed [460]. Prose [637]. Prose2Poem [637]. Prosodic [492, 692]. Prosody [930, 1010]. Prospective [964]. Protection [443, 613]. Proximity [205]. Pruning [24]. Pseudo [184, 556]. Pseudo-parallel [184]. psycholinguistic [833]. Psychological [854, 814]. Psychology [618]. PU [545]. Publishing [536]. Punctuation [548]. Punjabi [188, 211, 506, 458, 553, 646, 460]. Purchase [841, 296]. purpose [795]. Pyramid [1018, 878].

Q [422, 319]. Q-Learning [422, 319]. QA [115]. QEST [599]. QM [964]. Qualitative [330]. Quality [472, 80, 160, 197, 296, 964, 593, 803, 1009]. Quantitative [361, 826]. Quantity [825]. Quantization [232]. Quantized [599]. Quaternion [932]. Query [62, 922, 230, 346, 309, 51, 347, 673]. Question [672, 924, 323, 525, 327, 419, 185, 389, 755, 941, 958, 204, 785, 238, 861, 59, 836, 919, 549, 580, 208, 578, 673, 943]. Question-Answering [672]. Question-Schema [578]. Question-Specific [208]. Questions [389]. Qur’an [238, 875].

R [973, 336]. Radial [498]. Radiology [647]. Random [22, 10]. Rank [689].

Ranking [847, 36, 542, 305]. ray [864]. RDF [565, 641]. Re [60, 540, 186]. Re-sample [540]. Re-Segmentation [60]. Re-Training [186]. Readability [759, 609, 562, 235]. readable [565]. Reader [963, 261, 502]. Reading [963, 594, 430, 332, 517, 715, 463, 195, 466, 511, 829, 663]. Real [261, 531, 593]. Real-time [261, 531]. Realistic [645]. Realization [551]. Reasoning [571, 627]. REBAR [465]. Recognition [903, 907, 211, 851, 582, 496, 401, 732, 973, 808, 123, 1005, 472, 498, 711, 58, 881, 66, 766, 480, 622, 839, 300, 503, 736, 729, 440, 838, 801, 494, 723, 625, 476, 1008, 168, 458, 665, 651, 646, 655, 645, 657, 734, 67, 609, 219, 479, 954, 273, 558, 428, 731, 79, 848, 464, 193, 411, 914, 901, 495, 2, 405, 287, 84, 16, 403, 6, 95, 133, 1001, 187, 534, 8, 653, 318, 929, 767, 110, 283, 585, 818, 977, 660, 811, 699, 452, 703, 705, 750, 459]. Recognize [285, 97]. Recognized [360, 693]. Recognizing [590]. Recommend [102]. Recommendation [777, 334, 338, 252, 928]. Recommendations [979]. Recommender [847]. Reconstruction [815]. Reconstructor [630]. Records [283, 590]. Recovery [281]. Rectification [188]. Recurrent [226, 722, 797, 394, 288, 246, 125, 172, 500, 409]. Recursive [73, 71]. Reduce [422, 26]. Reducing [366]. Reduction [727, 266]. Reduplication [448]. Reevaluation [8]. Reference [804, 883]. Refining [103]. Reform [473]. Region [539]. Regional [853]. Regions [496]. Regression [771]. Regularizing [143]. Reinforce [892]. Reinforce-selected [892]. Reinforced [315]. Reinforcement [491, 256]. Rejoining [343]. Related [135, 102, 779]. Relatedness [205, 128]. Relation [656, 494, 92, 755, 67, 219, 411, 518, 681, 107, 312, 557, 410, 485, 237, 829, 790, 571, 1015]. Relational [905, 729, 987, 578, 641, 872]. Relations [27, 98]. Relationship [605, 91]. Relationships [82]. Relative [123, 896]. Relevance [658, 689, 950]. Relevant [65, 772, 798]. Reliable [576]. Religiously [843]. Remote [820]. Removal [266]. Reordering [314, 191]. Reply [770]. Report [647, 864]. Representation [1003, 712, 389, 67, 1011, 969, 320, 901, 368, 303, 275, 764, 789, 64, 91, 218, 930, 899, 993]. Representations [132, 106, 744, 138, 646, 497, 576, 126, 825]. Reread [194]. Rescoring [37]. Research [135, 538, 536, 45, 473, 361, 474, 328, 471, 926, 488, 318, 343, 818, 791, 810, 819]. Residual [439, 995, 359]. Resolution [1014, 162, 710, 243, 163]. Resource [73, 674, 62, 240, 377, 973, 608, 135, 176, 229, 936, 960, 66, 367, 698, 822, 801, 336, 358, 738, 125, 877, 858, 731, 150, 353, 426, 93, 490, 170, 202, 124, 870, 121, 518, 564, 451, 86, 402, 87, 77, 174, 623, 359, 961, 597, 955, 245, 549, 144, 1004, 789, 385, 814, 194, 134, 543, 809, 788, 337, 207, 976, 799, 598, 525, 658, 711, 940, 164, 572, 432, 326, 643, 184, 425, 587, 863, 315, 676, 946, 462, 760, 277, 362, 737, 783, 461, 752, 784, 580, 547]. Resource-Poor [490, 336]. Resource-Scarce [62]. Resourced [127, 901, 736, 776, 551, 965]. Resources [139, 503, 690, 515, 423, 341, 562, 740, 144, 1]. Response [484, 535, 564, 115, 178]. Restoration [418, 140]. Restoring [424]. Results [378, 390]. Retrieval [427, 937, 113, 608, 31, 419, 945, 720, 470, 1017, 309, 740, 537, 115, 342, 613]. Retrieval-Augmented [1017]. Retrieval-Based [115]. Retrieving [772]. Retrospective [663]. Reusable [608, 238]. Reuse [384, 640]. Review [561, 282, 745, 747, 31, 656, 655, 654, 434, 653, 685]. Reviews [667, 915, 744, 670, 328, 296, 992, 369, 842, 925]. Revisiting [436]. Rewriting [883]. Rhetorical [356]. Rhyme [880]. Rhythm [652, 874]. Rich [245, 281]. Riemannian [995]. right [693]. Right-hand [693]. Risk [817]. RNN [966]. Roadmap [563]. RoBERTa [671]. Robot [818]. Robust [226, 1016, 197, 387, 649, 511, 213, 339]. Robustness [568]. Role [146, 25, 432, 569, 482, 199, 344, 629, 78, 790, 28, 47]. Roman [188, 371, 236, 1016, 516, 257, 170, 386, 916]. Romanized [1007, 762]. Round [150]. RPEPL [896]. Rubbings [343]. Rule [130, 146, 978, 620, 940, 14, 397, 255, 806, 104, 460]. Rule-Base [460]. Rule-Based [130, 978, 14, 104, 146, 940, 397, 255, 806]. Rules [308, 845]. Rumor [365, 370, 358]. Rural [823]. Russian [495].

SA [351]. SA-MPCNN [351]. SACNN [295]. Salar [335]. Saliency [416]. Same [693]. Same-right-and-different-left [693]. Sample [948, 540]. Sampling [547]. Sandhi [607]. Sanskrit [587, 271]. Sarcasm [935, 837, 600, 918]. Saudi [921]. Say [578]. Scalable [212, 196]. Scale [230, 680, 204, 961, 240, 946, 697, 662]. Scarce [62, 326, 951]. Scarce-resource [326]. SCBG [830]. Scenario [315, 948]. Scenarios [300]. Scene [994, 715, 599, 853, 611, 878]. scenes [886]. Scenic [441]. Schema [578]. Schemes [241, 250]. Scholarly [252]. School [531]. Scientific [1006]. Score [880]. Scoring [442, 984, 836]. Scratch [572, 552]. Script [582, 45, 506, 1018, 665, 15, 220, 441, 192, 853, 832, 287, 390, 42, 522]. SCS [521]. SCT [772]. SE4ExSum [354]. Seal [881]. Search [608, 102, 945, 230, 346, 942, 43, 19]. Second [824]. section [269]. Secure [945]. Security [812]. Seed [499, 53]. Seed-Based [53]. SEEUNRS [777]. Segmentation [903, 48, 797, 229, 41, 399, 12, 453, 996, 299, 99, 387, 10, 89, 187, 682, 60, 489, 118]. Segmentation-based [996]. Seizure [745]. Selected [675, 892]. Selecting [699]. Selection [458, 555, 294, 115, 134, 198, 1009]. Self [620, 729, 476, 219, 969, 564, 195, 295, 682, 869, 420, 979]. Self-Attention [869, 979, 729, 219]. Self-attentive [295]. Self-Learning [476]. Self-Matching [195]. Self-Supervised [620, 969, 564, 682, 420]. SelfSeg [682]. Semantic [786, 323, 608, 603, 716, 266, 520, 936, 374, 383, 656, 348, 397, 734, 54, 482, 717, 950, 199, 550, 344, 629, 681, 107, 611, 395, 642, 115, 149, 910, 719, 354, 167, 237, 293, 998, 143, 830, 28, 47, 18, 9, 478, 912, 557]. Semantic-aware [354]. Semantic-Constrained [830]. Semantically [904, 777]. Semantically-Informed [904]. Semantics [575, 625, 55, 89, 364, 98, 1004, 790]. Semantics-Aware [1004]. Sememe [895]. Semi [709, 629, 664, 590]. Semi-Automatic [709]. Semi-Supervised [629, 664, 590]. Semisupervised [56]. Sense [427, 947, 890, 620, 675, 21, 437, 395, 152, 396, 714]. Sensitive [178]. Sensor [333]. Sensor-based [333]. Sentence [400, 513, 23, 588, 643, 486, 205, 195, 149, 917, 307, 91, 259, 943]. Sentences [278, 363, 991, 532]. Sentential [997]. SentiFars [181]. Sentiment [127, 667, 132, 146, 234, 71, 103, 574, 189, 959, 744, 670, 696, 620, 566, 153, 628, 745, 181, 616, 212, 712, 391, 326, 523, 724, 109, 956, 151, 330, 896, 379, 336, 734, 995, 315, 423, 119, 545, 355, 567, 895, 224, 871, 467, 462, 170, 450, 388, 765, 372, 832, 842, 918, 721, 386, 304, 414, 700, 737, 664, 316, 563, 251, 752, 719, 975, 307, 685, 633, 813, 925, 263, 934, 159, 779, 788, 805]. Sentimental [746]. Separation [382]. Seq2Seq [349]. Seq2Set2Seq [770]. Sequence [525, 216, 505, 893, 284, 706, 319, 408, 678, 868, 692, 829]. Sequence-to-Sequence [216, 505, 893, 284, 868, 692]. Sequences [63]. Sequential [736, 213]. Serendipity [102]. series [827]. Service [612]. Set [445, 151, 49]. Sets [53, 490]. Setting [652]. Settings [564]. Severity [623]. SGATS [348]. Shahmukhi [211]. Shallow [860, 372, 94]. Share [832]. Shared [497]. Sharing [760, 270]. Shift [422, 26]. Shift-Reduce [422]. Shifting [183]. Short [323, 223, 63, 588, 221, 203, 933, 355, 241, 128, 386, 239, 939, 586]. Short-Path [203]. Short-Term [323]. Shortcut [719]. Shot [804, 410, 627, 1015, 691, 336, 996, 763, 790, 547]. Siamese [947, 555]. SIAT [876]. Side [333, 314]. Side-channel [333]. Sign [732, 509, 561, 766, 191, 645, 459, 72, 495, 735, 262, 214, 679, 869]. Signal [607, 174]. Signals [693, 745]. Signs [72]. Similar [921, 258]. Similarity [519, 383, 100, 341, 550, 250, 259]. Simple [398, 906]. Simplification [952]. Sina [610]. Sindhi [40]. Singer [507]. Single [266]. Single-letter [266]. Singleton [983]. Singlish [161]. Singular [321]. Sinhala [72, 304, 952, 796]. SiTSE [952]. Skeleton [237]. Skip [754]. Skip-gram [754]. Sliding [167]. Slogan [849]. Slogans [356]. Slot [998]. Smartphone [333]. SMDDH [983]. SMS [684]. SMT [78, 32]. So2al [672]. So2al-wa-Gwab [672]. Social [695, 605, 762, 615, 360, 603, 366, 841, 570, 529, 776, 236, 843, 838, 617, 618, 738, 770, 568, 986, 369, 447, 918, 386, 888, 604, 733, 818, 143, 263, 209, 619, 739, 856, 889]. Social-sum-Mal [889]. Softmax [647]. Solution [951]. Solve [334]. Solving [12, 825]. Sorting [415]. Sound [815]. Sounds [479]. Source [314, 521]. Source-side [314]. South [690, 754, 926]. Space [106, 24, 9]. Spaces [270]. Spacing [649]. Spam [894, 369, 539]. Span [594, 482, 557]. Span-based [557]. Spanish [14]. Sparse [440, 647]. Sparsity [334]. Spatial [766, 495]. Spatiality [876]. Spatiality-Augmented [876]. Spatio [193]. Spatio-temporal [193]. Speaker [569, 723, 547]. Speaker-Aware [723]. Speakers [859]. Special [358, 34, 530, 877, 317, 249, 597]. Specific [1002, 109, 807, 208, 946, 110]. Speech [903, 81, 509, 973, 808, 711, 601, 634, 129, 686, 652, 324, 982, 736, 575, 1016, 945, 431, 96, 350, 458, 553, 257, 657, 856, 453, 609, 75, 845, 866, 335, 871, 731, 464, 662, 99, 914, 154, 434, 500, 1000, 405, 84, 961, 926, 6, 95, 1007, 1001, 653, 977, 930, 421, 1010, 222, 18, 699, 692, 198, 884, 799, 165, 702, 303, 201]. Speech-Driven [222]. Speech-to-Text [961]. Spell [34]. Spelling [857, 216, 35, 38, 120, 36, 802, 591, 276, 887, 37, 69]. Spider [617]. SPK [555]. SPK-CG [555]. Spoken [492, 865, 480, 105, 657, 435, 834]. Spontaneous [105, 914]. Sport [475]. Sports [475]. Spotting [524, 416, 834]. Spread [366]. SQL [578]. SRU [195]. SRU-Based [195]. ST [961]. Stable [957]. Stack [160]. Stacked [947, 1008, 553, 888]. Stacking [857, 714]. Stage [264, 338, 535, 596, 770, 294, 721]. Stance [626, 508]. Stances [738]. Standard [182, 868, 209, 552]. Standardized [70]. Started [564]. State [39, 605, 561, 378, 379, 320, 868]. State-of-the-Art [378, 320, 605, 379]. Statements [907]. Static [846, 495]. Stationary [401]. Statistical [472, 940, 14, 634, 33, 289, 148, 13, 54, 408, 93, 228, 584, 880, 187, 144, 24, 108, 701]. Statistics [341]. STCM [933]. Stemmer [898, 806, 104]. Stemmers [747]. Stemming [442, 937, 137, 241, 89, 927]. STEMUR [393]. Step [88]. Stickiness [332]. Stock [621, 606, 329]. Stopword [669]. Story [190, 264, 909]. Story-Ending [264]. Storytelling [962]. Strategies [662, 357, 862]. Strategy [749, 345, 8, 788, 337, 1009]. Streams [738]. Stress [512]. Strictly [938]. String [167]. Stroke [123, 881, 44]. Strokes [49]. Structural [48, 356, 97]. Structurally [218]. Structure [62, 890, 145, 484, 950, 604, 912]. Structure-Aware [484, 950]. Structured [298, 317, 249]. Structures [83, 835, 46]. Student [533, 791]. Students [476, 531, 725]. Study [127, 117, 727, 913, 176, 472, 23, 311, 331, 980, 569, 425, 476, 197, 55, 408, 531, 546, 948, 992, 232, 253, 861, 373, 500, 78, 285, 669, 783, 664, 133, 390, 258, 633, 376, 64, 699, 475]. Studying [476]. Style [482, 1011, 382, 663]. Styled [74]. Stylistic [826]. Stylistics [880]. Stylometric [921, 50, 196]. StyloThai [196]. Sub [123, 807, 682, 279]. Sub-character [279]. Sub-layer [807]. Sub-Stroke-Wise [123]. Sub-word [682]. Subtask [19]. Subword [903, 189]. Subwords [272]. Suggested [43]. sum [889]. Summaries [55, 949]. Summarization [519, 1012, 990, 348, 4, 706, 689, 946, 782, 75, 205, 951, 785, 889, 858, 965, 870, 291, 527, 798, 891, 142, 354, 143, 244, 521, 647, 305]. Summary [865, 772]. Sundanese [806, 104]. Supertags [3]. Supervised [573, 48, 620, 120, 61, 544, 969, 313, 629, 564, 76, 664, 606, 803, 619, 112, 179, 596, 682, 420, 590]. Supervision [247, 883, 202]. Supplementary [151]. Support [830]. Supporting [751]. Supports [478]. Supposititious [918]. Surface [551]. Survey [1014, 560, 937, 139, 985, 227, 622, 260, 138, 379, 515, 320, 926, 604, 809]. Sustainable [268]. SUSTEM [806]. SVM [467]. Swahili [657, 303, 342, 580]. Swarm [1007]. switched [300]. Switching [731, 986, 141]. Syllable [165, 486, 299, 528, 700]. Syllable-based [165]. Syllables [126, 303]. Sync [683]. Synchronous [54]. Synergistic [713]. Synonymy [373]. Synsets [376]. Syntactic [13, 94, 449, 719, 237, 718]. Syntactically [65, 191]. Syntactically-guided [191]. Syntax [148, 273, 167]. Syntax-Based [148]. Syntaxes [790]. Synthesis [936, 866, 595, 434, 340, 1000, 930, 1010, 692, 198]. System [130, 786, 139, 323, 14, 775, 636, 766, 233, 129, 327, 374, 391, 1016, 922, 974, 893, 820, 120, 476, 941, 777, 256, 847, 334, 470, 535, 204, 845, 79, 199, 412, 741, 584, 352, 537, 852, 214, 549, 602, 286, 468, 225]. Systematic [561, 282, 747, 655, 653, 685]. Systems [849, 937, 841, 655, 958, 228, 124, 84, 115, 789, 178].

T [818]. T-A-GCNIIT [818]. Table [93, 988]. Table-to-Text [988]. Taboos [481]. Tag [56, 158]. Tag-enhanced [158]. Tag-Transition-Based [56]. Tagger [748, 397, 154, 52, 702]. Tagging [799, 210, 137, 265, 176, 634, 129, 165, 284, 96, 350, 453, 353, 299, 99, 121, 10, 527, 649, 500, 642, 303, 201, 213, 161, 421, 787, 112]. Taguchi [582]. Tai [440]. Talents [473]. TALLIP [39, 5]. TAM [595]. Tamil [287, 903, 573, 797, 529, 242, 595, 923, 721, 340, 8]. Tamil-Brahmi [287]. Taming [270]. TAMIZHI [287, 800]. Tanzania [342]. Target [510, 13, 80, 642, 557, 508, 375]. Target-aware [557]. Target-Language [13]. Target-Oriented [510]. Task [711, 905, 491, 859, 946, 644, 426, 860, 867, 435, 602, 339, 391, 160, 717, 462, 649, 585, 660]. Task-aware [435]. Task-based [711]. Task-oriented [491, 602, 339]. Task-specific [946]. Tasks [560, 378, 490, 926, 19]. Teacher [533]. Teaching [473, 533, 812, 822, 476, 488, 816, 293, 469, 814, 810]. Tech [720]. Technique [30, 401, 146, 236, 734, 765, 772, 316, 739]. Techniques [442, 727, 14, 745, 529, 379, 515, 848, 949, 554, 95, 733, 468]. Technological [488]. Technology [823, 839, 264, 471, 253, 488, 816, 791, 810, 819]. Tells [311]. Telugu [3, 689, 490, 662, 700, 534]. Template [716]. Template-based [716]. Tempo [131]. Tempo-HindiWordNet [131]. Temporal [793, 766, 131, 715, 495, 2, 193]. Term [1002, 7, 323]. Terminologies [346]. Test [238, 235]. Text [130, 786, 226, 674, 978, 884, 857, 762, 85, 797, 890, 716, 266, 223, 538, 492, 601, 45, 894, 519, 775, 628, 505, 233, 374, 378, 990, 212, 391, 254, 523, 40, 221, 422, 431, 516, 333, 322, 348, 854, 655, 269, 995, 706, 850, 689, 983, 933, 644, 268, 172, 355, 531, 1011, 595, 889, 546, 715, 981, 599, 965, 694, 988, 654, 193, 384, 291, 128, 527, 853, 294, 434, 611, 942, 542, 309, 312, 455, 548, 386, 952, 239, 368, 316, 961, 563, 640, 390, 852, 920, 733, 351, 565, 142, 354, 493, 869, 218, 589, 143, 743, 878, 803, 244, 586, 382]. Text [578, 844, 209, 840]. Text-Based [850, 942, 563]. Text-line [390]. Text-to-Speech [434]. Text-to-SQL [578]. TextCNN [567]. Textile [974]. Texts [560, 418, 759, 837, 519, 499, 55, 241, 778, 568, 986, 562, 76, 826, 796, 939]. Textual [843, 862]. TFT [960]. TFT-TL [960]. Thai [751, 22, 4, 879, 783, 664, 196, 213]. Their [246, 341]. Them [341]. Theories [271]. Theory [381, 437, 779]. Things [614, 612, 318]. Think [625]. Three [415, 264, 661, 341]. Three-Stage [264]. Thresholding [876]. Throwing [488]. Tibetan [278, 674, 267, 896, 269, 272, 453, 486, 273, 802, 882, 255, 235, 274, 275, 920, 270, 692, 943]. Tier [45]. Time [793, 686, 683, 261, 531]. Time-frequency [686]. Time-Sync [683]. TinyCheXReport [864]. TL [960]. to/from [377]. Token [960, 494]. Token-Level [960]. Tokenization [129, 165, 350, 661, 927, 583]. Tonal [773]. Tone [607, 22, 901]. Tones [512]. Tool [247, 235]. Tools [139, 515, 678, 701]. Top [416]. Top-Down [416]. Topic [266, 465, 11, 544, 241, 467, 89, 297, 752, 855, 577, 342, 493, 252, 613, 638, 875, 32]. Topic-aware [855]. Topic-Based [544]. Topic-Enhanced [638]. Topic-language [342]. Topological [97]. Topology [182, 995]. Topology-based [182]. Tourism [327, 328]. Tourist [821]. Town [475]. Toxic [409]. TPoet [638]. Tradeoff [813]. Tradition [269]. Traditional [970, 928]. Traffic [569]. Trained [672, 408, 884, 400, 566, 748, 553, 676, 644, 944, 870, 556, 761]. Training [799, 58, 247, 960, 473, 624, 177, 465, 470, 166, 172, 858, 426, 228, 564, 318, 602, 186, 788, 234, 406, 657, 264, 717, 883, 867, 780, 972, 862]. TransBERT [264]. Transcription [865, 609, 174]. Transducers [868]. Transfer [691, 989, 960, 227, 164, 288, 651, 336, 738, 995, 1011, 428, 944, 150, 518, 832, 916, 1004, 161, 385, 382, 619]. Transfer-Based [995]. Transferring [449]. Transform [401, 164]. Transformation [308, 6, 274, 869]. Transformer [966, 947, 732, 978, 367, 982, 893, 941, 854, 336, 987, 357, 868, 869, 592, 585, 875, 872]. Transformer-Based [978, 367, 982, 875, 732, 941, 854]. Transformers [924, 744, 748, 646, 637, 564]. Transforming [239]. Transient [631]. Transition [65, 394, 88, 126, 56]. Transition-Based [65, 88, 126]. Translating [637, 77, 245]. Translation [851, 314, 377, 727, 658, 620, 487, 483, 984, 14, 155, 960, 636, 794, 33, 374, 931, 289, 541, 148, 698, 13, 175, 687, 974, 116, 624, 807, 184, 587, 177, 191, 417, 552, 80, 197, 690, 306, 676, 54, 29, 114, 203, 272, 756, 166, 845, 944, 981, 778, 749, 426, 93, 228, 760, 124, 384, 867, 94, 783, 584, 961, 740, 461, 593, 902, 682, 449, 955, 258, 784, 917, 68, 868, 60, 144, 20, 24, 108, 630, 158, 559, 156, 194, 186, 380, 9, 42, 279, 809, 892, 225, 1009, 532]. Translation-Based [984]. Translations [175, 875]. Translator [50, 679]. Translators [116]. Transliterated [349, 309]. Transliteration [188, 130, 762, 940, 506, 125, 192]. Transnational [253]. TransQAM [941]. Travel [859]. Treatment [445]. Tree [148, 422, 273, 925]. Tree-LSTM-Based [422]. Tree-to-Tree [148]. Treebank [381, 71, 200, 680, 274]. Trends [985, 621, 461]. TRGCN [872]. Tri [58, 406]. Tri-Training [58, 406]. Triangulation [995]. Trigger [835]. Triples [892]. Trust [292]. Truth [352]. TSK [391]. TTS [804]. Tuning [962, 263, 788, 626, 996, 633]. Tunisian [552, 140, 192, 661, 914, 868]. Turkish [907, 904, 744, 216, 311, 581, 505, 101, 523, 43, 583, 376, 559]. Turn [502]. Tweets [365, 326, 290, 671, 701]. Twitter [117, 921, 467, 897, 141, 879]. Two [88, 535, 770, 294, 329, 338, 307]. Two-channel [329]. Two-Stage [338, 535, 770, 294]. Two-Step [88]. Type [483, 389, 403, 485, 899].

UAV [488]. UIE [905]. UIE-Based [905]. Undergraduate [472, 473, 470]. Understand [741]. Understanding [371, 572, 666, 55, 734, 850, 435, 466, 511, 971]. Understanding-oriented [511]. Unified [12, 345, 998]. Uniform [415]. Uniformly [197]. Unify [971].

Unit [203, 409, 703, 198]. Unit-Selection-Based [198]. Units [355, 699]. Unity [810]. Universal [169, 310, 161]. Universities [473]. University [791]. Unknown [177]. Unravelling [991]. Unsupervised [730, 573, 137, 265, 675, 575, 676, 544, 228, 387, 450, 344, 51, 386, 258, 784, 489, 887, 532]. Updating [204]. upon [824, 908]. Urdu [371, 516, 989, 781, 947, 371, 898, 146, 598, 65, 45, 1012, 383, 393, 236, 588, 1016, 997, 168, 777, 257, 79, 170, 694, 450, 384, 550, 451, 402, 386, 152, 364, 396, 392, 916, 642, 640, 537, 908, 653, 875, 701, 976]. UrduAI [392]. Usage [135, 725, 826]. Use [331, 791]. used [880]. Useful [328]. User [610, 83, 330, 119, 596, 992, 613, 739]. User-based [610]. User-Generated [330]. Users [311, 570, 102]. Using [30, 793, 730, 924, 732, 605, 884, 921, 400, 598, 113, 210, 48, 349, 658, 266, 14, 634, 748, 281, 628, 766, 505, 773, 327, 707, 300, 101, 931, 529, 736, 44, 506, 365, 440, 569, 190, 893, 997, 21, 330, 746, 333, 1008, 96, 191, 458, 553, 854, 646, 217, 526, 336, 600, 846, 738, 734, 995, 706, 689, 334, 125, 958, 470, 535, 224, 242, 595, 871, 441, 150, 467, 79, 599, 93, 280, 949, 313, 992, 671, 384, 769, 126, 527, 765, 346, 147, 46, 754, 416, 631, 52, 702, 373, 542, 721, 435, 495, 412, 514, 287, 27, 359, 84]. Using [316, 916, 235, 303, 836, 187, 534, 908, 310, 917, 68, 20, 767, 276, 556, 421, 18, 279, 32, 198, 582, 401, 7, 745, 13, 437, 476, 617, 618, 827, 3, 983, 438, 870, 564, 901, 681, 414, 395, 554, 6, 888, 886, 115, 975, 733, 602, 739]. Utility [286]. Utilizing [62, 107, 912]. UTTAM [120]. Uyghur [866, 56, 243, 547].

Valence [423, 263]. Valence-Arousal [423]. Variable [807]. Variational [404, 172, 239, 577]. Variationally [440]. Variations [520, 982, 559]. Various [1013]. Vec [182]. Vector [106, 9, 232]. Vectors [113, 882]. Venue [252]. Verb [851, 675]. Verbal [271]. Verbalization [641]. VerbNet [751]. Verbs [313]. Verification [370, 6, 547]. Vernacular [747]. Versatile [220]. Verse [238]. Verse-Based [238]. versus [573]. VGGNET [316]. Via [106, 691, 58, 13, 484, 967, 676, 962, 725, 121, 754, 436, 885, 98, 593, 410, 876, 977, 928, 429, 683, 28, 522, 180, 701]. Vicious [366]. Video [1012, 683, 886]. VietMedKG [970]. Vietnamese [327, 707, 148, 643, 967, 946, 958, 408, 57, 463, 861, 942, 307, 685, 633, 60, 970, 795, 380]. View [570, 404, 769]. Views [61]. Virtual [815]. Vision [886, 795]. Vision-based [886]. Visual [821, 980, 688, 962, 969, 416, 909]. Visualizations [139]. Vocabulary [903, 994, 587, 77, 976]. Vocational [470]. Vocoding [697]. Voice [804, 607, 1010]. VoIP [81]. Vowels [446, 335]. vs [913].

wa [672]. WAD [691]. WAD-X [691]. Warm [564]. Warm-Started [564]. Wasf [182]. Wasf-Vec [182]. Wav2vec [973, 901]. Wavelet [401]. Weak [883, 539]. Weak-Region [539]. Weakly [112]. web [827, 614, 58, 247, 233, 102, 471, 612, 43, 740, 20, 810, 819]. Web-Based [740, 20]. Websites [457, 992]. WeChat [292]. Weibo [610]. Weight [756, 134]. Weighted [729, 847]. Weighting [241]. WFST [699]. WFST-based [699]. WH [612]. Wide [1]. Wide-Coverage [1]. Wikipedia [23, 66, 128, 919, 68]. Wikipedia-Based [128]. Wild [270]. Wise [123, 439, 186, 585]. with-or-without [725]. without [725, 112]. Word [648, 427, 182, 691, 947, 127, 401, 314, 106, 113, 890, 620, 229, 66, 773, 41, 206, 383, 393, 175, 38, 100, 399, 21, 437, 177, 191, 80, 526, 600, 12, 453, 650, 169, 567, 882, 441, 299, 193, 387, 550, 126, 320, 232, 649, 826, 700, 395, 152, 396, 714, 368, 632, 303, 390, 187, 825, 60, 82, 91, 108, 250, 489, 134, 118, 180, 976, 827, 682, 993]. Word-Character [91]. Word-Emoji [600]. Word-Pair-Dependency [206]. Word-vector [232]. WordNet [240, 101, 21, 437, 438, 373, 313]. Words [103, 113, 349, 266, 122, 526, 246, 450, 72, 152, 49, 133, 8]. Work [799]. Works [248]. Writeprints [392]. Writer [969, 753, 831]. Writers [215]. Writing [180]. WSD [989, 25, 432].

X [691, 864, 831, 897]. X-Phishing-Writer [831]. X-ray [864]. XLIT [867]. XLM [336]. XLM-R [336]. XLS [973]. XLS-R [973].

Yi [424]. Yoruba [901].

Zero [691, 804, 162, 336, 245, 790, 163]. Zero-Resource [245]. Zero-Shot [804, 691, 336, 790]. Zeros [1014]. Zipf [918].

References

Uematsu:2015:IMD

[1] Sumire Uematsu, Takuya Matsuzaki, Hiroki Hanaoka, Yusuke Miyao, and Hideki Mima. Integrating multiple dependency corpora for inducing wide-coverage Japanese CCG resources. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(1): 1:1–1:??, January 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ramrakhiyani:2015:ATE

[2] Nitin Ramrakhiyani and Prasenjit Majumder. Approaches to temporal expression recognition in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(1):2:1– 2:??, January 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kumari:2015:ITD

[3] B. Venkata Seshu Kumari and Ramisetty Rajeshwara Rao. Improving Telugu dependency parsing using combinatory categorial grammar supertags. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(1):3:1–3:??, January 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ketui:2015:EBA

[4] Nongnuch Ketui, Thanaruk Theeramunkong, and Chutamanee Onsuwan. An EDU-based approach for Thai multi-document summarization and its application. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(1):4:1–4:??, January 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sproat:2015:TPE

[5] Richard Sproat. TALLIP perspectives: Editorial commentary: The broadened focus of the journal. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(1):5:1–5:??, January 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Shen:2015:MGA

[6] Han ping Shen, Chung hsien Wu, and Pei shan Tsai. Model generation of accented speech using model transformation and verification for bilingual speech recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(2):6:1–6:??, March 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Awajan:2015:KEA

[7] Arafat Awajan. Keyword extraction from Arabic documents using term equivalence classes. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(2):7:1–7:??, March 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sundaram:2015:BLM

[8] Suresh Sundaram and A. G. Ramakrishnan. Bigram language models and reevaluation strategy for improved recognition of online handwritten Tamil words. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(2):8:1–8:??, March 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhang:2015:TMT

[9] Jiajun Zhang, Shujie Liu, Mu Li, Ming Zhou, and Chengqing Zong. Towards machine translation in semantic vector space. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(2):9:1–9:??, March 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Na:2015:CRF

[10] Seung-Hoon Na. Conditional random fields for Korean morpheme segmentation and POS tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(3):10:1–10:??, June 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2015:MTM

[11] Xiaodong Liu, Kevin Duh, and Yuji Matsumoto. Multilingual topic models for bilingual dictionary extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(3):11:1– 11:??, June 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2015:UMS

[12] Xiaoqing Li, Chengqing Zong, and Keh yih Su. A unified model for solving the OOV problem of Chinese word segmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(3):12:1–12:??, June 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Goto:2015:PUT

[13] Isao Goto, Masao Utiyama, Eiichiro Sumita, and Sadao Kurohashi. Preordering using a target-language parser via cross-language syntactic projection for statistical machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(3): 13:1–13:??, June 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Costa-Jussa:2016:DCS

[14] Marta R. Costa-Jussà and Jordi Centelles. Description of the Chineseto-Spanish rule-based machine translation system developed using a hybrid combination of human annotation and statistical techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(1):1:1–1:??, January 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Khanduja:2016:HFE

[15] Deepti Khanduja, Neeta Nain, and Subhash Panwar. A hybrid feature extraction algorithm for Devanagari script. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(1): 2:1–2:??, January 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Shatnawi:2016:IHA

[16] Maad Shatnawi and Sherief Abdallah. Improving handwritten Arabic character recognition by modeling human handwriting distortions. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(1):3:1–3:??, January 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wushouer:2016:CAP

[17] Mairidan Wushouer, Donghui Lin, Toru Ishida, and Katsutoshi Hirayama. A constraint approach to pivot-based bilingual dictionary induction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(1):4:1–4:??, January 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Yeh:2016:SAI

[18] Jui-Feng Yeh. Speech act identification using semantic dependency graphs with probabilistic context-free grammars. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(1):5:1–5:??, January 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2016:CCSa

[19] Ting-Xuan Wang and Wen-Hsiang Lu. Constructing complex search tasks with coherent subtask search goals. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(2): 6:1–6:??, February 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tsai:2016:CWB

[20] Richard Tzong-Han Tsai. Collective Web-based parenthetical translation extraction using Markov logic networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(2): 7:1–7:??, February 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jain:2016:FHW

[21] Amita Jain and D. K. Lobiyal. Fuzzy Hindi WordNet and word sense disambiguation using fuzzy graph connectivity measures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(2):8:1–8:??, February 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kertkeidkachorn:2016:AFH

[22] Natthawut Kertkeidkachorn, Proadpran Punyabukkana, and Atiwong Suchato. Acoustic features for hidden conditional random fields-based Thai tone classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(2):9:1–9:??, February 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Chu:2016:IPS

[23] Chenhui Chu, Toshiaki Nakazawa, and Sadao Kurohashi. Integrated parallel sentence and fragment extraction from comparable corpora: a case study on Chinese–Japanese Wikipedia. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(2): 10:1–10:??, February 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2016:CCSb

[24] Rui Wang, Masao Utiyama, Isao Goto, Eiichiro Sumita, Hai Zhao, and Bao-Liang Lu. Converting continuous-space language models into Ngram language models with efficient bilingual pruning for statistical machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):11:1–11:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Chakrabarty:2016:BBL

[25] Abhisek Chakrabarty and Utpal Garain. BenLem (a Bengali lemmatizer) and its role in WSD. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):12:1–12:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhou:2016:ESR

[26] Hao Zhou, Shujian Huang, Junsheng Zhou, Yue Zhang, Huadong Chen, Xinyu Dai, Chuan Cheng, and Jiajun Chen. Enhancing shift–reduce constituent parsing with action N-gram model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):13:1–13:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sadek:2016:EAC

[27] Jawad Sadek and Farid Meziane. Extracting Arabic causal relations using linguistic patterns. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):14:1–14:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Yang:2016:BSR

[28] Haitong Yang, Yu Zhou, and Chengqing Zong. Bilingual semantic role labeling inference via dual decomposition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3): 15:1–15:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2016:MMC

[29] Maoxi Li, Mingwen Wang, Hanxi Li, and Fan Xu. Modeling monolingual character alignment for automatic evaluation of Chinese translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):16:1–16:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Abuaiadah:2016:UBM

[30] Diab Abuaiadah. Using bisect K-means clustering technique in the analysis of Arabic documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):17:1–17:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Elayeb:2016:ACL

[31] Bilel Elayeb and Ibrahim Bounhas. Arabic cross-language information retrieval: a review. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):18:1–18:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhao:2016:ALM

[32] Yinggong Zhao, Shujian Huang, Xin-Yu Dai, and Jiajun Chen. Adaptation of language models for SMT using neural networks with topic information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):19:1–19:??, March 2016. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ding:2016:IIE

[33] Chenchen Ding, Keisuke Sakanushi, Hirona Touji, and Mikio Yamamoto. Inter-, intra-, and extra-chunk pre-ordering for statistical Japanese-to-English machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(3):20:1–20:??, March 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Lee:2015:ISI

[34] Lung-Hao Lee, Gina-Anne Levow, Shih-Hung Wu, and Chao-Lin Liu. Introduction to the special issue on Chinese spell checking. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):14:1–14:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Chen:2015:PFC

[35] Kuan-Yu Chen, Hsin-Min Wang, and Hsin-Hsi Chen. A probabilistic framework for Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):15:1– 15:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Liu:2015:HRA

[36] Xiaodong Liu, Fei Cheng, Kevin Duh, and Yuji Matsumoto. A hybrid ranking approach to Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4): 16:1–16:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Yeh:2015:CSC

[37] Jui-Feng Yeh, Wen-Yi Chen, and Mao-Chuan Su. Chinese spelling checker based on an inverted index list with a rescoring mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):17:1–17:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Hsieh:2015:CCS

[38] Yu-Ming Hsieh, Ming-Hong Bai, Shu-Ling Huang, and Keh-Jiann Chen. Correcting Chinese spelling errors with word lattice decoding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):18:1–18:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Anonymous:2015:TPE

[39] Anonymous. TALLIP perspectives: Editorial commentary: The state of the journal. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):19:1–19:??, October 2015. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Hakro:2016:PTI

[40] Dil Nawaz Hakro and Abdullah Zawawi Talib. Printed text image database for Sindhi OCR. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):21:1–21:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ding:2016:WSB

[41] Chenchen Ding, Ye Kyaw Thu, Masao Utiyama, and Eiichiro Sumita. Word segmentation for Burmese (Myanmar). ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):22:1–22:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhang:2016:ITP

[42] Tongtao Zhang, Aritra Chowdhury, Nimit Dhulekar, Jinjing Xia, Kevin Knight, Heng Ji, Bülent Yener, and Liming Zhao. From image to translation: Processing the endangered Nyushu script. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):23:1–23:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sarigil:2016:SPW

[43] Erdem Sarigil, Oguz Yilmaz, Ismail Sengor Altingovde, Rifat Ozcan, and ÖzgÜr Ulusoy. A “Suggested” picture of Web search in Turkish. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):24:1–24:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Goswami:2016:CPG

[44] Mukesh M. Goswami and Suman K. Mitra. Classification of printed Gujarati characters using low-level stroke features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):25:1–25:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Choudhary:2016:FTA

[45] Prakash Choudhary and Neeta Nain. A four-tier annotated Urdu handwritten text image dataset for multidisciplinary research on Urdu script. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):26:1–26:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Norimatsu:2016:FCL

[46] Jun-Ya Norimatsu, Makoto Yasuhara, Toru Tanaka, and Mikio Yamamoto. A fast and compact language model implementation using double-array structures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):27:1–27:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Yang:2016:LGF

[47] Haitong Yang and Chengqing Zong. Learning generalized features for semantic role labeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):28:1–28:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bhowmik:2016:BHC

[48] Tapan Kumar Bhowmik, Swapan Kumar Parui, Utpal Roy, and Lambert Schomaker. Bangla handwritten character segmentation using structural features: a supervised and bootstrapping approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 15(4):29:1–29:??, June 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Singh:2016:OHG

[49] Sukhdeep Singh, Anuj Sharma, and Indu Chhabra. Online handwritten Gurmukhi strokes dataset based on minimal set of words. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):1:1–1:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

El-Fiqi:2016:PCC

[50] Heba El-Fiqi, Eleni Petraki, and Hussein A. Abbass. Pairwise comparative classification for translator stylometric analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):2:1–2:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Qiao:2016:IUD

[51] Xiuming Qiao, Hailong Cao, and Tiejun Zhao. Improving unsupervised dependency parsing with knowledge from query logs. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 16(1):3:1–3:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Passban:2016:BNP

[52] Peyman Passban, Qun Liu, and Andy Way. Boosting neural POS tagger for Farsi using morphological information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1): 4:1–4:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2016:SBM

[53] Liangliang Liu and Cungen Cao. A seed-based method for generating Chinese confusion sets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):5:1–5:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2016:ISP

[54] Junhui Li, Muhua Zhu, Wei Lu, and Guodong Zhou. Improving semantic parsing with enriched synchronous context-free grammars in statistical machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):6:1–6:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Krishnamurthi:2016:UDS

[55] Karthik Krishnamurthi, Vijayapal Reddy Panuganti, and Vishnu Vardhan Bulusu. Understanding document semantics from summaries: a case study on Hindi texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):7:1–7:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tursun:2016:STT

[56] Eziz Tursun, Debasis Ganguly, Turghun Osman, Ya-Ting Yang, Ghalip Abdukerim, Jun-Lin Zhou, and Qun Liu. A semisupervised tagtransition-based Markovian model for Uyghur morphology analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2):8:1–8:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Nguyen:2016:ACN

[57] Long H. B. Nguyen, Dien Dinh, and Phuoc Tran. An approach to construct a named entity annotated English–Vietnamese bilingual corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2):9:1–9:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Chou:2016:BWN

[58] Chien-Lung Chou, Chia-Hui Chang, and Ya-Yun Huang. Boosted Web named entity recognition via tri-training. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2): 10:1–10:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sadek:2016:DBA

[59] Jawad Sadek and Farid Meziane. A discourse-based approach for Arabic question answering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2):11:1–11:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tran:2016:WRS

[60] Phuoc Tran, Dien Dinh, and Long H. B. Nguyen. Word re-segmentation in Chinese–Vietnamese machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2):12:1–12:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2016:MSC

[61] Peifeng Li, Guodong Zhou, and Qiaoming Zhu. Minimally supervised Chinese event extraction from multiple views. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(2):13:1–13:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Atreya:2016:QER

[62] Arjun Atreya V, Ashish Kankaria, Pushpak Bhattacharyya, and Ganesh Ramakrishnan. Query expansion in resource-scarce languages: a multilingual framework utilizing document structure. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16 (2):14:1–14:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Finch:2017:IBL

[63] Andrew Finch, Taisuke Harada, Kumiko Tanaka-Ishii, and Eiichiro Sumita. Inducing a bilingual lexicon from short parallel multiword sequences. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):15:1–15:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2017:CSC

[64] Shaonan Wang and Chengqing Zong. Comparison study on critical components in composition model for phrase representation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):16:1–16:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bhat:2017:ITB

[65] Riyaz Ahmad Bhat, Irshad Ahmad Bhat, and Dipti Misra Sharma. Improving transition-based dependency parsing of Hindi and Urdu by modeling syntactically relevant phenomena. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):17:1– 17:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Das:2017:NER

[66] Arjun Das, Debasis Ganguly, and Utpal Garain. Named entity recognition with word embeddings and Wikipedia categories for a low-resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):18:1–18:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2017:IDR

[67] Haoran Li, Jiajun Zhang, and Chengqing Zong. Implicit discourse relation recognition for English and Chinese with multiview modeling and effective representation learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):19:1–19:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tholpadi:2017:CBT

[68] Goutham Tholpadi, Chiranjib Bhattacharyya, and Shirish Shevade. Corpus-based translation induction in Indian languages using auxiliary language corpora from Wikipedia. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):20:1– 20:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhao:2017:HMC

[69] Hai Zhao, Deng Cai, Yang Xin, Yuzhu Wang, and Zhongye Jia. A hybrid model for Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):21:1– 21:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wali:2017:ECL

[70] Wafa Wali, Bilel Gargouri, and Adelmajid Ben Hamadou. Evaluating the content of LMF standardized dictionaries: a practical experiment on Arabic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):22:1–22:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Baly:2017:STM

[71] Ramy Baly, Hazem Hajj, Nizar Habash, Khaled Bashir Shaban, and Wassim El-Hajj. A sentiment Treebank and morphologically enriched recursive deep models for effective sentiment analysis in Arabic. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):23:1–23:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Punchimudiyanse:2017:AFW

[72] Malinda Punchimudiyanse and Ravinda Gayan Narendra Meegama. Animation of fingerspelled words and number signs of the Sinhala sign language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):24:1–24:??, September 2017. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Al-Sallab:2017:ARD

[73] Ahmad Al-Sallab, Ramy Baly, Hazem Hajj, Khaled Bashir Shaban, Wassim El-Hajj, and Gilbert Badaro. AROMA: a recursive deep learning model for opinion mining in Arabic as a low resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):25:1–25:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kong:2017:CSE

[74] Fang Kong and Guodong Zhou. A CDT-styled end-to-end Chinese discourse parser. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):26:1–26:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2017:PAL

[75] Shih-Hung Liu, Kuan-Yu Chen, Yu-Lun Hsieh, Berlin Chen, Hsin-Min Wang, Hsu-Chun Yen, and Wen-Lian Hsu. A position-aware language modeling framework for extractive broadcast news speech summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):27:1–27:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Phani:2017:SLA

[76] Shanta Phani, Shibamouli Lahiri, and Arindam Biswas. A supervised learning approach for authorship attribution of Bengali literary texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):28:1–28:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Passban:2017:TLR

[77] Peyman Passban, Qun Liu, and Andy Way. Translating low-resource languages by vocabulary adaptation from close counterparts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):29:1–29:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

S:2017:RMI

[78] Sreelekha S. and Pushpak Bhattacharyya. Role of morphology injection in SMT: a case study from Indian language perspective. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):1:1–1:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Malik:2017:UNE

[79] Muhammad Kamran Malik. Urdu named entity recognition and classification system using artificial neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):2:1–2:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kim:2017:PEN

[80] Hyun Kim, Hun-Young Jung, Hongseok Kwon, Jong-Hyeok Lee, and Seung-Hoon Na. Predictor–estimator: Neural quality estimation based on target word prediction for machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):3:1–3:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Almeman:2017:ABV

[81] Khalid Almeman. Automatically building VoIP speech parallel corpora for Arabic dialects. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):4:1–4:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tran:2017:LRB

[82] Phuoc Tran, Dien Dinh, Tan Le, and Long H. B. Nguyen. Linguisticrelationships-based approach for improving word alignment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):5:1–5:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Cheng:2017:ECC

[83] Xiyao Cheng, Ying Chen, Bixiao Cheng, Shoushan Li, and Guodong Zhou. An emotion cause corpus for Chinese microblogs with multipleuser structures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):6:1–6:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sarma:2017:DAS

[84] Himangshu Sarma, Navanath Saharia, and Utpal Sharma. Development and analysis of speech recognition systems for Assamese language using HTK. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):7:1–7:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bhattacharya:2017:COB

[85] Nilanjana Bhattacharya, Umapada Pal, and Partha Pratim Roy. Cleaning of online Bangla free-form handwritten text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(1):8:1–8:??, November 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Nasution:2018:GCA

[86] Arbi Haza Nasution, Yohei Murakami, and Toru Ishida. A generalized constraint approach to bilingual dictionary induction for low-resource language families. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(2):9:1–9:??, February 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Onyenwe:2018:BLR

[87] Ikechukwu E. Onyenwe, Mark Hepple, Uchechukwu Chinedu, and Ignatius Ezeani. A basic language resource kit implementation for the Igbo NLP project. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(2):10:1–10:??, February 2018. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jia:2018:IDP

[88] Yanyan Jia, Yansong Feng, Yuan Ye, Chao Lv, Chongde Shi, and Dongyan Zhao. Improved discourse parsing with two-step neural transition-based model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(2):11:1–11:??, February 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Naili:2018:CSS

[89] Marwa Naili, Anja Habacha Chaibi, and Henda Hajjami Ben Ghezala. The contribution of stemming and semantics in Arabic topic segmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(2):12:1–12:??, February 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Fujita:2018:EPL

[90] Atsushi Fujita and Pierre Isabelle. Expanding paraphrase lexicons by exploiting generalities. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(2):13:1–13:??, February 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2018:EEW

[91] Shaonan Wang, Jiajun Zhang, and Chengqing Zong. Empirical exploring word-character relationship for Chinese sentence representation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):14:1–14:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jia:2018:COR

[92] Shengbin Jia, Shijia E., Maozhen Li, and Yang Xiang. Chinese open relation extraction and knowledge base establishment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):15:1–15:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Marie:2018:PTI

[93] Benjamin Marie and Atsushi Fujita. Phrase table induction using monolingual data for low-resource statistical machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):16:1–16:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Salami:2018:ISS

[94] Shahram Salami and Mehrnoush Shamsfard. Integrating shallow syntactic labels in the phrase-boundary translation model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):17:1–17:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sherkawi:2018:ASA

[95] Lina Sherkawi, Nada Ghneim, and Oumayma Al Dakkak. Arabic speech act recognition techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):18:1–18:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jung:2018:EEK

[96] Sangkeun Jung, Changki Lee, and Hyunsun Hwang. End-to-end Korean part-of-speech tagging using copying mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):19:1–19:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sen:2018:AST

[97] Shibaprasad Sen, Ankan Bhattacharyya, Pawan Kumar Singh, Ram Sarkar, Kaushik Roy, and David Doermann. Application of structural and topological features to recognize online handwritten Bangla characters. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):20:1–20:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

She:2018:LHD

[98] Xiaohan She, Ping Jian, Pengcheng Zhang, and Heyan Huang. Leveraging hierarchical deep semantics to classify implicit discourse relations via a mutual learning method. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):21:1–21:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mohamed:2018:MSP

[99] Emad Mohamed. Morphological segmentation and part-of-speech tagging for the Arabic heritage. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):22:1–22:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Huang:2018:IPK

[100] Degen Huang, Jiahuan Pei, Cong Zhang, Kaiyu Huang, and Jianjun Ma. Incorporating prior knowledge into word embedding for Chinese word similarity measurement. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):23:1–23:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ehsani:2018:CWT

[101] Razieh Ehsani, Ercan Solak, and Olcay Taner Yildiz. Constructing a WordNet for Turkish using manual and automatic annotation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):24:1–24:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Huang:2018:LRR

[102] Jizhou Huang, Shiqiang Ding, Haifeng Wang, and Ting Liu. Learning to recommend related entities with serendipity for Web search users. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(3):25:1–25:??, May 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Basiri:2018:WII

[103] Mohammad Ehsan Basiri and Arman Kabiri. Words are important: Improving sentiment analysis in the Persian language by lexicon refining. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):26:1–26:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Suryani:2018:RBS

[104] Arie Ardiyanti Suryani, Dwi Hendratmo Widyantoro, Ayu Purwarianti, and Yayat Sudaryat. The rule-based Sundanese stemmer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):27:1–27:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

He:2018:DPS

[105] Ruifang He, Yaru Wang, Dawei Song, Peng Zhang, Yuan Jia, and Aijun Li. A dependency parser for spontaneous Chinese spoken language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):28:1–28:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bai:2018:IVS

[106] Xuefeng Bai, Hailong Cao, and Tiejun Zhao. Improving vector space word representations via kernel canonical correlation analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):29:1–29:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Park:2018:NCI

[107] Taekeun Park and Seung-Hoon Kim. Novel character identification utilizing semantic relation with animate nouns in Korean. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):30:1–30:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2018:GBB

[108] Rui Wang, Hai Zhao, Sabine Ploux, Bao-Liang Lu, Masao Utiyama, and Eiichiro Sumita. Graph-based bilingual word embedding for statistical machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):31:1–31:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Hamdi:2018:CCS

[109] Ali Hamdi, Khaled Shaban, and Anazida Zainal. CLASENTI: a classspecific sentiment analysis framework. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):32:1– 32:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2018:DSN

[110] Limin Wang, Shoushan Li, Qian Yan, and Guodong Zhou. Domainspecific named entity recognition with document-level optimization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):33:1–33:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Komiya:2018:CMA

[111] Kanako Komiya, Masaya Suzuki, Tomoya Iwakura, Minoru Sasaki, and Hiroyuki Shinnou. Comparison of methods to annotate named entity corpora. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):34:1–34:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhou:2018:WSP

[112] Deyu Zhou, Zhikai Zhang, Min-Ling Zhang, and Yulan He. Weakly supervised POS tagging without disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 17(4):35:1–35:??, August 2018. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bhattacharya:2019:UCW

[113] Paheli Bhattacharya, Pawan Goyal, and Sudeshna Sarkar. Using communities of words derived from multilingual word vectors for cross-language information retrieval in Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1): 1:1–1:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3208358.

Li:2019:OAE

[114] Maoxi Li and Mingwen Wang. Optimizing automatic evaluation of machine translation with the ListMLE approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):2:1–2:??, January 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3226045.

Su:2019:RSA

[115] Ming-Hsiang Su, Chung-Hsien Wu, Kun-Yi Huang, and Wu-Hsuan Lin. Response selection and automatic message-response expansion in retrieval-based QA systems using semantic dependency pair model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):3:1–3:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3229184.

Huang:2019:IMH

[116] Guoping Huang, Jiajun Zhang, Yu Zhou, and Chengqing Zong. Input method for human translators: a novel approach to integrate machine translation effectively and imperceptibly. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1): 4:1–4:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3230638.

Altakrori:2019:AAA

[117] Malik H. Altakrori, Farkhund Iqbal, Benjamin C. M. Fung, Steven H. H. Ding, and Abdallah Tubaishat. Arabic authorship attribution: an extensive study on Twitter posts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):5:1–5:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/ft_gateway.cfm?id=3236391.

Zhang:2019:WSB

[118] Shaoning Zhang, Cunli Mao, Zhengtao Yu, Hongbin Wang, Zhongwei Li, and Jiafu Zhang. Word segmentation for Burmese based on dual-layer CRFs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):6:1–6:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3232537.

Li:2019:IML

[119] Junjie Li, Haoran Li, Xiaomian Kang, Haitong Yang, and Chengqing Zong. Incorporating multi-level user preference into document-level sentiment classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):7:1–7:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3234512.

Jain:2019:UES

[120] Amita Jain, Minni Jain, Goonjan Jain, and Devendra K. Tayal. “UT-TAM”: an efficient spelling correction system for Hindi language based on supervised learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):8:1–8:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3264620.

Murthy:2019:INT

[121] Rudra Murthy, Mitesh M. Khapra, and Pushpak Bhattacharyya. Improving NER tagging performance in low-resource languages via multilingual learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):9:1–9:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3238797.

Jarrar:2019:DBM

[122] Mustafa Jarrar, Fadi Zaraket, Rami Asia, and Hamzeh Amayreh. Diacritic-based matching of Arabic words. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2): 10:1–10:??, February 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3242177.

Bhattacharya:2019:SSW

[123] Nilanjana Bhattacharya, Partha Pratim Roy, and Umapada Pal. Substroke-wise relative feature for online Indic handwriting recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):11:1–11:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3264735.

Mrinalini:2019:PBP

[124] K. Mrinalini, T. Nagarajan, and P. Vijayalakshmi. Pause-based phrase extraction and effective OOV handling for low-resource machine translation systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):12:1–12:??, February 2019. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3265751.

Le:2019:LRM

[125] Ngoc Tan Le, Fatiha Sadat, Lucie Menard, and Dien Dinh. Low-resource machine transliteration using recurrent neural networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):13:1–13:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3265752.

Na:2019:TBK

[126] Seung hoon Na, Jianri Li, Jong hoon Shin, and Kangil Kim. Transitionbased Korean dependency parsing using hybrid word representations of syllables and morphemes with LSTMs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):14:1– 14:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3241745.

Akhtar:2019:IWE

[127] Md Shad Akhtar, Palaash Sawant, Sukanta Sen, Asif Ekbal, and Pushpak Bhattacharyya. Improving word embedding coverage in less-resourced languages through multi-linguality and cross-linguality: a case study with aspect-based sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):15:1– 15:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3273931.

Nakamura:2019:WBR

[128] Tatsuya Nakamura, Masumi Shirakawa, Takahiro Hara, and Shojiro Nishio. Wikipedia-based relatedness measurements for multilingual short text clustering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):16:1–16:??, February 2019. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3276473.

Ding:2019:NFF

[129] Chenchen Ding, Masao Utiyama, and Eiichiro Sumita. NOVA: a feasible and flexible annotation system for joint tokenization and part-of-speech tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):17:1–17:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3276773.

Ahmadi:2019:RBK

[130] Sina Ahmadi. A rule-based Kurdish text transliteration system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):18:1–18:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3278623.

Kamila:2019:THL

[131] Sabyasachi Kamila, Mohammad Hasanuzzaman, Asif Ekbal, and Pushpak Bhattacharyya. Tempo-HindiWordNet: a lexical knowledge-base for temporal information processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(2):19:1–19:??, February 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/ft_gateway.cfm?id=3277504.

Alnawas:2019:SAI

[132] Anwar Alnawas and Nursal Arici. Sentiment analysis of Iraqi Arabic dialect on Facebook based on distributed representations of documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):20:1–20:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3278605.

Singh:2019:OHG

[133] Sukhdeep Singh and Anuj Sharma. Online handwritten Gurmukhi words recognition: an inclusive study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):21:1–21:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3282441.

Yucesoy:2019:COW

[134] Veysel Yücesoy and Aykut Koç. Co-occurrence weight selection in generation of word embeddings for low resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 18(3):22:1–22:??, July 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3282443.

Bounhas:2019:UCA

[135] Ibrahim Bounhas. On the usage of a classical Arabic corpus as a language resource: Related research and key challenges. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):23:1–23:??, July 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3277591.

Jung:2019:MPN

[136] Sangkeun Jung, Cheon-Eum Park, and Changki Lee. Multitask pointer network for Korean dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3): 24:1–24:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3282442.

Bolucu:2019:UJP

[137] Necva Bölücü and Burcu Can. Unsupervised joint PoS tagging and stemming for agglutinative languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):25:1–25:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3292398.

Kang:2019:SDR

[138] Xiaomian Kang, Chengqing Zong, and Nianwen Xue. A survey of discourse representations for Chinese discourse annotation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):26:1–26:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3293442.

Badaro:2019:SOM

[139] Gilbert Badaro, Ramy Baly, Hazem Hajj, Wassim El-Hajj, Khaled Bashir Shaban, Nizar Habash, Ahmad Al-Sallab, and Ali Hamdi. A survey of opinion mining in Arabic: a comprehensive system perspective covering challenges and advances in tools, resources, models, applications, and visualizations. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):27:1–27:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3295662.

Masmoudi:2019:ADR

[140] Abir Masmoudi, Salima Mdhaffar, Rahma Sellami, and Lamia Hadrich Belguith. Automatic diacritics restoration for Tunisian dialect. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):28:1–28:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3297278.

Rudra:2019:IAD

[141] Koustav Rudra, Ashish Sharma, Kalika Bali, Monojit Choudhury, and Niloy Ganguly. Identifying and analyzing different aspects of English– Hindi code-switching in Twitter. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):29:1–29:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3314935.

Verma:2019:CAH

[142] Pradeepika Verma, Sukomal Pal, and Hari Om. A comparative analysis on Hindi and English extractive text summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 18(3):30:1–30:??, July 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3308754.

Wei:2019:ROD

[143] Bingzhen Wei, Xuancheng Ren, Yi Zhang, Xiaoyan Cai, Qi Su, and Xu Sun. Regularizing output distribution of abstractive Chinese social media text summarization for improved semantic consistency. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):31:1–31:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3314934.

Trieu:2019:LAR

[144] Hai-Long Trieu, Duc-Vu Tran, Ashwin Ittoo, and Le-Minh Nguyen. Leveraging additional resources for improving statistical machine translation on Asian low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3): 32:1–32:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3314936.

Dehghan:2019:CDS

[145] Mohammad Hossein Dehghan and Heshaam Faili. Converting dependency structure into Persian phrase structure. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):33:1–33:??, July 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3314937.

Awais:2019:RDI

[146] Muhammad Awais and Muhammad Shoaib. Role of discourse information in Urdu sentiment classification: a rule-based method and machinelearning technique. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):34:1–34:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3300050.

Nongmeikapam:2019:HMM

[147] Kishorjit Nongmeikapam, Kanan Wahengbam, Oinam Nickson Meetei, and Themrichon Tuithung. Handwritten Manipuri Meetei–Mayek classification using convolutional neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):35:1–35:??, August 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3309497.

Gao:2019:SBC

[148] Shengxiang Gao, Jihao Huang, Mingya Xue, Zhengtao Yu, Zhuo Wang, and Yang Zhang. Syntax-based Chinese–Vietnamese tree-to-tree statistical machine translation with bilingual features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):36:1–36:??, August 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3314938.

Sun:2019:NSP

[149] Ruiyong Sun, Yijia Zhao, Qi Zhang, Keyu Ding, Shijin Wang, and Cui Wei. A neural semantic parser for math problems incorporating multisentence information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):37:1–37:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3314939.

Maimaiti:2019:MRT

[150] Mieradilijiang Maimaiti, Yang Liu, Huanbo Luan, and Maosong Sun. Multi-round transfer learning for low-resource NMT using multiple highresource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):38:1–38:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3314945.

Ihasz:2019:SFS

[151] Peter Lajos Ihasz, Mate Kovacs, Ian Piumarta, and Victor V. Kryssanov. A supplementary feature set for sentiment analysis in Japanese dialogues. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):39:1–39:??, August 2019. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3310283.

Saeed:2019:SAC

[152] Ali Saeed, Rao Muhammad Adeel Nawab, Mark Stevenson, and Paul Rayson. A sense annotated corpus for all-words Urdu word sense disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):40:1–40:??, August 2019. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3314940.

Dahou:2019:MCE

[153] Abdelghani Dahou, Shengwu Xiong, Junwei Zhou, and Mohamed Abd Elaziz. Multi-channel embedding convolutional neural network model for Arabic sentiment classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):41:1– 41:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3314941.

Onyenwe:2019:TEI

[154] Ikechukwu E. Onyenwe, Mark Hepple, Uchechukwu Chinedu, and Ignatius Ezeani. Toward an effective Igbo part-of-speech tagger. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):42:1–42:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3314942.

Costa-Jussa:2019:CCN

[155] Marta R. Costa-Jussà, Noé Casas, Carlos Escolano, and José A. R. Fonollosa. Chinese–Catalan: a neural machine translation approach based on pivoting and attention mechanisms. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):43:1– 43:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3312575.

Yu:2019:MTE

[156] Hui Yu, Weizhi Xu, Shouxun Lin, and Qun Liu. Machine translation evaluation metric based on dependency parsing model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):44:1–44:??, August 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3312573.

Liu:2019:EBC

[157] Yang Liu, Shaonan Wang, Jiajun Zhang, and Chengqing Zong. Experience-based causality learning for intelligent agents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):45:1–45:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3314943.

Yin:2019:PTE

[158] Yongjing Yin, Jinsong Su, Huating Wen, Jiali Zeng, Yang Liu, and Yidong Chen. POS tag-enhanced coarse-to-fine attention for neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):46:1–46:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3321124.

Yang:2019:MEA

[159] Jun Yang, Runqi Yang, Hengyang Lu, Chongjun Wang, and Junyuan Xie. Multi-entity aspect-based sentiment analysis with context, entity, aspect memory and dependency information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):47:1–47:??, August 2019. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3321125.

Kim:2019:MTS

[160] Hyun Kim, Jong-Hyeok Lee, and Seung-Hoon Na. Multi-task stack propagation for neural quality estimation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(4):48:1– 48:??, August 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=3321127.

Wang:2020:GCL

[161] Hongmin Wang, Jie Yang, and Yue Zhang. From Genesis to Creole language: Transfer learning for Singlish universal dependencies parsing and POS tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–29, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3321128.

Kong:2020:CZP

[162] Fang Kong, Min Zhang, and Guodong Zhou. Chinese zero pronoun resolution: a chain-to-chain approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–21, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3321129.

Yin:2020:CZP

[163] Qingyu Yin, Weinan Zhang, Yu Zhang, and Ting Liu. Chinese zero pronoun resolution: a collaborative filtering-based approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–20, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3325884.

Das:2020:TCT

[164] Ayan Das and Sudeshna Sarkar. Transform, combine, and transfer: Delexicalized transfer parser for low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–30, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3325886.

Ding:2020:TBM

[165] Chenchen Ding, Hnin Thu Zar Aye, Win Pa Pa, Khin Thandar Nwet, Khin Mar Soe, Masao Utiyama, and Eiichiro Sumita. Towards Burmese (Myanmar) morphological analysis: Syllable-based tokenization and part-of-speech tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–34, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3325885.

Liu:2020:AMC

[166] Dayiheng Liu, Kexin Yang, Qian Qu, and Jiancheng Lv. Ancient–modern Chinese translation with a new large training dataset. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–13, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3325887.

Wang:2020:CSP

[167] Wei Wang, Degen Huang, and Jingxiang Cao. Chinese syntax parsing based on sliding match of semantic string. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1): 1–14, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3329707.

Kanwal:2020:UNE

[168] Safia Kanwal, Kamran Malik, Khurram Shahzad, Faisal Aslam, and Zubair Nawaz. Urdu named entity recognition: Corpus generation and deep learning applications. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–13, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3329710.

Liu:2020:DCW

[169] Yijia Liu, Wanxiang Che, Yuxuan Wang, Bo Zheng, Bing Qin, and Ting Liu. Deep contextualized word embeddings for universal dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–17, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3326497.

Mehmood:2020:SAR

[170] Khawar Mehmood, Daryl Essam, Kamran Shafi, and Muhammad Kamran Malik. Sentiment analysis for a resource poor language —Roman Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–15, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3329709.

Bakhshaei:2020:MGM

[171] Somayeh Bakhshaei, Reza Safabakhsh, and Shahram Khadivi. Matching graph, a method for extracting parallel information from comparable corpora. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–29, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3329713.

Liu:2020:FTV

[172] Dayiheng Liu, Yang Xue, Feng He, Yuanyuan Chen, and Jiancheng Lv. μforcing: Training variational recurrent autoencoders for text generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–17, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3341110.

Srivastava:2020:AMA

[173] Jyoti Srivastava, Sudip Sanyal, and Ashish Kumar Srivastava. An automatic and a machine-assisted method to clean bilingual corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–19, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3342351.

Prakash:2020:ISP

[174] Jeena J. Prakash, Golda Brunet Rajan, and Hema A. Murthy. Importance of signal processing cues in transcription correction for low-resource Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–26, January 2020. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3342352.

Han:2020:EMW

[175] Dong Han, Junhui Li, Yachao Li, Min Zhang, and Guodong Zhou. Explicitly modeling word translations in neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–17, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3342353.

Chakrabarty:2020:NNM

[176] Abhisek Chakrabarty, Akshay Chaturvedi, and Utpal Garain. Neu-Morph: Neural morphological tagging for low-resource languages — an experimental study for Indic languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1): 1–19, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3342354.

Ji:2020:ATU

[177] Yatu Ji, Hongxu Hou, Junjie Chen, and Nier Wu. Adversarial training for unknown word problems in neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(1):1–12, January 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/abs/

10.1145/3342482.

Zhu:2020:OSK

[178] Qingfu Zhu, Weinan Zhang, Lei Cui, and Ting Liu. Order-sensitive keywords based response generation in open-domain conversational systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):18:1–18:18, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3343258.

Zhou:2020:NCG

[179] Guangyou Zhou, Yizhen Fang, Yehong Peng, and Jiaheng Lu. Neural conversation generation with auxiliary emotional supervised models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):19:1–19:17, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3344788.

Zhu:2020:EDC

[180] Wenhao Zhu, Xin Jin, Shuang Liu, Zhiguo Lu, Wu Zhang, Ke Yan, and Baogang Wei. Enhanced double-carrier word embedding via phonetics and writing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):20:1–20:18, March 2020. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3344920.

Dehkharghani:2020:SPP

[181] Rahim Dehkharghani. SentiFars: a Persian polarity lexicon for sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):21:1–21:12, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3345627.

Abdulhameed:2020:WVT

[182] Tiba Zaki Abdulhameed, Imed Zitouni, and Ikhlas Abdel-Qader. Wasf-Vec: Topology-based word embedding for modern standard Arabic and Iraqi dialect ontology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):22:1–22:27, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3345517.

Xu:2020:EPS

[183] Ge Xu, Xiaoyan Yang, Yuanzheng Cai, Zhiqiang Ruan, Tao Wang, and Xiangwen Liao. Extracting polarity shifting patterns from any corpus based on natural annotation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):23:1–23:16, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3345518.

Imankulova:2020:FPP

[184] Aizhan Imankulova, Takayuki Sato, and Mamoru Komachi. Filtered pseudo-parallel corpus improves low-resource neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):24:1–24:16, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3341726.

Gupta:2020:DNN

[185] Deepak Gupta, Asif Ekbal, and Pushpak Bhattacharyya. A deep neural network framework for English Hindi question answering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):25:1–25:22, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3359988.

Yu:2020:LWT

[186] Hongfei Yu, Xiaoqing Zhou, Xiangyu Duan, and Min Zhang. Layerwise de-training and re-training for ConvS2S machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):26:1–26:15, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3358414.

Somsap:2020:IDW

[187] Sittichai Somsap and Pusadee Seresangtakul. Isarn Dharma word segmentation using a statistical approach with named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):27:1–27:16, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3359990.

Abbas:2020:PIR

[188] Muhammad Raihan Abbas and Khadim Hussain Asif. Punjabi to ISO 15919 and Roman transliteration with phonetic rectification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):28:1–28:20, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3359991.

Beseiso:2020:SAM

[189] Majdi Beseiso and Haytham Elmousalami. Subword attentive model for Arabic sentiment analysis: a deep learning approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):29:1–29:17, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3360016.

Harikrishna:2020:CSC

[190] D. M. Harikrishna and K. Sreenivasa Rao. Children’s story classification in Indian languages using linguistic and keyword-based features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):30:1–30:22, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3342356.

Jung:2020:WRT

[191] Hun-Young Jung, Jong-Hyeok Lee, Eunju Min, and Seung-Hoon Na. Word reordering for translation into Korean sign language using syntactically-guided classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):31:1– 31:20, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3357612.

Masmoudi:2020:TAA

[192] Abir Masmoudi, Mariem Ellouze Khmekhem, Mourad Khrouf, and Lamia Hadrich Belguith. Transliteration of Arabizi into Arabic script for Tunisian dialect. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):32:1–32:21, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3364319.

Mukherjee:2020:FST

[193] Subham Mukherjee, Pradeep Kumar, and Partha Pratim Roy. Fusion of spatio-temporal information for Indic word recognition combining online and offline text data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(2):33:1–33:24, March 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3364533.

Yu:2020:ELR

[194] Zhiqiang Yu, Zhengtao Yu, Junjun Guo, Yuxin Huang, and Yonghua Wen. Efficient low-resource neural machine translation with reread and feedback mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):34:1–34:13, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3365244.

Park:2020:NSB

[195] Cheoneum Park, Heejun Song, and Changki Lee. S^{3}-NET: SRU-based sentence and self-matching networks for machine reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):35:1–35:14, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3365679.

Sarwar:2020:SSF

[196] Raheem Sarwar, Thanasarn Porthaveepong, Attapol Rutherford, Thanawin Rakthanmanon, and Sarana Nutanong. StyloThai:: a scalable framework for stylometric authorship identification of Thai documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):36:1–36:15, February 2020. CODEN ????

ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3365832.

Kim:2020:UIB

[197] Hyun Kim and Seung-Hoon Na. Uniformly interpolated balancing for robust prediction in translation quality estimation: a case study of English– Korean translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):37:1–37:27, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3365916.

Zhou:2020:LMU

[198] Xiao Zhou, Zhen-Hua Ling, and Li-Rong Dai. Learning and modeling unit embeddings using deep neural networks for unit-selection-based Mandarin speech synthesis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):38:1–38:14, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3372244.

Mirzaei:2020:SRL

[199] Azadeh Mirzaei, Fatemeh Sedghi, and Pegah Safari. Semantic role labeling system for Persian language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):39:1– 39:12, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3372246.

Ding:2020:BMT

[200] Chenchen Ding, Sann Su Su Yee, Win Pa Pa, Khin Mar Soe, Masao Utiyama, and Eiichiro Sumita. A Burmese (Myanmar) Treebank: Guideline and analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):40:1–40:13, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3373268.

Song:2020:KPS

[201] Hyun-Je Song and Seong-Bae Park. Korean part-of-speech tagging based on morpheme generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):41:1–41:10, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3373608.

Mi:2020:LIL

[202] Chenggang Mi, Lei Xie, and Yanning Zhang. Loanword identification in low-resource languages with minimal supervision. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):43:1–43:22, February 2020. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3374212.

Li:2020:INM

[203] Yachao Li, Junhui Li, Min Zhang, Yixin Li, and Peng Zou. Improving neural machine translation with linear interpolation of a short-path unit. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):44:1–44:16, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3377851.

Liu:2020:DUK

[204] Xiao-Yang Liu, Yimeng Zhang, Yukang Liao, and Ling Jiang. Dynamic updating of the knowledge base for a large-scale question answering system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):45:1–45:13, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3377708.

Liu:2020:ELM

[205] Shih-Hung Liu, Kuan-Yu Chen, and Berlin Chen. Enhanced language modeling with proximity and sentence relatedness information for extractive broadcast news summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):46:1– 46:19, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3377407.

Du:2020:CNL

[206] Qianlong Du, Chengqing Zong, and Keh-Yih Su. Conducting natural language inference with word-pair-dependency and local context. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(3):47:1–47:23, February 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3377704.

Zitouni:2020:ENE

[207] Imed Zitouni. Editorial from the new Editor-in-Chief: the era of natural language processing innovations on Asian and low-resource languages.

ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):48e:1–48e:2, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3397501.

Yang:2020:OEQ

[208] Jingxuan Yang, Haotian Cui, Si Li, Sheng Gao, Jun Guo, and Zhengdong Lu. Outline extraction with question-specific memory cells. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):48:1–48:17, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3377707.

Zarnoufi:2020:MNB

[209] Randa Zarnoufi, Hamid Jaafar, and Mounia Abik. Machine normalization: Bringing social media text from non-standard to standard form. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):49:1–49:30, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3378414.

Bhattu:2020:ICM

[210] S. Nagesh Bhattu, Satya Krishna Nunna, D. V. L. N. Somayajulu, and Binay Pradhan. Improving code-mixed POS tagging using code-mixed embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):50:1–50:31, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3380967.

Ahmad:2020:NER

[211] Muhammad Tayyab Ahmad, Muhammad Kamran Malik, Khurram Shahzad, Faisal Aslam, Asif Iqbal, Zubair Nawaz, and Faisal Bukhari. Named entity recognition and classification for Punjabi Shahmukhi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):51:1–51:13, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3383306.

Garg:2020:HES

[212] Kanika Garg and D. K. Lobiyal. Hindi EmotionNet: a scalable emotion lexicon for sentiment classification of Hindi text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):52:1–52:35, July 2020. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3383330.

Udomcharoenchaikit:2020:AER

[213] Can Udomcharoenchaikit, Prachya Boonkwan, and Peerapon Vateekul. Adversarial evaluation of robust neural sequential tagging methods for Thai language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):53:1–53:25, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3383201.

Sugandhi:2020:SLG

[214] Sugandhi, Parteek Kumar, and Sanmeet Kaur. Sign language generation system based on Indian sign language grammar. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):54:1–54:26, July 2020. CODEN ???? ISSN 2375-4699 (print),

2375-4702 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3384202.

Sarwar:2020:NLI

[215] Raheem Sarwar, Attapol T. Rutherford, Saeed-Ul Hassan, Thanawin Rakthanmanon, and Sarana Nutanong. Native language identification of fluent and advanced non-native writers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4): 55:1–55:19, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/abs/10.1145/3383202.

Buyuk:2020:CDS

[216] Osman Büyük. Context-dependent sequence-to-sequence Turkish spelling correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):56:1–56:16, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3383200.

Khalil:2020:EAC

[217] Hussein Khalil, Taha Osman, and Mohammed Miltan. Extracting Arabic composite names using genitive principles of Arabic grammar. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):57:1–57:16, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3382187.

Wang:2020:SCH

[218] Kexin Wang, Yu Zhou, Jiajun Zhang, Shaonan Wang, and Chengqing Zong. Structurally comparative hinge loss for dependency-based neural text representation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):58:1–58:19, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3387633.

Liu:2020:JME

[219] Maofu Liu, Yukun Zhang, Wenjie Li, and Donghong Ji. Joint model of entity recognition and relation extraction with self-attention mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):59:1–59:19, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3387634.

Kumar:2020:LGV

[220] H. R. Shiva Kumar and A. G. Ramakrishnan. Lipi Gnani: a versatile OCR for documents in any language printed in Kannada script. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4):60:1–60:23, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3387632.

Hao:2020:CST

[221] Ming Hao, Bo Xu, Jing-Yi Liang, Bo-Wen Zhang, and Xu-Cheng Yin. Chinese short text classification with mutual-attention convolutional neural networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):61:1–61:13, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3388970.

Xu:2020:SDE

[222] Fan Xu, Jian Luo, Mingwen Wang, and Guodong Zhou. Speechdriven end-to-end language discrimination toward Chinese dialects. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):62:1–62:24, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3389021.

Chen:2020:IIF

[223] Junjie Chen, Hongxu Hou, and Jing Gao. Inside importance factors of graph-based keyword extraction on Chinese short text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):63:1–63:15, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3388971.

Lou:2020:EBS

[224] Yinxia Lou, Yue Zhang, Fei Li, Tao Qian, and Donghong Ji. Emojibased sentiment analysis using attention networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 19(5):64:1–64:13, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3389035.

Zhou:2020:DNN

[225] Long Zhou, Jiajun Zhang, Xiaomian Kang, and Chengqing Zong. Deep neural network–based machine translation system combination. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):65:1–65:19, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3389791.

Ameur:2020:RAT

[226] Mohamed Seghir Hadj Ameur, Riadh Belkebir, and Ahmed Guessoum. Robust Arabic text categorization by combining convolutional and recurrent neural networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):66:1–66:16, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3390092.

Das:2020:SMT

[227] Ayan Das and Sudeshna Sarkar. A survey of the model transfer approaches to cross-lingual dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 19(5):67:1–67:60, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3383772.

Marie:2020:ITU

[228] Benjamin Marie and Atsushi Fujita. Iterative training of unsupervised neural and statistical machine translation systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 19(5):68:1–68:21, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3389790.

Chimalamarri:2020:MSI

[229] Santwana Chimalamarri, Dinkar Sitaram, and Ashritha Jain. Morphological segmentation to improve crosslingual word embeddings for low resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):69:1–69:15, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3390298.

Li:2020:PQA

[230] Ying Li, Jizhou Huang, Miao Fan, Jinyi Lei, Haifeng Wang, and Enhong Chen. Personalized query auto-completion for large-scale POI search at Baidu maps. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):70:1–70:16, August 2020. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3394137.

Alkhatib:2020:DLA

[231] Manar Alkhatib, Azza Abdel Monem, and Khaled Shaalan. Deep learning for Arabic error detection and correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 19(5):71:1–71:13, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3373266.

Orhan:2020:LWV

[232] Umut Orhan and Enis Arslan. Learning word-vector quantization: a case study in morphological disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):72:1– 72:18, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3397967.

Dhar:2020:CSC

[233] Ankita Dhar, Himadri Mukherjee, Niladri Sekhar Dash, and Kaushik Roy. CESS — a system to categorize Bangla Web text documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):73:1–73:18, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3398070.

Bai:2020:NCT

[234] Ruirui Bai, Zhongqing Wang, Fang Kong, Shoushan Li, and Guodong Zhou. Neural co-training for sentiment classification with product attributes. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):74:1–74:17, August 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3394113.

Schmidt:2020:GTC

[235] Dirk Schmidt. Grading Tibetan children’s literature: a test case using the NLP readability tool “Dakje”. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):75:1– 75:19, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3392046.

Habiba:2020:TCN

[236] Rabia Habiba, Muhammad Awais, and Muhammad Shoaib. A technique to calculate national happiness index by analyzing Roman Urdu messages posted on social media. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):76:1–76:16, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3400712.

Wang:2020:EFL

[237] Hao Wang, Qiongxing Tao, Siyuan Du, and Xiangfeng Luo. An extensible framework of leveraging syntactic skeleton for semantic relation classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):77:1–77:21, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3402885.

Malhas:2020:ABR

[238] Rana Malhas and Tamer Elsayed. AyaTEC: Building a reusable versebased test collection for Arabic question answering on the Holy Qur’an. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):78:1–78:21, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3400396.

Ruan:2020:CTV

[239] Yu-Ping Ruan, Zhen-Hua Ling, and Xiaodan Zhu. Conditiontransforming variational autoencoder for generating diverse short text conversations. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):79:1–79:13, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3402884.

Badaro:2020:LPA

[240] Gilbert Badaro, Hazem Hajj, and Nizar Habash. A link prediction approach for accurately mapping a large-scale Arabic lexical resource to English WordNet. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):80:1–80:38, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3404854.

Ma:2020:IWS

[241] Tinghuai Ma, Raeed Al-Sabri, Lejun Zhang, Bockarie Marah, and Najla Al-Nabhan. The impact of weighting schemes and stemming process on topic modeling of Arabic long and short texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):81:1–81:23, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3405843.

M:2020:CAH

[242] Poornima Devi. M. and M. Sornam. Classification of ancient handwritten Tamil characters on palm leaf inscription using modified adaptive backpropagation neural network with GLCM features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):82:1–82:24, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3406209.

Yang:2020:AMU

[243] Qimeng Yang, Long Yu, Shengwei Tian, and Jinmiao Song. Attention mechanism for Uyghur personal pronouns resolution. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):83:1–83:13, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3412323.

Xi:2020:GEL

[244] Xuefeng Xi, Zhou Pi, and Guodong Zhou. Global encoding for long Chinese text summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):84:1–84:17, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3407911.

Tanwar:2020:TMR

[245] Ashwani Tanwar and Prasenjit Majumder. Translating morphologically rich Indian languages under zero-resource conditions. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 19(6):85:1–85:15, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3407912.

Laatar:2020:DAW

[246] Rim Laatar, Chafik Aloulou, and Lamia Hadrich Belguith. Disambiguating Arabic words according to their historical appearance in the document based on recurrent neural networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6): 86:1–86:16, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3410569.

Chou:2020:CWN

[247] Chien-Lung Chou, Chia-Hui Chang, Yuan-Hao Lin, and Kuo-Chun Chien. On the construction of Web NER model training tool based on distant supervision. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):87:1–87:28, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3422817.

Wu:2020:DEW

[248] Chuhan Wu, Fangzhao Wu, Tao Qi, Junxin Liu, Yongfeng Huang, and Xing Xie. Detecting entities of works for Chinese chatbot. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(6):88:1–88:13, November 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3414901.

Manogaran:2021:SID

[249] Gunasekaran Manogaran, Hassan Qudrat-Ullah, and Qin Xin. Special issue on deep structured learning for natural language processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):1:1–1:2, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3436206.

Wang:2021:CDB

[250] Kun Wang, Yanpeng Cui, Jianwei Hu, Yu Zhang, Wei Zhao, and Luming Feng. Cyberbullying detection, based on the FastText and word similarity schemes. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):2:1–2:15, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3398191.

Sun:2021:DIM

[251] Chengai Sun, Liangyu Lv, Gang Tian, and Tailu Liu. Deep interactive memory network for aspect-level sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):3:1–3:12, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3402886.

Wang:2021:VTM

[252] Wei Wang, Zhiguo Gong, Jing Ren, Feng Xia, Zhihan Lv, and Wei Wei. Venue topic model-enhanced joint graph modelling for citation recommendation in scholarly big data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):4:1–4:15, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3404995.

Peng:2021:THS

[253] Lingxi Peng, Haohuai Liu, Yangang Nie, Ying Xie, Xuan Tang, and Ping Luo. The transnational happiness study with big data technology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):5:1–5:12, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3412497.

Guangce:2021:KDN

[254] Ruan Guangce and Xia Lei. Knowledge discovery of news text based on artificial intelligence. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):6:1–6:18, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3418062.

Meelen:2021:OLA

[255] Marieke Meelen, Élie Roux, and Nathan Hill. Optimisation of the largest annotated Tibetan corpus combining rule-based, memory-based, and deep-learning methods. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):7:1–7:11, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3409488.

Kumar:2021:MNF

[256] S. Rakesh Kumar, S. Muthuramalingam, and Fadi Al-Turjman. Multimodal news feed evaluation system with deep reinforcement learning approaches. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):8:1–8:12, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3414523.

Khan:2021:HSD

[257] Muhammad Moin Khan, Khurram Shahzad, and Muhammad Kamran Malik. Hate speech detection in Roman Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):9:1–9:19, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3414524.

Sun:2021:UNM

[258] Haipeng Sun, Rui Wang, Masao Utiyama, Benjamin Marie, Kehai Chen, Eiichiro Sumita, and Tiejun Zhao. Unsupervised neural machine translation for similar and distant language pairs: an empirical study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):10:1–10:17, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3418059.

Zhang:2021:HBN

[259] Peiying Zhang, Xingzhe Huang, Maozhen Li, and Yu Xue. Hybridization between neural computing and nature-inspired algorithms for a sentence similarity model based on the attention mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):11:1–11:21, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3447756.

Husain:2021:SOL

[260] Fatemah Husain and Ozlem Uzuner. A survey of offensive language detection for the Arabic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):12:1– 12:44, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3421504.

Alzubaidi:2021:RTA

[261] Mohammad A. Alzubaidi, Mwaffaq Otoom, and Nouran S. Ahmad. Realtime assistive reader pen for Arabic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (1):13:1–13:30, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3423133.

Sidig:2021:KAS

[262] Ala Addin I. Sidig, Hamzah Luqman, Sabri Mahmoud, and Mohamed Mohandes. KArSL: Arabic sign language database. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 20(1):14:1–14:19, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3423420.

Wijayanti:2021:AIS

[263] Rini Wijayanti and Andria Arisal. Automatic Indonesian sentiment lexicon curation with sentiment valence tuning for social media sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):15:1–15:16, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3425632.

Li:2021:TTS

[264] Zhongyang Li, Xiao Ding, and Ting Liu. TransBERT: a three-stage pretraining technology for story-ending prediction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (1):16:1–16:20, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3427669.

Bolucu:2021:CUM

[265] Necva Bölücü and Burcu Can. A cascaded unsupervised model for PoS tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):17:1–17:23, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3447759.

Chauhan:2021:ISC

[266] Uttam Chauhan and Apurva Shah. Improving semantic coherence of Gujarati text topic model using inflectional forms reduction and single-letter words removal. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(1):18:1–18:18, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3447760.

Congjun:2021:RDT

[267] Long Congjun and Nathan W. Hill. Recent developments in Tibetan NLP. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):19:1–19:3, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3453692.

List:2021:TSH

[268] Johann-Mattis List, Nathaniel A. Sims, and Robert Forkel. Toward a sustainable handling of interlinear-glossed text in language documentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):20:1–20:15, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3389010.

Krishna:2021:ATA

[269] Ravi Krishna, Norman Mu, and Kurt Keutzer. Applying text analytics to the mind-section literature of the Tibetan tradition of the Great Perfection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):21:1–21:32, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3392047.

Trinley:2021:TWE

[270] Ngawang Trinley, Tenzin, Dirk Schmidt, Helios Hildt, and Tenzin Kaldan. Taming the wild etext: Managing, annotating, and sharing Tibetan corpora in open spaces. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):22:1–22:23, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3418060.

Kulkarni:2021:SPF

[271] Amba Kulkarni. Sanskrit parsing following Indian theories of verbal cognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):23:1–23:38, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3418061.

Li:2021:FBS

[272] Yachao Li, Jing Jiang, Jia Yangji, and Ning Ma. Finding better subwords for Tibetan neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):24:1– 24:11, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3448216.

Long:2021:RTM

[273] Congjun Long, Xuewen Zhou, and Maoke Zhou. Recognition of Tibetan maximal-length noun phrases based on syntax tree. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):25:1–25:13, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3423324.

Shi:2021:MLC

[274] Shumin Shi, Dan Luo, Xing Wu, Congjun Long, and Heyan Huang. Multi-level chunk-based constituent-to-dependency Treebank transformation for Tibetan dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2): 26:1–26:12, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3424247.

Sun:2021:JMR

[275] Yuan Sun, Andong Chen, Chaofan Chen, Tianci Xia, and Xiaobing Zhao. A joint model for representation learning of Tibetan knowledge graph based on encyclopedia. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):27:1–27:17, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3447248.

Wang:2021:CSE

[276] Hao Wang, Bin Wang, Jianyong Duan, and Jiajun Zhang. Chinese spelling error detection using a fusion lattice LSTM. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):28:1–28:11, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3426882.

Nasution:2021:POB

[277] Arbi Haza Nasution, Yohei Murakami, and Toru Ishida. Plan optimization to bilingual dictionary induction for low-resource language families. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):29:1–29:28, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3448215.

An:2021:NDP

[278] Bo An and Congjun Long. Neural dependency parser for Tibetan sentences. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):30:1–30:16, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3429456.

Zhang:2021:USC

[279] Longtu Zhang and Mamoru Komachi. Using sub-character level information for neural machine translation of logographic languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):31:1–31:15, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3431727.

Mishra:2021:HIC

[280] Santosh Kumar Mishra, Rijul Dhir, Sriparna Saha, and Pushpak Bhattacharyya. A Hindi image caption generation framework using deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):32:1–32:19, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3432246.

Darwish:2021:ADR

[281] Kareem Darwish, Ahmed Abdelali, Hamdy Mubarak, and Mohamed Eldesouki. Arabic diacritic recovery using a feature-rich biLSTM model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):33:1–33:18, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3434235.

Binbeshr:2021:SRH

[282] Farid Binbeshr, Amirrudin Kamsin, and Manal Mohammed. A systematic review on hadith authentication and classification methods. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):34:1–34:17, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3434236.

Wang:2021:HMN

[283] Yu Wang, Yining Sun, Zuchang Ma, Lisheng Gao, and Yang Xu. A hybrid model for named entity recognition on Chinese electronic medical records. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):35:1–35:12, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3436819.

Jin:2021:HSS

[284] Guozhe Jin and Zhezhou Yu. A hierarchical sequence-to-sequence model for Korean POS tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(2):36:1–36:13, April 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3421762.

S:2021:SAA

[285] Dhivya S. and Usha Devi G. Study on automated approach to recognize characters for handwritten and historical document. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):37:1–37:24, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3396167.

Yang:2021:DDH

[286] Xiaodong Yang and Xiaoxia Lin. Design and development of heuristic utility management algorithm for Chinese library management system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):38:1–38:13, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3397968.

S:2021:THT

[287] Dhivya S. and Usha Devi G. TAMIZHI: Historical tamil-brahmi script recognition using CNN and MobileNet. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3): 39:1–39:26, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3402891.

Jiang:2021:TLB

[288] Peipei Jiang, Liailun Chen, and Min-Feng Wang. Transfer learning based recurrent neural network algorithm for linguistic analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):40:1–40:16, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3406204.

Fang:2021:HBG

[289] Hui Fang, Hongmei Shi, and Jiuzhou Zhang. Heuristic bilingual graph corpus network to improve English instruction methodology based on statistical translation approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):41:1–41:14, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3406205.

Jamal:2021:DLB

[290] Nasir Jamal, Chen Xianqiao, Fadi Al-Turjman, and Farhan Ullah. A deep learning-based approach for emotions classification in big corpus of imbalanced tweets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):42:1–42:16, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3410570.

Muthu:2021:FET

[291] Balaanand Muthu, Sivaparthipan Cb, Priyan Malarvizhi Kumar, Seifedine Nimer Kadry, Ching-Hsien Hsu, Oscar Sanjuan, and Ruben Gonzalez Crespo. A framework for extractive text summarization based on deep learning modified neural network classifier. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3): 45:1–45:20, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3392048.

Wang:2021:AEA

[292] Ailing Wang, Jie Sun, and Leiming Li. An analysis for elements of affecting the establishment and promotion of micro-business trust in C2C model under WeChat circumstance. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):46:1– 46:11, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3398011.

Wang:2021:SGD

[293] Erlu Wang, Priyan Malarvizhi Kumar, and R. Dinesh Jackson Samuel. Semantic graphical dependence parsing model in improving English teaching abilities. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):48:1–48:14, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3425633.

P:2021:TST

[294] Ashokkumar P., Siva Shankar G., Gautam Srivastava, Praveen Kumar Reddy Maddikunta, and Thippa Reddy Gadekallu. A two-stage text feature selection algorithm for improving text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):49:1–49:19, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3425781.

Quamer:2021:SSA

[295] Waris Quamer, Praphula Kumar Jain, Arpit Rai, Vijayalakshmi Saravanan, Rajendra Pamula, and Chiranjeev Kumar. SACNN: Self-attentive convolutional neural network model for natural language inference. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):50:1–50:16, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3426884.

Liao:2021:ENO

[296] Hsiu-Li Liao, Zhen-Yu Huang, and Su-Houn Liu. The effects of negative online reviews on consumer perception, attitude and purchase intention: Experimental investigation of the amount, quality, and presentation order of eWOM. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):51:1–51:21, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3426883.

Seifollahi:2021:EBT

[297] Sattar Seifollahi, Massimo Piccardi, and Alireza Jolfaei. An embeddingbased topic model for document classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):52:1–52:13, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3431728.

Li:2021:DSL

[298] Yong Li, Xiaojun Yang, Min Zuo, Qingyu Jin, Haisheng Li, and Qian Cao. Deep structured learning for natural language processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(3):53:1–53:14, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3433538.

Mao:2021:NJM

[299] Cunli Mao, Zhibo Man, Zhengtao Yu, Shengxiang Gao, Zhenhan Wang, and Hongbin Wang. A neural joint model with BERT for Burmese syllable segmentation, word segmentation, and POS tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):54:1–54:23, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3436818.

E:2021:AMP

[300] Manjunath K. E., Srinivasa Raghavan K. M., K. Sreenivasa Rao, Dinesh Babu Jayagopi, and V. Ramasubramanian. Approaches for multilingual phone recognition in code-switched and non-code-switched scenarios using Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):55:1–55:19, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3437256.

Kumar:2021:NAM

[301] Mohinder Kumar, Manish Kumar Jindal, and Munish Kumar. A novel attack on monochrome and greyscale Devanagari CAPTCHAs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):56:1–56:30, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3439798.

Lin:2021:FIG

[302] Nankai Lin, Boyu Chen, Xiaotian Lin, Kanoksak Wattanachote, and Shengyi Jiang. A framework for Indonesian grammar error correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):57:1–57:12, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3440993.

Shivachi:2021:LSU

[303] Casper Shikali Shivachi, Refuoe Mokhosi, Zhou Shijie, and Liu Qihe. Learning syllables using Conv-LSTM model for Swahili word representation and part-of-speech tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):58:1–58:25, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3445975.

Ranathunga:2021:SAS

[304] Surangika Ranathunga and Isuru Udara Liyanage. Sentiment analysis of Sinhala news comments. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):59:1–59:23, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3445035.

Zhu:2021:GBM

[305] Junnan Zhu, Lu Xiang, Yu Zhou, Jiajun Zhang, and Chengqing Zong. Graph-based multimodal ranking models for multimodal summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):60:1–60:21, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3445794.

Lalrempuii:2021:IEM

[306] Candy Lalrempuii, Badal Soni, and Partha Pakray. An improved English-to-Mizo neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):61:1–61:21, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3445974.

Thin:2021:TNL

[307] Dang Van Thin, Ngan Luu-Thuy Nguyen, Tri Minh Truong, Lac Si Le, and Duy Tin Vo. Two new large corpora for Vietnamese aspect-based sentiment analysis at sentence level. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):62:1– 62:22, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3446678.

Alian:2021:BAP

[308] Marwah Alian, Arafat Awajan, Ahmad Al-Hasan, and Raeda Akuzhia. Building Arabic paraphrasing benchmark based on transformation rules. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):63:1–63:17, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3446770.

Prabhakar:2021:QET

[309] Dinesh Kumar Prabhakar, Sukomal Pal, and Chiranjeev Kumar. Query expansion for transliterated text retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):

64:1–64:34, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3447649.

Taghizadeh:2021:CLA

[310] Nasrin Taghizadeh and Heshaam Faili. Cross-lingual adaptation using universal dependencies. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):65:1–65:23, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3448251.

Coban:2021:FTM

[311] Önder Çoban, Ali Inan, and Selma Ayse Özel. Facebook tells me your gender: an exploratory study of gender prediction for Turkish Facebook users. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):66:1–66:38, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3448253.

Qi:2021:DPB

[312] Shanshan Qi, Limin Zheng, and Feiyu Shang. Dependency parsingbased entity relation extraction over Chinese complex text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):67:1–67:34, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3450273.

Mousavi:2021:DPW

[313] Zahra Mousavi and Heshaam Faili. Developing the Persian Wordnet of verbs using supervised learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):68:1–68:18, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3450969.

Arora:2021:SSR

[314] Karunesh Kumar Arora and Shyam Sunder Agrawal. Source-side reordering to improve machine translation between languages with distinct word orders. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):69:1–69:18, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3448252.

Kumari:2021:RNS

[315] Divya Kumari, Asif Ekbal, Rejwanul Haque, Pushpak Bhattacharyya, and Andy Way. Reinforced NMT for sentiment and content preservation in low-resource scenario. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(4):70:1–70:27, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3450970.

Sathish:2021:ISA

[316] R. Sathish and P. Ezhumalai. Intermodal sentiment analysis for images with text captions using the VGGNET technique. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (4):71:1–71:14, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3450971.

Manogaran:2021:ISI

[317] Gunasekaran Manogaran, Hassan Qudrat-Ullah, and Qin Xin. Introduction to the special issue on deep structured learning for natural language processing, Part 3. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):72e:1–72e:3, September 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3476464.

Tan:2021:RPT

[318] Junyang Tan, Dan Xia, Shiyun Dong, Honghao Zhu, and Binshi Xu. Research on pre-training method and generalization ability of big data recognition model of the Internet of Things. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5): 72:1–72:15, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3433539.

Li:2021:SAQ

[319] Yarong Li. Sequence alignment with Q-learning based on the actor– critic model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):73:1–73:7, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3433540.

Naseem:2021:CSW

[320] Usman Naseem, Imran Razzak, Shah Khalid Khan, and Mukesh Prasad. A comprehensive survey on word representation models: From classical to state-of-the-art word representation language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):74:1–74:35, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3434237.

Ji:2021:OAP

[321] Xiaowen Ji and Jincheng Ni. An OT-ET analysis of Polish singular– plural pairs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):75:1–75:12, May 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3434238.

Jing:2021:GGM

[322] Weipeng Jing, Xianyang Song, Donglin Di, and Houbing Song. geoGAT: Graph model based on attention mechanism for geographic text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):76:1–76:18, September 2021. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3434239.

Bi:2021:BDL

[323] Mingwen Bi, Qingchuan Zhang, Min Zuo, Zelong Xu, and Qingyu Jin. Bi-directional long short-term memory model with semantic positional attention for the question answering system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5): 77:1–77:13, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3439800.

Fan:2021:DNN

[324] Xiaoqian Fan, Bowen Yang, Wenzhi Chen, and Quanfang Fan. Deep neural network based noised Asian speech enhancement and its implementation on a hearing aid app. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):78:1–78:14, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3439797.

Zhao:2021:MOH

[325] Chunhe Zhao, Balaanand Muthu, and P. Mohamed Shakeel. Multiobjective heuristic decision making and benchmarking for mobile applications in English language learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):79:1– 79:16, September 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3439799.

Gupta:2021:TIC

[326] Vedika Gupta, Nikita Jain, Shubham Shubham, Agam Madan, Ankit Chaudhary, and Qin Xin. Toward integrated CNN-based sentiment analysis of tweets for scarce-resource language-Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (5):80:1–80:23, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3450447.

Do:2021:DVT

[327] Phuc Do, Truong H. V. Phan, and Brij B. Gupta. Developing a Vietnamese tourism question answering system using knowledge graph and deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):81:1–81:18, June 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3453651.

Li:2021:REU

[328] Meng Li. Research on extraction of useful tourism online reviews based on multimodal feature fusion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):82:1–82:16, September 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3453694.

Sun:2021:TCA

[329] Lin Sun, Wenzheng Xu, and Jimin Liu. Two-channel attention mechanism fusion model of stock price prediction based on CNN-LSTM. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):83:1–83:12, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3453693.

Jain:2021:HCL

[330] Praphula Kumar Jain, Vijayalakshmi Saravanan, and Rajendra Pamula. A hybrid CNN-LSTM: a deep learning approach for consumer sentiment analysis using qualitative user-generated contents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (5):84:1–84:15, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3457206.

Deng:2021:CCB

[331] Fei Deng and Timothy V. Rasinski. A computer corpus-based study of Chinese EFL learners’ use of adverbial connectors and its implications for building a language-based learning environment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20 (5):85:1–85:16, June 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3457987.

Deng:2021:CPC

[332] Yongliang Deng and Hua Zhang. Configurational path to Chinese reading stickiness of digital library. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):86:1–86:18, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3459092.

Javed:2021:BSS

[333] Abdul Rehman Javed, Saif Ur Rehman, Mohib Ullah Khan, Mamoun Alazab, and Habib Ullah Khan. Betalogger: Smartphone sensor-based side-channel attack detection and text inference using language modeling and dense MultiLayer neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):87:1– 87:17, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3460392.

Lavanya:2021:MRS

[334] R. Lavanya and B. Bharathi. Movie recommendation system to solve data sparsity using collaborative filtering approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):88:1–88:14, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3459091.

Ma:2021:IAV

[335] Jun Ma, Hongzhi Yu, Yan Xu, and Kaiying Deng. An investigational approach for vowels of the Salar language based on a database of speech acoustic parameters. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):89:1–89:10, September 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3459927.

Kumar:2021:SAU

[336] Akshi Kumar and Victor Hugo C. Albuquerque. Sentiment analysis using XLM-R transformer and zero-shot transfer learning on resource-poor Indian language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):90:1–90:13, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3461764.

Zhou:2021:NRO

[337] Zhou Zhou, Fangmin Li, and Shuiqiao Yang. A novel resource optimization algorithm based on clustering and improved differential evolution strategy under a cloud environment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(5):91:1– 91:15, July 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3462761.

Tan:2021:BBT

[338] Minghuan Tan, Jing Jiang, and Bing Tian Dai. A BERT-based twostage model for Chinese Chengyu recommendation. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 20(6):92:1–92:18, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3453185.

Xiang:2021:RCL

[339] Lu Xiang, Junnan Zhu, Yang Zhao, Yu Zhou, and Chengqing Zong. Robust cross-lingual task-oriented dialogue. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6): 93:1–93:24, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3457571.

Premjith:2021:DLA

[340] B. Premjith and K. P. Soman. Deep learning approach for the morphological synthesis in Malayalam and Tamil at the character level. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):94:1–94:17, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3457976.

Mundotiya:2021:LRB

[341] Rajesh Kumar Mundotiya, Manish Kumar Singh, Rahul Kapur, Swasti Mishra, and Anil Kumar Singh. Linguistic resources for Bhojpuri, Magahi, and Maithili: Statistics about them, their similarity estimates, and baselines for three applications. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):95:1–95:37, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3458250.

Telemala:2021:ETL

[342] Joseph P. Telemala and Hussein Suleman. Exploring topic-language preferences in multilingual Swahili information retrieval in Tanzania. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):96:1–96:30, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3458671.

Tian:2021:RRO

[343] Yaolin Tian, Weize Gao, Xuxing Liu, Shanxiong Chen, and Bofeng Mo. The research on rejoining of the oracle bone rubbings based on curve matching. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):97:1–97:17, November 2021. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3460393.

Munir:2021:NUS

[344] Kashif Munir, Hai Zhao, and Zuchao Li. Neural unsupervised semantic role labeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):98:1–98:16, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3461613.

Saha:2021:UDM

[345] Tulika Saha, Dhawal Gupta, Sriparna Saha, and Pushpak Bhattacharyya. A unified dialogue management strategy for multi-intent dialogue conversations in multiple languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6): 99:1–99:22, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3461763.

Ni:2021:MDT

[346] Weijian Ni, Tong Liu, Qingtian Zeng, and Nengfu Xie. Mining domain terminologies using search engine’s query log. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 20(6):100:1–100:32, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3462327.

Xu:2021:CDG

[347] Jun Xu, Zeyang Lei, Haifeng Wang, Zheng-Yu Niu, Hua Wu, Wanxiang Che, Jizhou Huang, and Ting Liu. Coherent dialog generation with query graph. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):101:1–101:23, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3462551.

Joshi:2021:SSG

[348] Manju Lata Joshi, Nisheeth Joshi, and Namita Mittal. SGATS: Semantic graph-based automatic text summarization from Hindi text documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):102:1–102:32, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3464381.

Byambadorj:2021:NTM

[349] Zolzaya Byambadorj, Ryota Nishimura, Altangerel Ayush, and Norihide Kitaoka. Normalization of transliterated Mongolian words using Seq2Seq model with limited data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):103:1– 103:19, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3464361.

Kaing:2021:TTP

[350] Hour Kaing, Chenchen Ding, Masao Utiyama, Eiichiro Sumita, Sethserey Sam, Sopheap Seng, Katsuhito Sudoh, and Satoshi Nakamura. Towards tokenization and part-of-speech tagging for Khmer: Data and discussion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):104:1–104:16, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3464378.

Tian:2021:NCM

[351] Xiuxia Tian, Can Li, and Bo Zhao. A novel classification model SA-MPCNN for power equipment defect text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):105:1–105:21, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3464380.

Sen:2021:BGT

[352] Shibaprasad Sen, Ankan Bhattacharyya, Ram Sarkar, and Kaushik Roy. BYANJON: a ground truth preparation system for online handwritten Bangla documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):106:1–106:16, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3464379.

Maimaiti:2021:IDA

[353] Mieradilijiang Maimaiti, Yang Liu, Huanbo Luan, Zegao Pan, and Maosong Sun. Improving data augmentation for low-resource NMT guided by POS-tagging and paraphrase embedding. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 20(6):107:1–107:21, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3464427.

Vo:2021:SIS

[354] Tham Vo. SE4ExSum: an integrated semantic-aware neural approach with graph convolutional network for extractive text summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):108:1–108:22, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3464426.

Liu:2021:MCS

[355] Lei Liu, Hao Chen, and Yinghong Sun. A multi-classification sentiment analysis model of Chinese short text based on gated linear units and attention mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):109:1–109:13, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3464425.

Niwa:2021:CCR

[356] Ayana Niwa, Naoaki Okazaki, Kohei Wakimoto, Keisuke Nishiguchi, and Masataka Mouri. Construction of a corpus of rhetorical devices in slogans and structural analysis of antitheses. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):110:1– 110:26, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3465218.

Shin:2021:EEA

[357] Jaehun Shin, Wonkee Lee, Byung-Hyun Go, Baikjin Jung, Youngkil Kim, and Jong-Hyeok Lee. Exploration of effective attention strategies for neural automatic post-editing with transformer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 20(6):111:1–111:17, November 2021. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3465383.

Kumar:2022:ISI

[358] Akshi Kumar, Christian Esposito, and Dimitrios A. Karras. Introduction to special issue on misinformation, fake news and rumor detection in low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):1e:1–1e:3, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3505588.

Sangwan:2022:DCD

[359] Saurabh R. Sangwan and M. P. S. Bhatia. Denigrate comment detection in low-resource Hindi language using attention-based residual networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):1:1–1:14, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3431729.

Bhowmick:2022:MDF

[360] Rajat Subhra Bhowmick, Isha Ganguli, Jayanta Paul, and Jaya Sil. A multimodal deep framework for derogatory social media post identification of a recognized person. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):2:1–2:19, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3447651.

Jain:2022:FNC

[361] Rachna Jain, Deepak Kumar Jain, Dharana, and Nitika Sharma. Fake news classification: a quantitative research description. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):3:1–3:17, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3447650.

Ranasinghe:2022:MOL

[362] Tharindu Ranasinghe and Marcos Zampieri. Multilingual offensive language identification for low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (1):4:1–4:13, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3457610.

Das:2022:DLA

[363] Soma Das, Pooja Rai, and Sanjay Chatterji. Deep level analysis of legitimacy in Bengali news sentences. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):5:1– 5:18, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3459928.

Saeed:2022:ECE

[364] Ramsha Saeed, Hammad Afzal, Haider Abbas, and Maheen Fatima. Enriching conventional ensemble learner with deep contextual semantics to detect fake news in Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):6:1–6:19, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3461614.

Gumaei:2022:EAR

[365] Abdu Gumaei, Mabrook S. Al-Rakhami, Mohammad Mehedi Hassan, Victor Hugo C. De Albuquerque, and David Camacho. An effective approach for rumor detection of Arabic tweets using eXtreme gradient boosting method. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):7:1–7:16, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3461697.

Dhall:2022:BBF

[366] Sakshi Dhall, Ashutosh Dhar Dwivedi, Saibal K. Pal, and Gautam Srivastava. Blockchain-based framework for reducing fake or vicious news spread on social media/messaging platforms. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (1):8:1–8:33, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3467019.

De:2022:TBA

[367] Arkadipta De, Dibyanayan Bandyopadhyay, Baban Gain, and Asif Ekbal. A transformer-based approach to multilingual fake news detection in lowresource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):9:1–9:20, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3472619.

Samadi:2022:PFN

[368] Mohammadreza Samadi, Maryam Mousavian, and Saeedeh Momtazi. Persian fake news detection: Neural representation and classification at word and text levels. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):10:1–10:11, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3472620.

Najadat:2022:DAS

[369] Hassan Najadat, Mohammad A. Alzubaidi, and Islam Qarqaz. Detecting Arabic spam reviews in social networks based on classification algorithms. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):11:1–11:13, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3476115.

Jahanbakhsh-Nagadeh:2022:DCB

[370] Zoleikha Jahanbakhsh-Nagadeh, Mohammad-Reza Feizi-Derakhshi, and Arash Sharifi. A deep content-based model for Persian rumor verification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):12:1–12:29, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3487289.

Alam:2022:RUP

[371] Mehreen Alam and Sibt Ul Hussain. Roman–Urdu–Parl: Roman–Urdu and Urdu parallel corpus for Urdu language understanding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):13:1–13:20, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3464424.

Nassif:2022:EES

[372] Ali Bou Nassif, Abdollah Masoud Darya, and Ashraf Elnagar. Empirical evaluation of shallow and deep learning classifiers for Arabic sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):14:1–14:25, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3466171.

Phukon:2022:SEU

[373] Bornali Phukon, Akash Anil, Sanasam Ranbir Singh, and Priyankoo Sarmah. Synonymy expansion using link prediction methods: a case study of Assamese WordNet. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):15:1–15:21, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3467966.

Eddine:2022:NCE

[374] Meftah Mohammed Charaf Eddine. A new concept of electronic text based on semantic coding system for machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):16:1–16:16, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3469655.

Xiang:2022:EGN

[375] Yan Xiang, Zhengtao Yu, Junjun Guo, Yuxin Huang, and Yantuan Xian. Event graph neural network for opinion target classification of microblog comments. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):17:1–17:13, January 2022. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3469725.

Turan:2022:CIA

[376] Erhan Turan and Umut Orhan. Confidence indexing of automated detected synsets: a case study on contemporary Turkish dictionary. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):18:1–18:19, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3469724.

Baruah:2022:LRN

[377] Rupjyoti Baruah, Rajesh Kumar Mundotiya, and Anil Kumar Singh. Low resource neural machine translation: Assamese to/from other Indo– Aryan (Indic) languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):19:1–19:32, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3469721.

Fadel:2022:NAT

[378] Ali Fadel, Ibraheem Tuffaha, and Mahmoud Al-Ayyoub. Neural Arabic text diacritization: State-of-the-art results and a novel approach for Arabic NLP downstream tasks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):20:1–20:25, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3470849.

Kulkarni:2022:SAH

[379] Dhanashree S. Kulkarni and Sunil S. Rodd. Sentiment analysis in Hindi — a survey on the state-of-the-art techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(1):21:1–21:46, January 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3469722.

Yu:2022:ICV

[380] Zhiqiang Yu, Zhengtao Yu, Yantuan Xian, Yuxin Huang, and Junjun Guo. Improving Chinese–Vietnamese neural machine translation with linguistic differences. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):22:1–22:12, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3477536.

Halabi:2022:INA

[381] Dana Halabi, Ebaa Fayyoumi, and Arafat Awajan. I3rab: a new Arabic dependency treebank based on Arabic grammatical theory. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):23:1–23:32, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3472295.

Yang:2022:ASN

[382] Haitong Yang, Guangyou Zhou, and Tingting He. Adversarial separation network for text style transfer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):24:1– 24:14, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3472621.

Fatima:2022:DCL

[383] Ghazeefa Fatima, Rao Muhammad Adeel Nawab, Muhammad Salman Khan, and Ali Saeed. Developing a cross-lingual semantic word similarity corpus for English–Urdu language pair. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):25:1– 25:16, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3472618.

Muneer:2022:CLT

[384] Iqra Muneer and Rao Muhammad Adeel Nawab. Cross-lingual text reuse detection using translation plus monolingual analysis for English-Urdu language pair. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):26:1–26:18, March 2022. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3473331.

Xu:2022:LRL

[385] Fan Xu, Yangjie Dan, Keyu Yan, Yong Ma, and Mingwen Wang. Lowresource language discrimination toward Chinese dialects with transfer learning and data augmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):27:1–27:21, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3473499.

Rana:2022:UAS

[386] Toqir A. Rana, Kiran Shahzadi, Tauseef Rana, Ahsan Arshad, and Mohammad Tubishat. An unsupervised approach for sentiment analysis on social media short text classification in Roman Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):28:1–28:16, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3474119.

Mukherjee:2022:URL

[387] Jayati Mukherjee, Swapan K. Parui, and Utpal Roy. An unsupervised and robust line and word segmentation method for handwritten and degraded printed document. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):29:1–29:31, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3474118.

Mukta:2022:CGB

[388] Md. Saddam Hossain Mukta, Md. Adnanul Islam, Faisal Ahamed Khan, Afjal Hossain, Shuvanon Razik, Shazzad Hossain, and Jalal Mahmud. A comprehensive guideline for Bengali sentiment annotation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):30:1–30:19, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3474363.

Hu:2022:GFQ

[389] Yue Hu, Haitong Yang, Guangyou Zhou, and Jimmy Xiangji Huang. Generating factoid questions with question type enhanced representation and attention-based copy mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):31:1– 31:18, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3474555.

Singh:2022:CSH

[390] Pawan Kumar Singh, Ram Sarkar, Ajith Abraham, and Mita Nasipuri. A case study on handwritten Indic script classification: Benchmarking of the results at page, block, text-line, and word levels. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):32:1–32:36, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3476102.

Gu:2022:MTF

[391] Xiaoqing Gu, Kaijian Xia, Yizhang Jiang, and Alireza Jolfaei. Multi-task fuzzy clustering-based multi-task TSK fuzzy system for text sentiment classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):33:1–33:24, March 2022. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3476103.

Sarwar:2022:UWU

[392] Raheem Sarwar and Saeed-Ul Hassan. UrduAI: Writeprints for Urdu authorship identification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):34:1–34:18, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3476467.

Fatima:2022:SAW

[393] Tayyaba Fatima, Raees Ul Islam, Muhammad Waqas Anwar, M. Hasan Jamal, M. Tayyab Chaudhry, and Zeeshan Gillani. STEMUR: an automated word conflation algorithm for the Urdu language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):35:1–35:20, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3476226.

Hiraoka:2022:RNH

[394] Tatsuya Hiraoka, Sho Takase, Kei Uchiumi, Atsushi Keyaki, and Naoaki Okazaki. Recurrent neural hidden Markov model for high-order transition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):36:1–36:15, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3476511.

S:2022:IWS

[395] Sruthi S., B. Kannan, and Binu Paul. Improved word sense determination in Malayalam using latent Dirichlet allocation and semantic features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):37:1–37:11, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3476978.

Saeed:2022:IFD

[396] Ali Saeed, Rao Muhammad Adeel Nawab, and Mark Stevenson. Investigating the feasibility of deep learning methods for Urdu word sense disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):38:1–38:16, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3477578.

Katyayan:2022:DAR

[397] Pragya Katyayan and Nisheeth Joshi. Development of automatic rulebased semantic tagger and Karaka analyzer for Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):39:1–39:25, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3479155.

Abbad:2022:SED

[398] Hamza Abbad and Shengwu Xiong. Simple extensible deep learning model for automatic Arabic diacritization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2): 40:1–40:16, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3480938.

Huang:2022:DAW

[399] Kaiyu Huang, Keli Xiao, Fengran Mo, Bo Jin, Zhuang Liu, and Degen Huang. Domain-aware word segmentation for Chinese language: a document-level context-aware model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):41:1– 41:16, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3481298.

Bai:2022:UPT

[400] Guirong Bai, Shizhu He, Kang Liu, and Jun Zhao. Using pre-trained language model to enhance active learning for sentence matching. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(2):42:1–42:19, March 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3480937.

Al-Shatnawi:2022:AHW

[401] Atallah Mahmoud Al-Shatnawi, Faisal Al-Saqqar, and Alireza Souri. Arabic handwritten word recognition based on stationary wavelet transform technique using machine learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3): 43:1–43:21, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3474391.

Nazir:2022:AAR

[402] Zulqarnain Nazir, Khurram Shahzad, Muhammad Kamran Malik, Waheed Anwar, Imran Sarwar Bajwa, and Khawar Mehmood. Authorship attribution for a resource poor language — Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (3):44:1–44:23, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3487061.

Sheikhaei:2022:JTL

[403] Mohammad Sadegh Sheikhaei, Hasan Zafari, and Yuan Tian. Joined type length encoding for nested named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):45:1–45:23, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3487057.

Li:2022:DVC

[404] Mei Li, Jiajun Zhang, Xiang Lu, and Chengqing Zong. Dual-view conditional variational auto-encoder for emotional dialogue generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):46:1–46:18, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3481890.

Raval:2022:IDL

[405] Deepang Raval, Vyom Pathak, Muktan Patel, and Brijesh Bhatt. Improving deep learning based automatic speech recognition for Gujarati. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):47:1–47:18, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3483446.

Jiang:2022:TTD

[406] Shu Jiang, Zuchao Li, Hai Zhao, Bao-Liang Lu, and Rui Wang. Tritraining for dependency parsing domain adaptation. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(3):48:1–48:17, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3488367.

Mishra:2022:ECA

[407] Santosh Kumar Mishra, Gaurav Rai, Sriparna Saha, and Pushpak Bhattacharyya. Efficient channel attention based encoder-decoder approach for image captioning in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):49:1–49:17, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3483597.

Liao:2022:SLM

[408] Xianwen Liao, Yongzhong Huang, Peng Yang, and Lei Chen. A statistical language model for pre-trained sequence labeling: a case study on Vietnamese. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):50:1–50:21, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3483524.

Wang:2022:TCC

[409] Zhongguo Wang and Bao Zhang. Toxic comment classification based on bidirectional gated recurrent unit and convolutional neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):51:1–51:12, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3488366.

Sun:2022:OSR

[410] Jian Sun, Yu Zhou, and Chengqing Zong. One-shot relation learning for knowledge graphs via neighborhood aggregation and paths encoding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):52:1–52:19, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3484729.

Munir:2022:MAI

[411] Kashif Munir, Hongxiao Bai, Hai Zhao, and Junhan Zhao. Memorizing all for implicit discourse relation recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):53:1–53:20, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3485016.

Reddy:2022:FBA

[412] A. Pramod Reddy and Vijayarajan V. Fusion based AER system using deep learning approach for amplitude and frequency analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):54:1–54:19, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3488369.

Zhang:2022:LJE

[413] Hu Zhang, Bangze Pan, and Ru Li. Legal judgment elements extraction approach with law article-aware mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (3):55:1–55:15, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3485244.

Rani:2022:ABS

[414] Sujata Rani and Parteek Kumar. Aspect-based sentiment analysis using dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):56:1–56:19, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3485243.

Ishraq:2022:TDU

[415] Mir Ragib Ishraq, Nitesh Khadka, Asif Mohammed Samir, and M. Shahidur Rahman. Towards developing uniform lexicon based sorting algorithm for three prominent Indo–Aryan languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):57:1–57:20, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3488371.

Pandey:2022:HAS

[416] Shilpa Pandey and Gaurav Harit. Handwritten annotation spotting in printed documents using top-down visual saliency models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):58:1–58:25, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3485468.

Kang:2022:ELT

[417] Xiaomian Kang, Yang Zhao, Jiajun Zhang, and Chengqing Zong. Enhancing lexical translation consistency for document-level neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):59:1–59:21, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3485469.

Azmi:2022:LDR

[418] Aqil M. Azmi, Rehab M. Alnefaie, and Hatim A. Aboalsamh. Light diacritic restoration to disambiguate homographs in modern Arabic texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):60:1–60:14, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3486675.

Guo:2022:ACD

[419] Aibo Guo, Xinyi Li, Ning Pang, and Xiang Zhao. Adversarial crossdomain community question retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):61:1– 61:22, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3487291.

Wang:2022:CSS

[420] Shaolei Wang, Zhongyuan Wang, Wanxiang Che, Sendong Zhao, and Ting Liu. Combining self-supervised learning and active learning for disfluency detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):62:1–62:25, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3487290.

Warjri:2022:PSP

[421] Sunita Warjri, Partha Pakray, Saralin A. Lyngdoh, and Arnab Kumar Maji. Part-of-speech (POS) tagging using deep learning-based approaches on the designed Khasi POS corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(3):63:1–63:24, May 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3488381.

Hastuti:2022:QLS

[422] Rochana Prih Hastuti, Yohanes Suyanto, and Anny Kartika Sari. Qlearning for shift-reduce parsing in Indonesian tree-LSTM-based text generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):64:1–64:15, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3490501.

Lee:2022:CEB

[423] Lung-Hao Lee, Jian-Hong Li, and Liang-Chih Yu. Chinese EmoBank: Building valence-arousal resources for dimensional sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):65:1–65:18, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3489141.

Chen:2022:DDG

[424] Shanxiong Chen, Ye Yang, Xuxin Liu, and Shiyu Zhu. Dual discriminator GAN: Restoring ancient yi characters. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):66:1– 66:23, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3490031.

Jana:2022:HDL

[425] Abhik Jana, Gopalakrishnan Venkatesh, Seid Muhie Yimam, and Chris Biemann. Hypernymy detection for low-resource languages: a study for Hindi, Bengali, and Amharic. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):67:1–67:21, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3490389.

Mao:2022:LDM

[426] Zhuoyuan Mao, Chenhui Chu, and Sadao Kurohashi. Linguistically driven multi-task pre-training for low-resource neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):68:1–68:29, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3491065.

Abderrahim:2022:AWS

[427] Mohammed Alaeddine Abderrahim and Mohammed El-Amine Abderrahim. Arabic word sense disambiguation for information retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):69:1–69:19, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3510451.

Ma:2022:ERC

[428] Hongchao Ma, Zhongqing Wang, Xiabing Zhou, Guodong Zhou, and Qinglei Zhou. Emotion recognition with conversational generation transfer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):70:1–70:17, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3494532.

Wu:2022:CEE

[429] Xiaohua Wu, Tengrui Wang, Youping Fan, and Fangjian Yu. Chinese event extraction via graph attention network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):71:1–71:12, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3494533.

Cui:2022:IGD

[430] Yiming Cui, Wanxiang Che, Ziqing Yang, Ting Liu, Bing Qin, Shijin Wang, and Guoping Hu. Interactive gated decoder for machine reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):72:1–72:19, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3501399.

Husain:2022:IEP

[431] Fatemah Husain and Ozlem Uzuner. Investigating the effect of preprocessing Arabic text on offensive language and hate speech detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):73:1–73:20, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3501398.

Gogoi:2022:LLR

[432] Arjun Gogoi and Nomi Baruah. A lemmatizer for low-resource languages: WSD and its role in the Assamese language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4): 74:1–74:22, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3502157.

Harrag:2022:AFN

[433] Fouzi Harrag and Mohamed Khalil Djahli. Arabic fake news detection: a fact checking based deep learning approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4): 75:1–75:34, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3501401.

P:2022:TSS

[434] Jasir M. P. and Kannan Balakrishnan. Text-to-speech synthesis: Literature review with an emphasis on Malayalam language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):76:1–76:56, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3501397.

Qin:2022:MDS

[435] Libo Qin, Fuxuan Wei, Minheng Ni, Yue Zhang, Wanxiang Che, Yangming Li, and Ting Liu. Multi-domain spoken language understanding using domain- and task-aware parameterization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (4):77:1–77:17, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3502198.

Qin:2022:ACE

[436] Yanxia Qin, Zhongqing Wang, Yue Zhang, Kehai Chen, and Min Zhang. Advancing Chinese event detection via revisiting character information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):78:1–78:9, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3502197.

Jain:2022:WSD

[437] Goonjan Jain and D. K. Lobiyal. Word sense disambiguation using cooperative game theory and fuzzy Hindi WordNet based on ConceptNet. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):79:1–79:25, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3502739.

Manerkar:2022:KWC

[438] Sanjana Manerkar, Kavita Asnani, Preeti Ravindranath Khorjuvenkar, Shilpa Desai, and Jyoti D. Pawar. Konkani WordNet: Corpus-based enhancement using crowdsourcing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):80:1– 80:18, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3503156.

Chen:2022:MMR

[439] Junyi Chen, Lan Du, Ming Liu, and Xiabing Zhou. Mulan: a multiple residual article-wise attention network for legal judgment prediction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):81:1–81:15, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3503157.

Guo:2022:HNT

[440] H. Guo, N. Dong, J. Y. Zhao, and Y. F. Liu. Handwritten new Tai Lue character recognition using convolutional prior features and deep variationally sparse Gaussian process modeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21 (4):82:1–82:25, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3506700.

Mahajan:2022:WLS

[441] Shilpa Mahajan and Rajneesh Rani. Word level script identification using convolutional neural network enhancement for scenic images. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):83:1–83:29, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3506699.

Alshammari:2022:CNS

[442] Nasser O. Alshammari and Fawaz D. Alharbi. Combining a novel scoring approach with Arabic stemming techniques for Arabic chatbots conversation engine. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(4):84:1–84:21, July 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3511215.

Shang:2022:IHD

[443] Rui Shang and Xia Li. Improved heuristic data management and protection algorithm for digital China cultural datasets. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(5):85:1–85:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3394114.

Canhasi:2022:AFN

[444] Ercan Canhasi, Rexhep Shijaku, and Erblin Berisha. Albanian fake news detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):86:1–86:??, September 2022. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3487288.

Ahmed:2022:FCS

[445] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Fuzzy contrast set based deep attention network for lexical analysis and mental health treatment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):87:1–87:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3506701.

Fadte:2022:AAV

[446] Swapnil Fadte, Edna Vaz Fernandes, Ramdas Karmali, and Jyoti D. Pawar. Acoustic analysis of vowels in Konkani. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(5):88:1–88:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3474358.

Otoom:2022:NSI

[447] Mwaffaq Otoom, Mohammad A. Alzubaidi, and Lina Nasr Abd Al-Raziq. A novel social interaction assistive device for Arab deaf people. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):89:1–89:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3508374.

Pathak:2022:RAI

[448] Dhrubajyoti Pathak, Sukumar Nandi, and Priyankoo Sarmah. Reduplication in Assamese: Identification and modeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(5):90:1–90:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3510419.

Su:2022:INM

[449] Chao Su, Heyan Huang, Shumin Shi, and Ping Jian. Improving neural machine translation by transferring knowledge from syntactic constituent alignment learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):91:1–91:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3510580.

Mukhtar:2022:IUA

[450] Neelam Mukhtar, Mohammad Abid Khan, Nadia Chiragh, Shah Nazir, and Asim Ullah Jan. An intelligent unsupervised approach for handling context-dependent words in Urdu sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):92:1–92:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3510830.

Nasim:2022:ALC

[451] Zarmeen Nasim and Sajjad Haider. Automatic labeling of clusters for a low-resource Urdu language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):93:1–93:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511097.

Zhang:2022:EMC

[452] Songming Zhang, Ying Zhang, Yufeng Chen, Du Wu, Jinan Xu, and Jian Liu. Exploiting morpheme and cross-lingual knowledge to enhance Mongolian named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):94:1– 94:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3511098.

Li:2022:CBJ

[453] Yan Li, Xiaomin Li, Yiru Wang, Hui Lv, Fenfang Li, and La Duo. Character-based joint word segmentation and part-of-speech tagging for Tibetan based on deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):95:1–95:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511600.

Hu:2022:CAI

[454] Gangqiang Hu, Shengfei Lyu, Xingyu Wu, Jinlong Li, and Huanhuan Chen. Contextual-aware information extractor with adaptive objective for Chinese medical dialogues. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):96:1–96:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511602.

Rafi-Ur-Rashid:2022:BCC

[455] Md. Rafi-Ur-Rashid, Mahim Mahbub, and Muhammad Abdullah Adnan. Breaking the curse of class imbalance: Bangla text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):97:1–97:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3511601.

Chaudhury:2022:DOD

[456] Ayan Chaudhury, Partha Sarathi Mukherjee, Sudip Das, Chandan Biswas, and Ujjwal Bhattacharya. A deep OCR for degraded Bangla documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):98:1–98:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3511807.

Ahmadi:2022:LMN

[457] Sina Ahmadi, Hossein Hassani, and Daban Q. Jaff. Leveraging multilingual news Websites for building a Kurdish parallel corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):99:1–99:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3511806.

Kaur:2022:IFE

[458] Kamaldeep Kaur and Parminder Singh. Impact of feature extraction and feature selection algorithms on Punjabi speech emotion recognition using convolutional neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):100:1– 100:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3511888.

Kumar:2022:DBD

[459] Krishan Kumar. DEAF-BSL: Deep lEArning Framework for British Sign Language recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):101:1– 101:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3513004.

Verma:2022:CAE

[460] Vikas Verma and S. K. Sharma. Critical analysis of existing Punjabi grammar checker and a proposed hybrid framework involving machine learning and rule-base criteria. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):102:1– 102:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3514237.

Shi:2022:LRN

[461] Shumin Shi, Xing Wu, Rihai Su, and Heyan Huang. Low-resource neural machine translation: Methods and trends. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):103:1–103:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3524300.

Mamta:2022:EML

[462] Mamta, Asif Ekbal, and Pushpak Bhattacharyya. Exploring multilingual, multi-task, and adversarial learning for low-resource sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):104:1–104:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3514498.

Nguyen:2022:NVC

[463] Kiet Van Nguyen, Tin Van Huynh, Duc-Vu Nguyen, Anh Gia-Tuan Nguyen, and Ngan Luu-Thuy Nguyen. New Vietnamese corpus for machine reading comprehension of health news articles. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):105:1–105:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3527631.

Mehra:2022:BMB

[464] Pramod Mehra and Shashi Kant Verma. BERIS: an mBERT-based emotion recognition algorithm from Indian speech. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(5):106:1–106:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3517195.

Kumar:2022:NTM

[465] Amit Kumar, Nazanin Esmaili, and Massimo Piccardi. Neural topic model training with the REBAR gradient estimator. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(5):107:1–107:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3517336.

Ren:2022:DUB

[466] Feiliang Ren, Yongkang Liu, Bochao Li, Zhibo Wang, Yu Guo, Shilei Liu, Huimin Wu, Jiaqi Wang, Chunchao Liu, and Bingchao Wang. Deep understanding based multi-document machine reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):108:1–108:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3519296.

Maity:2022:TSA

[467] Shuverthi Maity and Kamal Sarkar. Topic sentiment analysis for Twitter data in Indian languages using composite kernel SVM and deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(5):109:1–109:??, September 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3519297.

Yinying:2022:DMT

[468] Cai Yinying, Juan Li, and Bo Wang. Data mining techniques and machine learning algorithms in the multimedia system to enhance engineering education. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):112:1–112:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3517805.

Xu:2022:EMA

[469] Bo Xu and Jing Jiang. Exploitation for multimedia Asian information processing and artificial intelligence-based art design and teaching in colleges. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):114:1–114:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3526219.

Li:2022:CMC

[470] Zhiqin Li. Construction of marketing curriculum system based on blending learning “3+2” joint training of higher vocational and undergraduate education using NLP for marketing document management and information retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):115:1–115:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3524113.

Li:2022:RIA

[471] Changzhen Li and Xin Zhao. Research on the influence of artificial intelligence technology with Web 3.0 on accounting education and its countermeasures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):116:1–116:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3527666.

Chen:2022:ESI

[472] Zhaojun Chen, Yanhai Bao, and Tongxun Zhu. An empirical study on IPO model construction of undergraduate education quality evaluation in China from the statistical pattern recognition approach in NLP. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):117:1–117:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3543851.

Ding:2022:RTR

[473] Yanhong Ding and Na Yang. Research on teaching reform of innovative and entrepreneurial talents training course in local undergraduate universities based on intelligent blended learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 21(6):119:1–119:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3529393.

Jiang:2022:RBL

[474] Wenxia Jiang, Qingna Lin, and Wei Qin. Research on blended learning evaluation method of overseas Chinese education in the post-epidemic era. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):120:1–120:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3542926.

Zhong:2022:CSS

[475] Deren Zhong. A case study for sports characteristic town: Enlightenment to the sport industry. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):121:1– 121:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3517915.

Jiang:2022:SCE

[476] Guolei Jiang. A study on the construction of the evaluation system of the teaching ability of students using pattern recognition for studying majoring in badminton in the mixed learning model of physical education majors and self-learning system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):123:1– 123:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3527607.

Li:2022:MAC

[477] Yingtao Li. Modeling and analysis of Chinese culture and communication. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):124:1–124:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3514238.

Zhao:2022:FSP

[478] Yun Zhao, Dexi Liu, Changxuan Wan, Xiping Liu, Xiangqing Qiu, and Jianyun Nie. Find supports for the post about mental issues: More than semantic matching. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):126:1–126:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3508373.

Liu:2022:ANE

[479] Bo Liu, Zhuo Su, and Guangzhi Qu. AIP: a named entity recognition method combining glyphs and sounds. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):127:1– 127:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3522736.

Dey:2022:OIS

[480] Spandan Dey, Md Sahidullah, and Goutam Saha. An overview of Indian spoken language recognition from machine learning perspective. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):128:1–128:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3523179.

Niraula:2022:LTE

[481] Nobal B. Niraula, Saurab Dulal, and Diwa Koirala. Linguistic taboos and euphemisms in Nepali. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):129:1– 129:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3524111.

Li:2022:DSC

[482] Zuchao Li, Hai Zhao, Junru Zhou, Kevin Parnow, and Shexia He. Dependency and span, cross-style semantic role labeling on PropBank and NomBank. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):130:1–130:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3526214.

Chen:2022:OTC

[483] Linqing Chen, Junhui Li, Zhengxian Gong, Min Zhang, and Guodong Zhou. One type context is not enough: Global context-aware neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):131:1–131:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3526215.

Guan:2022:RGS

[484] Mengyu Guan, Zhongqing Wang, and Guodong Zhou. Response generation via structure-aware constraints. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):132:1– 132:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3526216.

Wan:2022:CCC

[485] Qizhi Wan, Changxuan Wan, Keli Xiao, Dexi Liu, Qing Liu, Jiangling Deng, Wenkang Luo, and Rong Hu. Construction of a Chinese corpus for multi-type economic event relation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):133:1– 133:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3527240.

Li:2022:SBD

[486] Fenfang Li, Hui Lv, Duo La, Binbin Yong, and Qingguo Zhou. Sentence boundary disambiguation for Tibetan based on attention mechanism at the syllable level. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):134:1–134:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3527663.

Chen:2022:CFO

[487] Qi Chen, Oi Yee Kwong, Yinqiao Li, Tong Xiao, and Jingbo Zhu. Coarseto-fine output predictions for efficient decoding in neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 21(6):135:1–135:??, November 2022. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3527664.

Shijun:2023:RUT

[488] Xu Shijun, Yi Li, Zeng Cong, and Jin Yong Tang. Research on UAV teaching application and technological innovation with 5G technology and development of high-pole throwing hydrangea. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(1):14:1–14:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3529391.

Wang:2023:UWS

[489] Lihao Wang and Xiaoqing Zheng. Unsupervised word segmentation with bi-directional neural language model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):17:1– 17:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3529387.

Marreddy:2023:RPL

[490] Mounika Marreddy, Subba Reddy Oota, Lakshmi Sireesha Vakada, Venkata Charan Chinni, and Radhika Mamidi. Am I a resource-poor language? Data sets, embeddings, models and analysis for four different NLP tasks in Telugu language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):18:1– 18:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3531535.

Hsueh:2023:TOC

[491] Yu-Ling Hsueh and Tai-Liang Chou. A task-oriented chatbot based on LSTM and reinforcement learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):19:1– 19:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3529649.

Cho:2023:TIP

[492] Won Ik Cho and Nam Soo Kim. Text implicates prosodic ambiguity: a corpus for intention identification of the Korean spoken language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):21:1–21:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3529648.

Vo:2023:ITM

[493] Tham Vo. An integrated topic modelling and graph neural network for improving cross-lingual text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1): 22:1–22:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3530260.

Huang:2023:TRA

[494] Zeyu Huang, Wenge Rong, Xiaofeng Zhang, Yuanxin Ouyang, Chenghua Lin, and Zhang Xiong. Token relation aware Chinese named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):24:1–24:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3531534.

Rajalakshmi:2023:SDI

[495] E. Rajalakshmi, R. Elakkiya, Alexey L. Prikhodko, M. G. Grif, Maxim A. Bakaev, Jatinderkumar R. Saini, Ketan Kotecha, and V. Subramaniyaswamy. Static and dynamic isolated Indian and Russian sign language recognition with spatial and temporal feature detection using hybrid neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):26:1–26:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3530989.

Al-Barhamtoshy:2023:AMR

[496] Hassanin M. Al-Barhamtoshy, Kamal M. Jambi, Mohsen A. Rashwan, and Sherif M. Abdou. An Arabic manuscript regions detection, recognition and its applications for OCRing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):27:1– 27:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3532609.

Li:2023:EJC

[497] Zezhong Li, Fuji Ren, Xiao Sun, Degen Huang, and Piao Shi. Exploiting Japanese-Chinese cognates with shared private representations for NMT. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):28:1–28:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3533429.

Chen:2023:RBF

[498] Jiusheng Chen, Xingkai Xu, and Xiaoyu Zhang. Radial basis function attention for named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):30:1– 30:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3539014.

Elghannam:2023:MLA

[499] Fatma Elghannam. Multi-label annotation and classification of Arabic texts based on extracted seed keyphrases and bi-gram alphabet feed forward neural networks model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(1):31:1–31:??, January 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3539607.

Priyadarshi:2023:SPR

[500] Ankur Priyadarshi and Sujan Kumar Saha. A study on the performance of recurrent neural network based models in Maithili part of speech tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):32:1–32:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3540260.

Mishra:2023:OLB

[501] Santosh Kumar Mishra, Harshit, Sriparna Saha, and Pushpak Bhattacharyya. An object localization-based dense image captioning framework in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):33:1–33:??, February 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3558391.

Yang:2023:MTM

[502] Hang Yang, Yubo Chen, Kang Liu, Jun Zhao, Zuyu Zhao, and Weijian Sun. Multi-turn and multi-granularity reader for document-level event extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):34:1–34:??, February 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3542925.

Eiselen:2023:INE

[503] Roald Eiselen and Andiswa Bukula. IsiXhosa named entity recognition resources. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):35:1–35:??, February 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3531478.

Rai:2023:APB

[504] Pooja Rai and Sanjay Chatterji. Annotation projection-based dependency parser development for Nepali. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):36:1– 36:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3542696.

Demir:2023:TDT

[505] Seniz Demir. Turkish data-to-text generation using sequence-to-sequence neural networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):37:1–37:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3543826.

Goyal:2023:FBT

[506] Kapil Dev Goyal, Muhammad Raihan Abbas, Vishal Goyal, and Yasir Saleem. Forward-backward transliteration of Punjabi Gurmukhi script using N-gram language model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):38:1–38:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3542924.

Gomaa:2023:LAA

[507] Walid Gomaa. Lyrics analysis of the Arab singer Abdel ElHalim Hafez. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):39:1–39:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3544100.

Sun:2023:SDM

[508] Qingying Sun, Xuefeng Xi, Jiajun Sun, Zhongqing Wang, and Huiyan Xu. Stance detection with a multi-target adversarial attention network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):40:1–40:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3544490.

Alzubaidi:2023:NAG

[509] Mohammad A. Alzubaidi, Mwaffaq Otoom, and Areen M. Abu Rwaq. A novel assistive glove to convert Arabic sign language into speech. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):41:1–41:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3545113.

Do:2023:TOK

[510] Heejin Do and Gary Geunbae Lee. Target-oriented knowledge distillation with language-family-based grouping for multilingual NMT. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):42:1–42:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3546067.

Ren:2023:UOR

[511] Feiliang Ren, Yongkang Liu, Bochao Li, Shilei Liu, Bingchao Wang, Jiaqi Wang, Chunchao Liu, and Qi Ma. An understanding-oriented robust machine reading comprehension model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2): 43:1–43:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3546190.

Ramdinmawii:2023:PAC

[512] Esther Ramdinmawii and Sanghamitra Nath. A preliminary analysis on the correlates of stress and tones in mizo. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2): 44:1–44:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3546950.

Bai:2023:BSO

[513] Guirong Bai, Shizhu He, Kang Liu, and Jun Zhao. Bidirectional sentence ordering with interactive decoding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):45:1– 45:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3561510.

Roy:2023:MEC

[514] Aniruddha Roy, Isha Sharma, Sudeshna Sarkar, and Pawan Goyal. Meta-ED: Cross-lingual event detection using meta-learning for Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):46:1–46:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3555340.

Lahoti:2023:SNR

[515] Pawan Lahoti, Namita Mittal, and Girdhari Singh. A survey on NLP resources, tools, and techniques for Marathi language processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):47:1–47:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3548457.

Ilyas:2023:EDC

[516] Abdullah Ilyas, Khurram Shahzad, and Muhammad Kamran Malik. Emotion detection in code-mixed Roman Urdu–English text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):48:1–48:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3552515.

Liu:2023:CLB

[517] Maofu Liu, Junyi Xiang, Xu Xia, and Huijun Hu. Contrastive learning between classical and modern Chinese for classical Chinese machine reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):49:1–49:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3551637.

Nag:2023:TLL

[518] Arijit Nag, Bidisha Samanta, Animesh Mukherjee, Niloy Ganguly, and Soumen Chakrabarti. Transfer learning for low-resource multilingual relation classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):50:1–50:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3554734.

DSilva:2023:ISM

[519] Jovi D’Silva and Uzzal Sharma. Impact of similarity measures in graphbased automatic text summarization of Konkani texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):51:1–51:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3554943.

Chiba:2023:AVE

[520] Yuya Chiba and Ryuichiro Higashinaka. Analyzing variations of everyday Japanese conversations based on semantic labels of functional expressions. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):52:1–52:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3552310.

Zhang:2023:GSG

[521] Mengli Zhang, Gang Zhou, Wanting Yu, Ningbo Huang, and Wenfen Liu. GA-SCS: Graph-augmented source code summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):53:1–53:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3554820.

Zhou:2023:SEP

[522] Bo Zhou, Yubo Chen, Kang Liu, and Jun Zhao. Script event prediction via multilingual event graph networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):54:1– 54:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3557893.

Guven:2023:CLM

[523] Zekeriya Anil Guven. The comparison of language models with a novel text filtering approach for Turkish sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):55:1–55:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3557892.

Lei:2023:LPF

[524] Lei Lei, Guoshun Yuan, Tianle Zhang, and Hongjiang Yu. Low-power feature-attention Chinese keyword spotting framework with distillation learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):56:1–56:??, February 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3558002.

Bidgoly:2023:CBS

[525] Amir Jalaly Bidgoly, Hossein Amirkhani, and Razieh Baradaran. Clustering-based sequence to sequence model for generative question answering in a low-resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):57:1– 57:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3563036.

Komiya:2023:CWE

[526] Kanako Komiya, Shinji Kono, Takumi Seito, and Teruo Hirabayashi. Composing word embeddings for compound words using linguistic knowledge. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):58:1–58:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3561299.

Nambiar:2023:AST

[527] Sindhya K. Nambiar, David Peter S., and Sumam Mary Idicula. Abstractive summarization of text document in Malayalam language: Enhancing attention model using POS tagging feature. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2): 59:1–59:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3561819.

P:2023:IEF

[528] Jasir M. P. and Kannan Balakrishnan. Identification and extraction of features from Malayalam poems for analyzing syllable duration patterns. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2):60:1–60:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3561298.

FHA:2023:DEM

[529] Shibly FHA, Uzzal Sharma, and HMM. Naleer. Development of an efficient method to detect mixed social media data with Tamil–English code using machine learning techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(2): 61:1–61:??, February 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3563775.

Lin:2023:ISI

[530] Jerry Chun-Wei Lin, Vicente García Díaz, and Juan Antonio Morente Molinera. Introduction to the special issue of recent advances in computational linguistics for Asian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):62:1– 62:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3588316.

Liu:2023:ELE

[531] Shuai Liu, Tenghui He, Jingyi Li, Yating Li, and Akshi Kumar. An effective learning evaluation method based on text data with real-time attribution — a case study for mathematical class with students of junior middle school in China. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):63:1–63:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3474367.

Zhu:2023:UPS

[532] Shaolin Zhu, Chenggang Mi, Tianqi Li, Yong Yang, and Chun Xu. Unsupervised parallel sentences of machine translation for Asian language pairs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):64:1–64:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3486677.

Fang:2023:ECE

[533] Hui Fang, Hongmei Shi, Jiuzhou Zhang, and Marimuthu Karuppiah. Effective college English teaching based on teacher–student interactive model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):65:1–65:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3486676.

Sonthi:2023:ITH

[534] Vijaya Krishna Sonthi, Nagarajan S., and Krishnaraj N. An intelligent Telugu handwritten character recognition using multi-objective mayfly optimization with deep learning-based DenseNet model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):67:1–67:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3520439.

Li:2023:TED

[535] Shaobo Li, Chengjie Sun, Zhen Xu, Prayag Tiwari, Bingquan Liu, Deepak Gupta, K. Shankar, Zhenzhou Ji, and Mingjiang Wang. Toward explainable dialogue system using two-stage response generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):68:1–68:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3551869.

Chen:2023:RIA

[536] Keliang Chen, Jianming Huang, Qi Zhang, and Yansong Cui. Research and implementation of automatic indexing method of PDF for digital publishing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):69:1–69:??, March 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3501400.

Shoaib:2023:CAU

[537] Umar Shoaib, Laiba Fiaz, Chinmay Chakraborty, and Hafiz Tayyab Rauf. Context-aware Urdu information retrieval system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):70:1–70:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3502854.

Chen:2023:RCA

[538] Keliang Chen, Jianming Huang, Yansong Cui, and Weizheng Ren. Research on Chinese audio and text alignment algorithm based on AIC-FCM and Doc2Vec. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):71:1–71:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3532852.

Nosrati:2023:WRE

[539] Vahid Nosrati, Mohsen Rahmani, Alireza Jolfaei, and Sattar Seifollahi. A weak-region enhanced Bayesian classification for spam content-based filtering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):72:1–72:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3510420.

Xue:2023:IHO

[540] Xingsi Xue and Wenyu Liu. Integrating heterogeneous ontologies in Asian languages through compact genetic algorithm with annealing resample inheritance mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):73:1–73:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3519298.

Feng:2023:DLC

[541] Hailin Feng, Shuxuan Xie, Wei Wei, Haibin Lv, and Zhihan Lv. Deep learning in computational linguistics for Chinese language translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):74:1–74:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3519386.

Prabadevi:2023:DMR

[542] B. Prabadevi, N. Deepa, K. Ganesan, and Gautam Srivastava. A decision model for ranking Asian higher education institutes using an NLPbased text analysis approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):75:1–75:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3534562.

Zhang:2023:DIF

[543] Tao Zhang, Congying Xia, Zhiwei Liu, Shu Zhao, Hao Peng, and Philip Yu. Domain-invariant feature progressive distillation with adversarial adaptive augmentation for low-resource cross-domain NER. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):76:1–76:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3570502.

Liu:2023:TBU

[544] Yuzhi Liu and Massimo Piccardi. Topic-based unsupervised and supervised dictionary induction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):77:1–77:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3564698.

Li:2023:KSL

[545] Chao Li, Xin Yan, Guangyi Xu, Zhongying Deng, and Yuanyuan Mo. Khmer sentiment lexicon based on PU learning and label propagation algorithm. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):78:1–78:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3564697.

Ma:2023:EIM

[546] Chunpeng Ma, Aili Shen, Hiyori Yoshikawa, Tomoya Iwakura, Daniel Beck, and Timothy Baldwin. On the effectiveness of images in multimodal text classification: an annotation study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (3):79:1–79:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3565572.

Zi:2023:BBS

[547] Yunfei Zi and Shengwu Xiong. BSML: Bidirectional sampling aggregation-based metric learning for low-resource Uyghur few-shot speaker verification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):80:1–80:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3564782.

Rahman:2023:PPB

[548] Habibur Rahman, Md. Rezwan Shahrior Rahin, Araf Mohammad Mahbub, Md. Adnanul Islam, Md. Saddam Hossain Mukta, and Md. Mahbubur Rahman. Punctuation prediction in Bangla text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):81:1–81:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3575804.

Tran:2023:BCD

[549] Phuoc Tran, Dat Nguyen, Huu-Anh Tran, Thien Nguyen, and Tram Tran. Building a closed-domain question answering system for a lowresource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):82:1–82:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3566123.

Muneer:2023:DLB

[550] Iqra Muneer, Ghazeefa Fatima, Muhammad Salman Khan, Rao Muhammad Adeel Nawab, and Ali Saeed. Developing a large benchmark corpus for Urdu semantic word similarity. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):83:1– 83:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3566124.

Mahlaza:2023:SRA

[551] Zola Mahlaza and C. Maria Keet. Surface realization architecture for low-resourced African languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):84:1–84:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3567594.

Kchaou:2023:HPB

[552] Saméh Kchaou, Rahma Boujelbane, and Lamia Hadrich. Hybrid pipeline for building Arabic Tunisian dialect-standard Arabic neural machine translation model from scratch. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):85:1–85:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3568674.

Kaur:2023:MSP

[553] Navdeep Kaur and Parminder Singh. Modelling of speech parameters of Punjabi by pre-trained deep neural network using stacked denoising autoencoders. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):86:1–86:??, March 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3568308.

S:2023:MNF

[554] Rakesh Kumar S., Gayathri Nagasubramanian, Muthuramalingam S., and Fadi Al-Turjman. Multilingual news feed analysis using intelligent linguistic particle filtering techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):87:1– 87:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3569899.

Ma:2023:SCS

[555] Tinghuai Ma, Zheng Zhang, Huan Rong, and Najla Al-Nabhan. SPK-CG: Siamese network based posterior knowledge selection model for knowledge driven conversation generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3): 88:1–88:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3569579.

Wang:2023:CGE

[556] Hongfei Wang, Michiki Kurosawa, Satoru Katsumata, Masato Mita, and Mamoru Komachi. Chinese grammatical error correction using pretrained models and pseudo data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):89:1–89:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3570209.

Su:2023:SBT

[557] Xuefeng Su, Ru Li, Xiaoli Li, Baobao Chang, Zhiwei Hu, Xiaoqi Han, and Zhichao Yan. A span-based target-aware relation model for framesemantic parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):90:1–90:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3569581.

Long:2023:DNN

[558] Kaifang Long, Han Zhao, Zengzhen Shao, Yang Cao, Yanfang Geng, Yintai Sun, Weizhi Xu, and Hui Yu. Deep neural network with embedding fusion for Chinese named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (3):91:1–91:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3570328.

Yirmibesoglu:2023:MMI

[559] Zeynep Yi̇rmi̇besoglu and Tunga Güngör. Morphologically motivated input variations and data augmentation in Turkish–English neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(3):92:1–92:??, March 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3571073.

Alqahtani:2023:SAI

[560] Fatimah Alqahtani and Mischa Dohler. Survey of authorship identification tasks on Arabic texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):93:1–93:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3564156.

Attar:2023:SAA

[561] Rakesh Kumar Attar, Vishal Goyal, and Lalit Goyal. State of the art of automation in sign language: a systematic review. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (4):94:1–94:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3564769.

Nassiri:2023:AMR

[562] Naoual Nassiri, Violetta Cavalli-Sforza, and Abdelhak Lakhouaja. Approaches, methods, and resources for assessing the readability of Arabic texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):95:1–95:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3571510.

Shammi:2023:CRB

[563] Shumaiya Akter Shammi, Sajal Das, Narayan Ranjan Chakraborty, Sumit Kumar Banshal, and Nishu Nath. A comprehensive roadmap on Bangla text-based sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):96:1– 96:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3572783.

Naous:2023:ODR

[564] Tarek Naous, Zahraa Bassyouni, Bassel Mousi, Hazem Hajj, Wassim El Hajj, and Khaled Shaban. Open-domain response generation in lowresource settings using self-supervised pre-training of warm-started transformers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):97:1–97:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3579164.

Touma:2023:AGH

[565] Roudy Touma, Hazem Hajj, Wassim El-Hajj, and Khaled Shaban. Automated generation of human-readable natural Arabic text from RDF data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):98:1–98:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3582262.

Chen:2023:FPT

[566] Jie Chen, Zhiqiang Yao, Shu Zhao, and Yanping Zhang. Fusion pretrained emoji feature enhancement for sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):99:1–99:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3578582.

Liu:2023:CAT

[567] Jiangwei Liu, Zian Yan, Sibao Chen, Xiao Sun, and Bin Luo. Channel attention TextCNN with feature word extraction for Chinese sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):100:1–100:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3571716.

Mi:2023:IRL

[568] Chenggang Mi. Improving the robustness of loanword identification in social media texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):101:1–101:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3572773.

Guo:2023:CSS

[569] Dongyue Guo, Jianwei Zhang, Bo Yang, and Yi Lin. A comparative study of speaker role identification in air traffic communication using deep learning approaches. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):102:1–102:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3572792.

Duwairi:2023:MVL

[570] Rehab Duwairi and Zain Halloush. A multi-view learning approach for detecting personality disorders among Arab social media users. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):103:1–103:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3572906.

Xu:2023:DLR

[571] Wang Xu, Kehai Chen, and Tiejun Zhao. Document-level relation extraction with path reasoning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):104:1–104:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3572898.

Di:2023:BDU

[572] Donglin Di, Xianyang Song, Weinan Zhang, Yue Zhang, and Fanglin Wang. Building dialogue understanding models for low-resource language Indonesian from scratch. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):105:1–105:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3575803.

Balakrishnan:2023:TOL

[573] Vimala Balakrishnan, Vithyatheri Govindan, and Kumanan N. Govaichelvan. Tamil offensive language detection: Supervised versus unsupervised learning approaches. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):106:1– 106:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3575860.

Berrimi:2023:AMA

[574] Mohamed Berrimi, Mourad Oussalah, Abdelouahab Moussaoui, and Mohamed Saidi. Attention mechanism architecture for Arabic sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):107:1–107:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3578265.

Ghosal:2023:HUH

[575] Sayani Ghosal and Amita Jain. HateCircle and unsupervised hate speech detection incorporating emotion and contextual semantics. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):108:1–108:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3576913.

Li:2023:LRN

[576] Yinqiao Li, Runzhe Cao, Qiaozhi He, Tong Xiao, and Jingbo Zhu. Learning reliable neural networks with distributed architecture representations. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):109:1–109:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3578709.

Tang:2023:NVG

[577] Yi-Kun Tang, Heyan Huang, Xuewen Shi, and Xian-Ling Mao. Neural variational Gaussian mixture topic model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4): 110:1–110:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3578583.

Yu:2023:BSD

[578] Wei Yu, Haiyan Yang, Mengzhu Wang, and Xiaodong Wang. Bravely say I don’t know: Relational question-schema graph for text-to-SQL answerability classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):111:1–111:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3579030.

Mishra:2023:DCB

[579] Santosh Kumar Mishra, Sushant Sinha, Sriparna Saha, and Pushpak Bhattacharyya. Dynamic convolution-based encoder-decoder framework for image captioning in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):112:1– 112:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3573891.

Wanjawa:2023:KQA

[580] Barack W. Wanjawa, Lilian D. A. Wanzare, Florence Indede, Owen Mconyango, Lawrence Muchemi, and Edward Ombui. KenSwQuAD — a question answering dataset for Swahili low-resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):113:1–113:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3578553.

Coban:2023:DCD

[581] Onder Coban, Selma Ayse Ozel, and Ali Inan. Detection and crossdomain evaluation of cyberbullying in Facebook activity contents for Turkish. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):114:1–114:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3580393.

Akhter:2023:HPO

[582] Shaheera Saba Mohd Naseem Akhter and Priti P. Rege. Hyper parameter optimization of CRNN for printed Devanagari script recognition using Taguchi’s method. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):115:1–115:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3578549.

Toraman:2023:ITL

[583] Cagri Toraman, Eyup Halit Yilmaz, Furkan Sahi̇nuç, and Oguzhan Ozcelik. Impact of tokenization on language models: an analysis for Turkish. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):116:1–116:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3578707.

Sebastian:2023:MNL

[584] Mary Priya Sebastian and Santhosh Kumar G. Malayalam natural language processing: Challenges in building a phrase-based statistical machine translation system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):117:1–117:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3579163.

Wang:2023:MTL

[585] Xuelei Wang, Xirong Xu, Degen Huang, and Ting Zhang. Multi-task label-wise transformer for Chinese named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):118:1–118:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3576025.

Xu:2023:STC

[586] Qianqian Xu, Junjie Peng, Cangzhi Zheng, Shuhua Tan, Fen Yi, and Feng Cheng. Short text classification of Chinese with label information assisting. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):119:1–119:??, April 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3582301.

Jha:2023:FEV

[587] Piyush Jha, Rashi Kumar, and Vineet Sahula. Filtering and extended vocabulary based translation for low-resource language pair of Sanskrit– Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):120:1–120:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3580495.

Hafeez:2023:USP

[588] Hamza Hafeez, Iqra Muneer, Muhammad Sharjeel, Muhammad Adnan Ashraf, and Rao Muhammad Adeel Nawab. Urdu short paraphrase detection at sentence level. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):121:1–121:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3586009.

Wang:2023:LCD

[589] Xiangyu Wang and Chengqing Zong. Learning category distribution for text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(4):122:1–122:??, April 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3585279.

Yan:2023:CLB

[590] Jinghui Yan, Chengqing Zong, and Jinan Xu. Combination of loss-based active learning and semi-supervised learning for recognizing entities in Chinese electronic medical records. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):123:1– 123:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3588314.

Lv:2023:GDA

[591] Qi Lv, Ziqiang Cao, Lei Geng, Chunhui Ai, Xu Yan, and Guohong Fu. General and domain-adaptive Chinese spelling check with errorconsistent pretraining. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):124:1–124:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3564271.

Wang:2023:BLN

[592] Shuheng Wang, Heyan Huang, and Shumin Shi. Better localness for non-autoregressive transformer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):125:1– 125:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3587266.

Shi:2023:ART

[593] Xuewen Shi, Heyan Huang, Ping Jian, and Yi-Kun Tang. Approximating to the real translation quality for neural machine translation via causal motivated methods. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):126:1–126:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3583684.

Biltawi:2023:ASE

[594] Mariam M. Biltawi, Arafat Awajan, and Sara Tedmori. Arabic span extraction-based reading comprehension benchmark (ASER) and neural baseline models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):127:1–127:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3579047.

M:2023:TGT

[595] Diviya M. and Karmel A. TAM GAN: Tamil text to naturalistic image synthesis using conventional deep adversarial networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):128:1–128:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3584019.

Liu:2023:OOS

[596] Yuanxia Liu, Tianyong Hao, Hai Liu, Yuanyuan Mu, Heng Weng, and Fu Lee Wang. OdeBERT: One-stage deep-supervised early-exiting BERT for fast inference in user intent classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (5):129:1–129:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3587464.

Srivastava:2023:ESI

[597] Gautam Srivastava, Jerry Chun-Wei Lin, and Yu-Dong Zhang. Editorial for the special issue on computational linguistics processing in low-resource indigenous languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):130:1– 130:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3591208.

Bashir:2023:CAE

[598] Muhammad Farrukh Bashir, Abdul Rehman Javed, Muhammad Umair Arshad, Thippa Reddy Gadekallu, Waseem Shahzad, and Mirza Omer Beg. Context-aware emotion detection from low-resource Urdu language using deep neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):131:1– 131:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3528576.

Manjari:2023:QQE

[599] Kanak Manjari, Madhushi Verma, Gaurav Singal, and Suyel Namasudra. QEST: Quantized and efficient scene text detector using deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):132:1–132:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3526217.

Kumar:2023:HDL

[600] Akshi Kumar, Saurabh Raj Sangwan, Adarsh Kumar Singh, and Gandharv Wadhwa. Hybrid deep learning model for sarcasm detection in Indian indigenous language using word-emoji embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):133:1–133:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3519299.

Chopra:2023:FOH

[601] Abhishek Chopra, Deepak Kumar Sharma, Aashna Jha, and Uttam Ghosh. A framework for online hate speech detection on code-mixed Hindi-English text and Hindi text in Devanagari. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (5):134:1–134:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3568673.

Wu:2023:JID

[602] Yirui Wu, Hao Li, Lilai Zhang, Chen Dong, Qian Huang, and Shaohua Wan. Joint intent detection model for task-oriented human-computer dialogue system using asynchronous training. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5): 135:1–135:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3558096.

Chakraborty:2023:EOB

[603] Chinmay Chakraborty, Shaohua Wan, and Mohammad R. Khosravi. Editorial: Ontology-based knowledge presentation and computational linguistics for semantic big social data analytics in Asian social networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):136:1–136:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3594719.

Singh:2023:SNA

[604] Shashank Sheshar Singh, Vishal Srivastava, Ajay Kumar, Shailendra Tiwari, Dilbag Singh, and Heung-No Lee. Social network analysis: a survey on measure, structure, language information analysis, privacy, and applications. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):137:1–137:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3539732.

Anwar:2023:SRA

[605] Sibgha Anwar, Mirza Omer Beg, Kiran Saleem, Zeeshan Ahmed, Abdul Rehman Javed, and Usman Tariq. Social relationship analysis using state-of-the-art embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):138:1– 138:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3539608.

Sharma:2023:SML

[606] Neha Sharma, Mukesh Soni, Sumit Kumar, Rajeev Kumar, Nabamita Deb, and Anurag Shrivastava. Supervised machine learning method for ontology-based financial decisions in the stock market. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):139:1–139:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3554733.

Jiang:2023:GVS

[607] Wei Jiang, Mengqi Li, Mohammad Shabaz, Ashutosh Sharma, and Mohd Anul Haq. Generation of voice signal tone sandhi and melody based on convolutional neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5): 140:1–140:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3545569.

Bibi:2023:RCR

[608] Nazia Bibi, Tauseef Rana, Ayesha Maqbool, Tamim Alkhalifah, Wazir Zada Khan, Ali Kashif Bashir, and Yousaf Bin Zikria. Reusable component retrieval: a semantic search approach for low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):141:1–141:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3564604.

Liao:2023:IRA

[609] Junwei Liao, Sefik Eskimez, Liyang Lu, Yu Shi, Ming Gong, Linjun Shou, Hong Qu, and Michael Zeng. Improving readability for automatic speech recognition transcription. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):142:1– 142:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3557894.

Chen:2023:UBH

[610] Qian Chen, Xiao Sun, Jiamin Wang, and Meng Wang. User-based hierarchical network of Sina Weibo emotion analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (5):143:1–143:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3579048.

Pei:2023:SGS

[611] Jiaming Pei, Kaiyang Zhong, Zhi Yu, Lukun Wang, and Kuruva Lakshmanna. Scene graph semantic inference for image and text matching. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):144:1–144:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3563390.

Liang:2023:WDD

[612] Kun Liang, Ruhui Ma, Yang Hua, Hao Wang, Ningxin Hu, Tao Song, Honghao Gao, and Haibing Guan. WH^{2}D^{2}N^{2}: Distributed AI-enabled OK-ASN service for Web of Things. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):145:1– 145:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3564242.

Wu:2023:CMP

[613] Zongda Wu, Jian Xie, Shigen Shen, Chongze Lin, Guandong Xu, and Enhong Chen. A confusion method for the protection of user topic privacy in Chinese keyword-based book retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):

146:1–146:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3571731.

Belhadi:2023:FAF

[614] Asma Belhadi, Youcef Djenouri, Gautam Srivastava, and Jerry Chun-Wei Lin. Fast and accurate framework for ontology matching in Web of Things. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):147:1–147:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3578708.

Berhoum:2023:IAB

[615] Adel Berhoum, Mohammed Charaf Eddine Meftah, Abdelkader Laouid, and Mohammad Hammoudeh. An intelligent approach based on cleaning up of inutile contents for extremism detection and classification in social networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):148:1–148:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3575802.

Deng:2023:CFA

[616] Xiaoheng Deng, Dingjie Han, and Ping Jiang. A context-focused attention evolution model for aspect-based sentiment classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):149:1–149:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3587465.

Kumar:2023:OLD

[617] Sanjay Kumar, Akshi Kumar, Abhishek Mallik, and Sakshi Dhall. Opinion leader detection in Asian social networks using modified spider monkey optimization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):150:1–150:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3555311.

Kumar:2023:PDU

[618] Akshi Kumar, Rohit Beniwal, and Dipika Jain. Personality detection using kernel-based ensemble model for leveraging social psychology in online networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):151:1–151:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3571584.

Zhang:2023:MSF

[619] Yuanpeng Zhang, Yizhang Jiang, and Jolfaei Alireza. Mutual supervised fusion & transfer learning with interpretable linguistic meaning for social data analytics. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):152:1–152:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3568675.

Chauhan:2023:RBF

[620] Shweta Chauhan, Jayashree Premkumar Shet, Shehab Mohamed Beram, Vishal Jagota, Mohammed Dighriri, Mohd Wazih Ahmad, Md Shamim Hossain, and Ali Rizwan. Rule based fuzzy computing approach on selfsupervised sentiment polarity classification with word sense disambiguation in machine translation for Hindi language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (5):153:1–153:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3574130.

Jin:2023:SPT

[621] Dawei Jin, Yiyi Hu, Jingyu Chen, and Mengran Xia. Stock price trends prediction based on the classical models with key information fusion of ontologies. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(5):154:1–154:??, May 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3592599.

Dey:2023:DBO

[622] Raghunath Dey and Rakesh Chandra Balabantaray. Development of a benchmark Odia handwritten character database for an efficient offline handwritten character recognition with a chronological survey. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):155:1–155:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3583988.

Saeed:2023:DOL

[623] Ramsha Saeed, Hammad Afzal, Sadaf Abdul Rauf, and Naima Iltaf. Detection of offensive language and ITS severity for low resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):156:1–156:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3580476.

Huang:2023:CAT

[624] Xin Huang, Jiajun Zhang, and Chengqing Zong. Contrastive adversarial training for multi-modal machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):157:1–157:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3587267.

Jia:2023:TMA

[625] Yue Jia, Wei Fang, and Heng-Yang Lu. Think more ambiguity less: a novel dual interactive model with local and global semantics for Chinese named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):158:1– 158:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3583685.

Huang:2023:KEP

[626] Hu Huang, Bowen Zhang, Yangyang Li, Baoquan Zhang, Yuxi Sun, Chuyao Luo, and Cheng Peng. Knowledge-enhanced prompt-tuning for stance detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):159:1–159:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3588767.

Zhao:2023:BBF

[627] Feng Zhao, Cheng Yan, Hai Jin, and Lifang He. BayesKGR: Bayesian few-shot learning for knowledge graph reasoning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):160:1–160:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3589183.

Das:2023:ITM

[628] Ringki Das and Thoudam Doren Singh. Image-text multimodal sentiment analysis framework of Assamese news articles using late fusion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):161:1–161:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3584861.

Munir:2023:SSS

[629] Kashif Munir, Hai Zhao, and Zuchao Li. Semi-supervised semantic role labeling with bidirectional language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):

162:1–162:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3587160.

Yi:2023:IRP

[630] Nian Yi, Chenze Shao, and Aishan Wumaier. Integrating reconstructor and post-Editor into neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):163:1–163:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3588766.

Parte:2023:EAD

[631] Smita Athanere Parte, Ankur Ratmele, and Ritesh Dhanare. An efficient and accurate detection of fake news using capsule transient auto encoder. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):164:1–164:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3589184.

Sharma:2023:LLF

[632] Richa Sharma and Arti Arya. LFWE: Linguistic feature based word embedding for Hindi fake news detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6): 165:1–165:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3589764.

Thin:2023:VSA

[633] Dang Van Thin, Duong Ngoc Hao, and Ngan Luu-Thuy Nguyen. Vietnamese sentiment analysis: an overview and comparative study of finetuning pretrained language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):166:1– 166:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3589131.

Dalai:2023:PST

[634] Tusarkanta Dalai, Tapas Kumar Mishra, and Pankaj K. Sa. Part-of-speech tagging of Odia language using statistical and deep learning based approaches. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):167:1–167:??, June 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3588900.

Singh:2023:KHN

[635] Niraj Kumar Singh, Komal Naaz, and Soubhik Chakraborty. Komala and Kat.hora: a novel approach towards classification of Hindi poetry. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):168:1–168:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3589249.

Das:2023:IMN

[636] Sudhansu Bala Das, Atharv Biradar, Tapas Kumar Mishra, and Bidyut Kr. Patra. Improving multilingual neural machine translation system for Indic languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):169:1–169:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3587932.

Khanmohammadi:2023:PBT

[637] Reza Khanmohammadi, Mitra Sadat Mirshafiee, Yazdan Rezaee Jouryabi, and Seyed Abolghasem Mirroshandel. Prose2Poem: The blessing of transformers in translating prose to Persian poetry. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):170:1–170:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3592791.

Yang:2023:TTE

[638] Liang Yang, Zhexu Shen, Fengqing Zhou, Hongfei Lin, and Junpeng Li. TPoet: Topic-enhanced Chinese poetry generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):171:1–171:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3593805.

Shamas:2023:MML

[639] Mohsen Shamas, Wassim El Hajj, Hazem Hajj, and Khaled Shaban. Metadial: a meta-learning approach for Arabic dialogue generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):172:1–172:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3590960.

Sharjeel:2023:CLT

[640] Muhammad Sharjeel, Iqra Muneer, Sumaira Nosheen, Rao Muhammad Adeel Nawab, and Paul Rayson. Cross-lingual text reuse detection at document level for English–Urdu language pair. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):173:1–173:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3592761.

Zhang:2023:ERV

[641] Fan Zhang, Meishan Zhang, Shuang Liu, Yueheng Sun, and Nan Duan. Enhancing RDF verbalization with descriptive and relational knowledge. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):174:1–174:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3595293.

Shafi:2023:STU

[642] Jawad Shafi, Rao Muhammad Adeel Nawab, and Paul Rayson. Semantic tagging for the Urdu language: Annotated corpus and multi-target classification methods. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):175:1–175:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3582496.

Huang:2023:CLS

[643] Yuxin Huang, Yin Liang, Zhaoyuan Wu, Enchang Zhu, and Zhengtao Yu. Cross-lingual sentence embedding for low-resource Chinese– Vietnamese based on contrastive learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6): 176:1–176:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3589341.

Liao:2023:TPC

[644] Junwei Liao, Shuai Cheng, and Minghuan Tan. Text polishing with Chinese idiom: Task, datasets and pre-trained baselines. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):177:1–177:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3593806.

Khellas:2023:ARD

[645] Kenza Khellas and Rachid Seghir. Alabib-65: a realistic dataset for Algerian sign language recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):178:1– 178:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3596909.

Khalid:2023:UDA

[646] Hamza Khalid, Ghulam Murtaza, and Qaiser Abbas. Using data augmentation and bidirectional encoder representations from transformers for improving Punjabi named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):179:1–179:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3595861.

Zhao:2023:SNN

[647] Shuai Zhao, Qing Li, Yuer Yang, Jinming Wen, and Weiqi Luo. From softmax to nucleusmax: a novel sparse language model for Chinese radiology report summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):180:1– 180:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3596219.

Abbache:2023:IAD

[648] Mohamed Abbache, Ahmed Abbache, Jingwen Xu, Farid Meziane, and Xianbin Wen. The impact of Arabic diacritization on word embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):181:1–181:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3592603.

Park:2023:RMT

[649] Cheoneum Park and Juae Kim. Robust multi-task learning-based Korean POS tagging to overcome word spacing errors. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (6):182:1–182:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3591206.

Li:2023:MBB

[650] Zezhong Li, Xiao Sun, Fuji Ren, Jianjun Ma, Degen Huang, and Piao Shi. Multilingual BERT-based word alignment by incorporating common Chinese characters. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):183:1–183:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3594634.

Khairunnisa:2023:DEM

[651] Siti Oryza Khairunnisa, Zhousi Chen, and Mamoru Komachi. Dataset enhancement and multilingual transfer for named entity recognition in the Indonesian language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(6):184:1–184:??, June 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3592854.

Droua-Hamdani:2023:MSR

[652] Ghania Droua-Hamdani. MSA speech rhythm pattern in a multilingual setting. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):185:1–185:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3593295.

Taj:2023:USE

[653] Soonh Taj, Ghulam Mujtaba, Sher Muhammad Daudpota, and Muhammad Hussain Mughal. Urdu speech emotion recognition: a systematic literature review. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):186:1–186:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3595377.

Mosa:2023:ELR

[654] Mohamed Atef Mosa. An exhaustive literature review of Hadith text mining. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):187:1–187:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3588315.

Khan:2023:ACT

[655] Sulaiman Khan, Shah Nazir, and Habib Ullah Khan. Analysis of cursive text recognition systems: a systematic literature review. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):188:1–188:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3592600.

Gharagozlou:2023:SRE

[656] Hamid Gharagozlou, Javad Mohammadzadeh, Azam Bastanfard, and Saeed Shiry Ghidary. Semantic relation extraction: a review of approaches, datasets, and evaluation methods with looking at the methods and datasets in the Persian language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):189:1– 189:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3592601.

Kivaisi:2023:SSD

[657] Alexander R. Kivaisi, Qingjie Zhao, and Jimmy T. Mbelwa. Swahili speech dataset development and improved pre-training method for spoken digit recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):190:1–190:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3597494.

Chakrabarty:2023:LRM

[658] Abhisek Chakrabarty, Raj Dabre, Chenchen Ding, Masao Utiyama, and Eiichiro Sumita. Low-resource multilingual neural translation using linguistic feature-based relevance mechanisms. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7): 191:1–191:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3594631.

Ding:2023:MMA

[659] Ling Ding, Xiaojun Chen, Jian Wei, and Yang Xiang. MABERT: Mask-attention-based BERT for Chinese event extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):192:1–192:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3597455.

Yan:2023:AMT

[660] Yibo Yan, Peng Zhu, Dawei Cheng, Fangzhou Yang, and Yifeng Luo. Adversarial multi-task learning for efficient Chinese named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):193:1–193:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3603626.

Mekki:2023:TTA

[661] Asma Mekki, Inès Zribi, Mariem Ellouze, and Lamia Hadrich Belguith. Tokenization of Tunisian Arabic: a comparison between three machine learning models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):194:1–194:??, July 2023.

CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3599234.

Mirishkar:2023:ICC

[662] Ganesh S. Mirishkar, Vishnu Vidyadhara Raju V, Meher Dinesh Naroju, Sudhamay Maity, Prakash Yalla, and Anil Kumar Vuppala. IIITH-CSTD Corpus: Crowdsourced strategies for the collection of a large-scale Telugu speech corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):195:1–195:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3600228.

Yue:2023:RMG

[663] Jianyu Yue, Yiwen Sun, Xiaojun Bi, Zheng Chen, and Yu Zhang. Retrospective multi-granularity fusion network for Chinese idiom cloze-style reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):196:1–196:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3603370.

Sangsavate:2023:ESL

[664] Suntarin Sangsavate, Sukree Sinthupinyo, and Achara Chandrachai. Experiments of supervised learning and semi-supervised learning in Thai financial news sentiment: a comparative study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22 (7):197:1–197:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3603499.

Kaur:2023:BEA

[665] Harmandeep Kaur, Munish Kumar, Aastha Gupta, Monika Sachdeva, Ajay Mittal, and Krishan Kumar. Bagging: an ensemble approach for recognition of handwritten place names in Gurumukhi script. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):198:1–198:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3593024.

Jin:2023:IAP

[666] Feihu Jin, Jinliang Lu, Jiajun Zhang, and Chengqing Zong. Instanceaware prompt learning for language understanding and generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):199:1–199:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3604613.

Al-Jarrah:2023:ABS

[667] Ibrahim Al-Jarrah, Ahmad M. Mustafa, and Hassan Najadat. Aspectbased sentiment analysis for Arabic food delivery reviews. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):200:1–200:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3605146.

Salloum:2023:NEA

[668] Said Salloum, Tarek Gaber, Sunil Vadera, and Khaled Shaalan. A new English/Arabic parallel corpus for phishing emails. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):201:1–201:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3606031.

Sahu:2023:SCB

[669] Siba Sankar Sahu and Sukomal Pal. A study on corpus-based stopword lists in Indian language IR. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(7):202:1– 202:??, July 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3606262.

Cao:2023:ORS

[670] Jihua Cao, Jie Li, Miao Yin, and Yunfeng Wang. Online reviews sentiment analysis and product feature improvement with deep learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):203:1–203:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3522575.

Mujahid:2023:ACT

[671] Muhammad Mujahid, Khadija Kanwal, Furqan Rustam, Wajdi Aljadani, and Imran Ashraf. Arabic ChatGPT tweets classification using RoBERTa and BERT ensemble model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):204:1– 204:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3605889.

Al-Omari:2023:SWG

[672] Hani Al-Omari and Rehab Duwairi. So2al-wa-Gwab: a new Arabic question-answering dataset trained on answer extraction models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):205:1–205:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3605550.

Zhu:2023:QCE

[673] Wenhao Zhu, Xiaoyu Zhang, Liang Ye, and Qiuhong Zhai. Query context expansion for open-domain question answering. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(8):206:1–206:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3603498.

An:2023:PBL

[674] Bo An. Prompt-based for low-resource Tibetan text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):207:1–207:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3603168.

Djaidri:2023:CSL

[675] Asma Djaidri, Hassina Aliane, and Hamid Azzoune. The contribution of selected linguistic markers for unsupervised Arabic verb sense disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):208:1–208:??, August 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3605777.

Lalrempuii:2023:IUN

[676] Candy Lalrempuii and Badal Soni. Investigating unsupervised neural machine translation for low-resource language pair English–Mizo via lexically enhanced pre-trained language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(8):209:1–209:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3609222.

Fenta:2023:AII

[677] Anduamlak Abebe Fenta and Seffi Gebeyehu. Automatic idiom identification model for Amharic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):210:1– 210:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3606864.

Rai:2023:DLB

[678] Pooja Rai, Sanjay Chatterji, and Byung-Gyu Kim. Deep learning-based sequence labeling tools for Nepali. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):211:1– 211:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3606696.

Trujillo-Romero:2023:MSL

[679] Felipe Trujillo-Romero and Gibran García-Bautista. Mexican sign language corpus: Towards an automatic translator. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(8):212:1–212:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3591471.

Lim:2023:ILM

[680] Ee Suan Lim, Wei Qi Leong, Thanh Ngan Nguyen, Wei Ming Kng, William Chandra Tjhi, Dea Adhista, and Ayu Purwarianti. ICON: a linguistically-motivated large-scale benchmark Indonesian constituency Treebank. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):213:1–213:??, August 2023. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3609798.

Osman:2023:EFD

[681] Taha Osman, Hussein Khalil, Mohammed Miltan, Khaled Shaalan, and Rowida Alfrjani. Exploiting functional discourse grammar to enhance complex Arabic relation extraction using a hybrid semantic knowledge base — machine learning approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):214:1– 214:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3610581.

Song:2023:SSS

[682] Haiyue Song, Raj Dabre, Chenhui Chu, Sadao Kurohashi, and Eiichiro Sumita. SelfSeg: a self-supervised sub-word segmentation method for neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):215:1– 215:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3610611.

Xie:2023:CGM

[683] Zheyong Xie, Weidong He, Tong Xu, Shiwei Wu, Chen Zhu, Ping Yang, and Enhong Chen. Comprehending the gossips: Meme explanation in time-sync video comment via multimodal cues. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(8):216:1–216:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3612920.

Dahou:2023:DSA

[684] Brahim Dahou, Leila Falek, Mourad Abbas, Slimane Mekaoui, Mohamed Lichouri, and Aicha Zitouni. DZ-SMS: an authentic corpus of Algerian SMS. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):217:1–217:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3610522.

Thin:2023:SLR

[685] Dang Van Thin, Duong Ngoc Hao, and Ngan Luu-Thuy Nguyen. A systematic literature review on Vietnamese aspect-based sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):218:1–218:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3610226.

Do:2023:SFE

[686] Duc-Hao Do, Thanh-Duc Chau, and Thai-Son Tran. Speech feature enhancement based on time-frequency analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(8):219:1–219:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3605549.

Haulai:2023:CME

[687] Thangkhanhau Haulai and Jamal Hussain. Construction of Mizo:English parallel corpus for machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(8):220:1– 220:??, August 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3610404.

Huang:2023:HCC

[688] Lihua Huang and Peng Zheng. Human-computer collaborative visual design creation assisted by artificial intelligence. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(9):221:1–221:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3554735.

L:2023:EST

[689] Anand Babu G. L. and Srinivasu Badugu. Extractive summarization of Telugu text using modified text rank and maximum marginal relevance. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):222:1–222:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3600224.

Kim:2023:NKN

[690] Hwichan Kim, Hirasawa Tosho, Sangwhan Moon, Naoaki Okazaki, and Mamoru Komachi. North Korean neural machine translation through South Korean resources. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):223:1– 223:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3608947.

Ahmat:2023:WXI

[691] Ahtamjan Ahmat, Yating Yang, Bo Ma, Rui Dong, Kaiwen Lu, and Lei Wang. WAD-X: Improving zero-shot cross-lingual transfer via adapter-based word alignment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):224:1– 224:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3610289.

Zhang:2023:ISS

[692] Weizhao Zhang and Hongwu Yang. Improving sequence-to-sequence Tibetan speech synthesis with prosodic information. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(9):225:1–225:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3616012.

Cao:2023:CSR

[693] Yidan Cao, Qingshan Wang, Qi Wang, and Peng Liu. Can same-right-and-different-left gestures be recognized with only right-hand signals?

ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):226:1–226:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3617370.

Mehmood:2023:EML

[694] Faiza Mehmood, Rehab Shahzadi, Hina Ghafoor, Muhammad Nabeel Asim, Muhammad Usman Ghani, Waqar Mahmood, and Andreas Dengel. EnML: Multi-label ensemble learning for Urdu text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):227:1–227:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3616111.

Al-Ibrahim:2023:DHS

[695] Rogayah M. Al-Ibrahim, Mostafa Z. Ali, and Hassan M. Najadat. Detection of hateful social media content for Arabic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):228:1–228:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3592792.

Chaudhari:2023:MTE

[696] Prasad Chaudhari, Pankaj Nandeshwar, Shubhi Bansal, and Nagendra Kumar. MahaEmoSen: Towards emotion-aware multimodal Marathi sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):229:1–229:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3618057.

Li:2023:IGA

[697] Wanting Li, Yiting Chen, and Buzhou Tang. Improving generative adversarial network-based vocoding through multi-scale convolution. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):230:1–230:??, September 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3610532.

Gezmu:2023:MBN

[698] Andargachew Mekonnen Gezmu and Andreas Nürnberger. Morphemebased neural machine translation models for low-resource fusion languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(9):231:1–231:??, September 2023.

CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3610773.

Yonghe:2023:CSS

[699] Wang Yonghe, Feilong Bao, and Gaunglai Gao. A comparative study on selecting acoustic modeling units for WFST-based Mongolian speech recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):232:1–232:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3617830.

Rayala:2023:SAC

[700] Upendar Rao Rayala, Karthick Seshadri, and Nagesh Bhattu Sristy. Sentiment analysis of code-mixed Telugu–English data leveraging syllable and word embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):233:1–233:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3620670.

Zoya:2023:AUL

[701] Zoya, Seemab Latif, Rabia Latif, Hammad Majeed, and Nor Shahida Mohd Jamail. Assessing Urdu language processing tools via statistical and outlier detection methods on Urdu tweets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10): 234:1–234:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3622939.

Pathak:2023:PST

[702] Dhrubajyoti Pathak, Sukumar Nandi, and Priyankoo Sarmah. Part-of-speech tagger for Assamese using ensembling approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):235:1–235:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3617653.

Zhao:2023:DDI

[703] Shu Zhao, Weifeng Liu, Jie Chen, and Xiao Sun. DIEU: a dynamic interaction emotion unit for emotion recognition in conversation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):236:1–236:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3616493.

Bensalem:2023:APP

[704] Raja Bensalem, Kais Haddar, and Philippe Blache. An Arabic probabilistic parser based on a property grammar. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10): 237:1–237:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3612921.

Zheng:2023:MEM

[705] Cheng Zheng, Haojie Xu, and Xiao Sun. MHG-ERC: Multi-hypergraph feature aggregation network for emotion recognition in conversations. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):238:1–238:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3622935.

Kumari:2023:HTS

[706] Namrata Kumari and Pardeep Singh. Hindi text summarization using sequence to sequence neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10): 239:1–239:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3624013.

Duong:2023:FCV

[707] Huong T. Duong, Van H. Ho, and Phuc Do. Fact-checking Vietnamese information using Knowledge Graph, Datalog, and KG-BERT. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10):240:1–240:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3624557.

Mishra:2023:GGA

[708] Santosh Kumar Mishra, Soham Chakraborty, Sriparna Saha, and Pushpak Bhattacharyya. GAGPT-2: a geometric attention-based GPT-2 framework for image captioning in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(10): 241:1–241:??, October 2023. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3622936.

BenMesmia:2023:SAB

[709] Fatma Ben Mesmia and Malek Mouhoub. Semi-automatic building and learning of a multilingual ontology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):242:1–242:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3615864.

Xie:2023:CED

[710] Dongdong Xie, Fei Li, Bobo Li, Chong Teng, Donghong Ji, and Meishan Zhang. Chinese event discourse deixis resolution: Design of the dataset and model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):243:1–243:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3618109.

Chen:2023:TBM

[711] Yaqi Chen, Wenlin Zhang, Hao Zhang, Dan Qu, and Xu-Kui Yang. Taskbased meta focal loss for multilingual low-resource speech recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):244:1–244:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3626187.

Ghasemi:2023:HDC

[712] Rouzbeh Ghasemi and Saeedeh Momtazi. How a deep contextualized representation and attention mechanism justifies explainable cross-lingual sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):245:1–245:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3626094.

Haffar:2023:SBL

[713] Nafaa Haffar and Mounir Zrigui. A synergistic bidirectional LSTM and N-gram multi-channel CNN approach based on BERT and FastText for Arabic event identification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):246:1– 246:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3626568.

Saidi:2023:SBC

[714] Rakia Saidi and Fethi Jarray. Stacking of BERT and CNN models for Arabic word sense disambiguation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):247:1–247:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3623379.

Ma:2023:RST

[715] Tianlong Ma, Xiangcheng Du, Xingjiao Wu, Zhao Zhou, Yingbin Zheng, and Cheng Jin. Reading scene text with aggregated temporal convolutional encoder. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):248:1–248:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3625822.

Chang:2023:STB

[716] Yung-Chun Chang, Siu Hin Ng, Jung-Peng Chen, Yu-Chi Liang, and Wen-Lian Hsu. Semantic template-based convolutional neural network for text classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):249:1– 249:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3627820.

Li:2023:MTP

[717] Da Li, Boqing Zhu, Sen Yang, Kele Xu, Ming Yi, Yukai He, and Huaimin Wang. Multi-task pre-training language model for semantic network completion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):250:1–250:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3627704.

Zhang:2023:FGD

[718] Meishan Zhang, Peiming Guo, Peijie Jiang, Dingkun Long, Yueheng Sun, Guangwei Xu, Pengjun Xie, and Min Zhang. Fine-grained domain adaptation for Chinese syntactic processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(11):251:1–251:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3629519.

Tang:2023:SES

[719] Xiaoyu Tang, Mengyun Zheng, Jiewen Feng, Jiazheng Huang, and Yayun Gong. Shortcut enhanced syntactic and semantic dual-channel network for aspect-based sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):252:1–252:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3629518.

Jin:2023:FKP

[720] Weiqiang Jin, Biao Zhao, Yu Zhang, Gege Sun, and Hang Yu. Fintech Key-Phrase: a new Chinese financial high-tech dataset accelerating expression-level information retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(11):253:1–253:??, November 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3627989.

Prakash:2023:CLS

[721] Jothi Prakash V and Arul Antran Vijay S. Cross-lingual sentiment analysis of Tamil language using a multi-stage deep learning architecture. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):254:1–254:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3631391.

Belay:2023:FRN

[722] Birhanu Hailu Belay, Gebeyehu Belay Gebremeskel, Belete Biazen Bezabih, and Seffi Gebeyehu. Factorized recurrent neural network with attention for language identification and content detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):255:1–255:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3630607.

Jia:2023:SAI

[723] Zhaohong Jia, Yunwei Shi, Weifeng Liu, Zhenhua Huang, and Xiao Sun. Speaker-aware interactive graph attention network for emotion recognition in conversation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):256:1–256:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3627806.

Hamadouche:2023:AAC

[724] Khaoula Hamadouche, Kheira Zineb Bousmaha, Mohamed Abdelwaret Bekkoucha, and Lamia Hadrich-Belguith. AlgBERT: Automatic construction of annotated corpus for sentiment analysis in Algerian dialect. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):257:1–257:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3632948.

Modak:2023:ALC

[725] Masooda M. Modak, Prachi Gharpure, and Sasikumar M. Adaptive learning and correlative assessment of differential usage patterns for students with-or-without learning disabilities via learning analytics. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):258:1–258:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3632365.

Li:2023:DBB

[726] Daiyi Li, Li Yan, and Zongmin Ma. Dependency-based BERT for Chinese event argument extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):259:1– 259:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3633306.

Bensalah:2023:CSD

[727] Nouhaila Bensalah, Habib Ayad, Abdellah Adib, and Abdelhamid Ibn El Farouk. A comparative study of different dimensionality reduction techniques for Arabic machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):260:1–260:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3634681.

Husain:2023:NKA

[728] Fatemah Husain. A novel knowledge-augmented model customization approach for Arabic offensive language detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 22(12):261:1–261:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3634702.

Ganjalipour:2023:NER

[729] Ebrahim Ganjalipour, Amir Hossein Refahi Sheikhani, Sohrab Kordrostami, and Ali Asghar Hosseinzadeh. Named entity recognition in Persian language based on self-attention mechanism with weighted relational position encoding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 22(12):262:1–262:??, December 2023. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3633513.

Ahmed:2024:EIA

[730] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Emotional intelligence attention unsupervised learning using lexicon analysis for irony-based advertising. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):1:1–1:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3580496.

Madan:2024:CFS

[731] Chetan Madan, Harshita Diddee, Deepika Kumar, and Mamta Mittal. CodeFed: Federated speech recognition for low-resource code-switching detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):2:1–2:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3571732.

Alyami:2024:IAS

[732] Sarah Alyami, Hamzah Luqman, and Mohammad Hammoudeh. Isolated Arabic sign language recognition using a transformer-based model and landmark keypoints. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):3:1–3:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3584984.

Tejaswini:2024:DDS

[733] Vankayala Tejaswini, Korra Sathya Babu, and Bibhudatta Sahoo. Depression detection from social media text analysis using natural language processing techniques and hybrid deep learning model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):4:1–4:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3569580.

Kumar:2024:SCU

[734] Ankit Kumar, Surbhi Bhatia, Mohammad R. Khosravi, Arwa Mashat, and Parul Agarwal. Semantic and context understanding for sentiment analysis in Hindi handwritten character recognition using a multiresolution technique. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):5:1–5:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3557895.

Ren:2024:ICC

[735] Tianyu Ren, Dengfeng Yao, Chaoran Yang, and Xinchen Kang. The influence of Chinese characters on Chinese sign language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):6:1–6:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3591465.

Gambhir:2024:EEM

[736] Pooja Gambhir, Amita Dev, Poonam Bansal, and Deepak Kumar Sharma. End-to-end multi-modal low-resourced speech keywords recognition using sequential Conv2D nets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):7:1– 7:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3606019.

Roy:2024:DEN

[737] Pradeep Kumar Roy. Deep ensemble network for sentiment analysis in bi-lingual low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):8:1– 8:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3600229.

Kumar:2024:NSD

[738] Sanjay Kumar. Negative stances detection from multilingual data streams in low-resource languages on social media using BERT and CNN-based transfer learning model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):9:1– 9:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3625821.

Zheng:2024:UII

[739] Jiangbo Zheng and Ying Liang. User interest identification with social media information using natural language and meta-heuristic technique. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):10:1–10:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3579165.

Sharma:2024:EWB

[740] Vijay Sharma, Namita Mittal, Ankit Vidyarthi, and Deepak Gupta. Exploring Web-based translation resources applied to Hindi–English crosslingual information retrieval. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):11:1–11:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3569010.

Rezaee:2024:CYU

[741] Khosro Rezaee, Hossein Ghayoumi Zadeh, Lianyong Qi, Hamidreza Rabiee, and Mohammad R. Khosravi. Can you understand why I am crying? A decision-making system for classifying infants’ cry languages based on DeepSVM model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):12:1–12:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3579032.

Chakraborty:2024:ECH

[742] Angana Chakraborty, Subhankar Joardar, and Arif Ahmed Sekh. Ensemble classifier for Hindi hostile content detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(1):13:1–13:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3591353.

Wei:2024:MLT

[743] Xiao Wei, Jianbao Huang, Rui Zhao, Hang Yu, and Zheng Xu. Multilabel text classification model based on multi-level constraint augmentation and label association attention. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):14:1– 14:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3586008.

Bozuyla:2024:SAT

[744] Mehmet Bozuyla. Sentiment analysis of Turkish drug reviews with bidirectional encoder representations from transformers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):15:1–15:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3626523.

Dash:2024:RMD

[745] Deba Prasad Dash, Maheshkumar Kolekar, Chinmay Chakraborty, and Mohammad R. Khosravi. Review of machine and deep learning techniques in epileptic seizure detection using physiological signals and sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):16:1–16:??, January 2024. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3552512.

Jain:2024:OBN

[746] Deepak Kumar Jain, Shamimul Qamar, Saurabh Raj Sangwan, Weiping Ding, and Anand J. Kulkarni. Ontology-based natural language processing for sentimental knowledge analysis using deep learning architectures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):17:1–17:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3624012.

Dave:2024:SRS

[747] Nakul R. Dave, Mayuri A. Mehta, and Ketan Kotecha. A systematic review of stemmers of Indian and non-Indian vernacular languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(1):18:1–18:??, January 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3604612.

Dalai:2024:DLB

[748] Tusarkanta Dalai, Tapas Kumar Mishra, and Pankaj K. Sa. Deep learning-based POS tagger and chunker for Odia language using pretrained transformers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):19:1–19:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3637877.

Man:2024:ESG

[749] Zhibo Man, Yujie Zhang, Yu Li, Yuanmeng Chen, Yufeng Chen, and Jinan Xu. An ensemble strategy with gradient conflict for multi-domain neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):20:1–20:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3638248.

Zheng:2024:HNN

[750] Cheng Zheng, Haojie Xu, and Xiao Sun. Hypergraph neural network for emotion recognition in conversations. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):21:1– 21:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3638760.

Chungnoi:2024:CMS

[751] Krittanut Chungnoi, Rachada Kongkachandra, and Sarun Gulyanon. The computational method for supporting Thai VerbNet construction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):22:1–22:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3638533.

Sweidan:2024:AFE

[752] Asmaa Hashem Sweidan, Nashwa El-Bendary, and Esraa Elhariri. Autoregressive feature extraction with topic modeling for aspect-based sentiment analysis of Arabic as a low-resource language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):23:1–23:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3638050.

Wang:2024:DBM

[753] Haixia Wang, Yingyu Mao, Qingran Miao, Qun Xiao, and Yilong Zhang. Dual-branch multitask fusion network for offline Chinese writer identification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):24:1–24:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3638554.

Oriola:2024:IDM

[754] Oluwafemi Oriola and Eduan Kotzé. Improving the detection of multilingual South African abusive language via skip-gram using joint multilevel domain adaptation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):25:1–25:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3638759.

Jiao:2024:REA

[755] Songlin Jiao, Zhenfang Zhu, Jiangtao Qi, Fuyong Xu, Hongli Pei, Wenling Wang, Ze Song, and Peiyu Liu. A relation embedding assistance networks for multi-hop question answering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2): 26:1–26:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3635114.

Li:2024:DNN

[756] Lingfang Li, Aijun Zhang, and Ming-Xing Luo. DSISA: a new neural machine translation combining dependency weight and neighbors. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):27:1–27:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3638762.

Lin:2024:LDG

[757] Tzu-Mi Lin, Man-Chen Hung, and Lung-Hao Lee. Leveraging dual gloss encoders in Chinese biomedical entity linking. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(2):28:1–28:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3638555.

Nazih:2024:IGI

[758] Waleed Nazih, Amany Fashwan, Amr El-Gendy, and Yasser Hifny. Ibn-Ginni: an improved morphological analyzer for Arabic. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):29:1–29:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3639050.

Bhogayata:2024:MLB

[759] Chandrakant K. Bhogayata. A machine learning-based readability model for Gujarati texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):30:1–30:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3637826.

Mi:2024:MGK

[760] Chenggang Mi, Shaoliang Xie, and Yi Fan. Multi-granularity knowledge sharing in low-resource neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):31:1–31:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3639930.

Yang:2024:EGK

[761] Zhao Yang, Yuanzhe Zhang, Dianbo Sui, Yiming Ju, Jun Zhao, and Kang Liu. Explanation guided knowledge distillation for pre-trained language model compression. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):32:1–32:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3639364.

Baruah:2024:TCR

[762] Hemanta Baruah, Sanasam Ranbir Singh, and Priyankoo Sarmah. Transliteration characteristics in Romanized Assamese language social media text and machine transliteration. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2): 33:1–33:??, February 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3639565.

Wang:2024:FSI

[763] Hao Wang, Hanwen Shi, and Jianyong Duan. Few-shot incremental event detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(2):34:1–34:??, February 2024. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3634747.

Tohidi:2024:PPA

[764] Nasim Tohidi, Chitra Dadkhah, Reza Nouralizadeh Ganji, Ehsan Ghaffari Sadr, and Hoda Elmi. PAMR: Persian abstract meaning representation corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):35:1–35:??, March 2024. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3638288.

Nasayreh:2024:ASA

[765] Ahmad Nasayreh, Rabia Emhamed Al Mamlook, Ghassan Samara, Hasan Gharaibeh, Mohammad Aljaidi, Dalia Alzu’bi, Essam Al-Daoud, and Laith Abualigah. Arabic sentiment analysis for ChatGPT using machine learning classification algorithms: a hyperparameter optimization technique. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):36:1–36:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3638285.

Das:2024:ESI

[766] Soumen Das, Saroj Kr. Biswas, and Biswajit Purkayastha. An expert system for Indian sign language recognition using spatial attention-based feature and temporal feature. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):37:1–37:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3643824.

V:2024:MNA

[767] Lakshmi Lalitha V. and Dinesh Kumar Anguraj. Modeling a novel approach for emotion recognition using learning and natural language processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):38:1–38:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3641851.

Li:2024:IBB

[768] Qing Li, Weibin Wan, Yuming Zhao, and Xiaoyan Jiang. Improved BIO-based Chinese automatic abstract-generation model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):39:1–39:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3643695.

N:2024:MVI

[769] Thenmoezhi N., Perumal B., and Lakshmi A. Multi-view image fusion using ensemble deep learning algorithm for MRI and CT images. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):40:1–40:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3640811.

Liu:2024:STS

[770] Jie Liu, Yaguang Li, Shizhu He, Shun Wu, Kang Liu, Shenping Liu, Jiong Wang, and Qing Zhang. Seq2Set2Seq: a two-stage disentangled method for reply keyword generation in social media. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (3):41:1–41:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3644074.

Banik:2024:IRA

[771] Debajyoty Banik, Rahul Paul, Rajkumar Singh Rathore, and Rutvij H. Jhaveri. Improved regression analysis with ensemble pipeline approach for applications across multiple domains. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3): 42:1–42:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3645110.

Rafi:2024:SSC

[772] Shaik Rafi and Ranjita Das. SCT: Summary caption technique for retrieving relevant images in alignment with multimodal abstractive summary. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):43:1–43:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3645029.

Devi:2024:DIM

[773] Thiyam Susma Devi and Pradip K. Das. Disambiguation of isolated Manipuri tonal contrast word pairs using acoustic features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):44:1–44:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3643830.

Bi:2024:CCL

[774] Zhen Bi, Jing Chen, Yinuo Jiang, Feiyu Xiong, Wei Guo, Huajun Chen, and Ningyu Zhang. CodeKGC: Code language model for generative knowledge graph construction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):45:1–45:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3641850.

Dargan:2024:GCS

[775] Shaveta Dargan and Munish Kumar. Gender classification system based on the behavioral biometric modality: Application of handwritten text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):46:1–46:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3626236.

Garg:2024:TMH

[776] Muskan Garg. Towards mental health analysis in social media for lowresourced languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):47:1–47:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3638761.

Kanwal:2024:SSE

[777] Safia Kanwal, Muhammad Kamran Malik, Zubair Nawaz, and Khawar Mehmood. SEEUNRS: Semantically enriched entity-based Urdu news recommendation system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):48:1–48:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3639049.

Mahata:2024:CBM

[778] Sainik Kumar Mahata, Dipankar Das, and Sivaji Bandyopadhyay. Consensus-based machine translation for code-mixed texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(3):49:1–49:??, March 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3628427.

Yu:2024:SAM

[779] Yang Yu, Dong Qiu, and Huanyu Wan. Sentiment analysis method of epidemic-related microblog based on hesitation theory. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):50:1–50:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3648360.

Yuan:2024:CLK

[780] Xiaowei Yuan, Kang Liu, and Yequan Wang. Contrastive languageknowledge graph pre-training. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):51:1–51:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3644820.

Ahmed:2024:EUN

[781] Anil Ahmed, Degen Huang, Syed Yasser Arafat, and Imran Hameed. Enriching Urdu NER with BERT embedding, data augmentation, and hybrid encoder-CNN architecture. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):52:1–52:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3648362.

Li:2024:BAA

[782] Jiuyi Li, Junpeng Liu, Jianjun Ma, Wei Yang, and Degen Huang. Boundary-aware abstractive summarization with entity-augmented attention for enhancing faithfulness. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):53:1– 53:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3641278.

San:2024:SEL

[783] Mya Ei San, Sasiporn Usanavasin, Ye Kyaw Thu, and Manabu Okumura. A study for enhancing low-resource Thai–Myanmar–English neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):54:1–54:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3645111.

Tayir:2024:UMM

[784] Turghun Tayir and Lin Li. Unsupervised multimodal machine translation for low-resource distant language pairs. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):55:1– 55:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3652161.

Lu:2024:MQS

[785] Wenpeng Lu, Sibo Wei, Xueping Peng, Yi-Fei Wang, Usman Naseem, and Shoujin Wang. Medical question summarization with entity-driven contrastive learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):56:1–56:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3652160.

Akanova:2024:NSS

[786] Akerke Akanova, Aisulu Ismailova, Zhanar Oralbekova, Zhanat Kenzhebayeva, and Galiya Anarbekova. Neurocomputer system of semantic analysis of the text in the Kazakh language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4): 57:1–57:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3652159.

Zaman:2024:LBL

[787] Farooq Zaman, Onaiza Maqbool, and Jaweria Kanwal. Leveraging bidirectionl LSTM with CRFs for Pashto tagging. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (4):58:1–58:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3649456.

Zhao:2024:CDA

[788] Chuanjun Zhao, Meiling Wu, Xinyi Yang, Xuzhuang Sun, Suge Wang, and Deyu Li. Cross-domain aspect-based sentiment classification with a pre-training and fine-tuning strategy for low-resource domains. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(4):59:1–59:??, April 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3653299.

Verdu:2024:MLR

[789] Elena Verdu, Yuri Vanessa Nieto, and Nasir Saleem. Multi-lingual representation of natural language processing for low resource Asian language processing systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):60:1–60:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3603169.

Wei:2024:MTS

[790] Kaiwen Wei, Li Jin, Zequn Zhang, Zhi Guo, Xiaoyu Li, Qing Liu, and Weimiao Feng. More than syntaxes: Investigating semantics to zeroshot cross-lingual relation extraction and event argument role labelling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):61:1–61:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3582261.

Yu:2024:RUS

[791] Shanshan Yu. A research on university students’ behavioral intention to use new-generation information technology in intelligent foreign language learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):62:1–62:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3563774.

Jain:2024:KBD

[792] Deepak Kumar Jain, Yamila GarcÍa-MartÍnez Eyre, Akshi Kumar, Brij B. Gupta, and Ketan Kotecha. Knowledge-based data processing for multilingual natural language analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5): 63:1–63:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3583686.

Ahmed:2024:ATL

[793] Usman Ahmed, Jerry Chun-Wei Lin, and Vicente Garcia Diaz. Automatically temporal labeled data generation using positional lexicon expansion for focus time estimation of news articles. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (5):64:1–64:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3568164.

Das:2024:MNM

[794] Sudhansu Bala Das, Divyajyoti Panda, Tapas Kumar Mishra, Bidyut Kr. Patra, and Asif Ekbal. Multilingual neural machine translation for Indic to Indic languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):65:1–65:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3652026.

Vu:2024:NPG

[795] Dinh Anh Vu, Quang Nhat Minh Pham, and Giang Son Tran. A novel pretrained general-purpose vision language model for the Vietnamese language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):66:1–66:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3654796.

Sarwar:2024:CLB

[796] Raheem Sarwar, Maneesha Perera, Pin Shen Teh, Raheel Nawaz, and Muhammad Umair Hassan. Crossing linguistic barriers: Authorship attribution in Sinhala texts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):67:1–67:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3655620.

C:2024:FRN

[797] Vinotheni C. and S. Lakshmana Pandian. Fast recurrent neural network with bi-LSTM for handwritten Tamil text segmentation in NLP. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):68:1–68:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3643808.

Rong:2024:MMM

[798] Huan Rong, Zhongfeng Chen, Zhenyu Lu, Fan Xu, and Victor S Sheng. Multization: Multi-modal summarization enhanced by multicontextually relevant and irrelevant attention alignment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):69:1–69:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3651983.

Baishya:2024:PST

[799] Diganta Baishya and Rupam Baruah. Part-of-speech tagging for lowresource languages: Activation function for deep learning network to work with minimal training data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):70:1–70:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3655023.

Munivel:2024:PBA

[800] Monisha Munivel and V. S. Felix Enigo. Performance of binarization algorithms on Tamizhi inscription images: an analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):71:1–71:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3656583.

Hou:2024:KEP

[801] Wenlong Hou, Weidong Zhao, Xianhui Liu, and Wenyan Guo. Knowledge-enriched prompt for low-resource named entity recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):72:1–72:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3659948.

Luo:2024:CJN

[802] Queenie Luo and Yung-Sung Chuang. Cleansing jewel: a neural spelling correction model built on Google OCR-ed Tibetan manuscripts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):73:1–73:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3654811.

Wu:2024:SCL

[803] Liang Wu, Fangfang Zhang, Chao Cheng, and Shinan Song. Supervised contrast learning text classification model based on data quality augmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):74:1–74:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3653300.

Chen:2024:MTM

[804] Yi Ting Chen, Wanting Li, and Buzhou Tang. MRMI-TTS: Multireference audios and mutual information driven zero-shot voice cloning.

ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):75:1–75:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3649501.

Wang:2024:ADG

[805] Chunmei Wang, Yuan Luo, Chunli Meng, and Feiniu Yuan. An adaptive dual graph convolution fusion network for aspect-based sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):76:1–76:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3659579.

Setiawan:2024:SIR

[806] Irwan Setiawan and Hung-Yu Kao. SUSTEM: an improved rulebased Sundanese stemmer. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):77:1–77:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3656342.

Huang:2024:LDS

[807] Shuanghong Huang, Chong Feng, Ge Shi, Zhengjun Li, Xuan Zhao, Xinyan Li, and Xiaomei Wang. Learning domain specific sub-layer latent variable for multi-domain adaptation neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):78:1–78:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3661305.

Bekarystankyzy:2024:IEE

[808] Akbayan Bekarystankyzy, Orken Mamyrbayev, and Tolganay Anarbekova. Integrated end-to-end automatic speech recognition for languages for agglutinative languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):79:1– 79:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3663568.

Zhang:2024:NMT

[809] Jinyi Zhang, Ke Su, Haowei Li, Jiannan Mao, Ye Tian, Feng Wen, Chong Guo, and Tadahiro Matsumoto. Neural machine translation for lowresource languages from a Chinese-centric perspective: a survey. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):80:1–80:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3665244.

Zhong:2024:RIA

[810] Zhengwu Zhong. Research on the implementation of advertising design teaching based on Unity3D development platform and Web3D technology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):81:1–81:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3595294.

Yin:2024:AII

[811] Jie Yin. Application of intelligent image recognition and digital media art in the inheritance of black pottery intangible cultural heritage. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):82:1–82:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3597430.

Gao:2024:IMN

[812] Yuan Gao, Yani Wu, and Jun Qian. Intelligent multimedia network security and PBL teaching mode in the basic course teaching of college design major. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):83:1–83:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3597429.

Wang:2024:SAC

[813] Jie Wang, Jiangjun Yuan, and Weinan Liu. Sentiment analysis and corpus: Cognitive perspective and overhead-accuracy tradeoff. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):84:1–84:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3594537.

Yang:2024:APE

[814] Lili Yang and Yue Zhang. Application of psychological effects in intelligent digital English teaching based on multimodal low resource language information processing from a psychological perspective. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):85:1–85:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3599726.

Wang:2024:VSI

[815] Yifang Wang. Virtual sound image reconstruction method for multiobjective optimization of folk music based on evolutionary algorithm. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):86:1–86:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3604614.

Tu:2024:COE

[816] Haihua Tu. Collaborative optimization of English online teaching informatization based on intelligent multimedia image technology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):87:1–87:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3599725.

Li:2024:EAE

[817] Tianhua Li and Shaowei Qu. Effectiveness analysis of entrepreneurial legal risk prevention based on multi-modal deep learning model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):88:1–88:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3622937.

Wang:2024:RRM

[818] Huaben Wang and Jixueyang Tang. Research on recognition method of social robot based on T-A-GCNIIT in the Metaverse. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):89:1–89:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3624014.

Zhong:2024:RPA

[819] Zhengwu Zhong. Research on product advertising design combining feature extraction technology and Web3D technology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (6):90:1–90:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3608948.

Huang:2024:MLN

[820] Yong Huang and Lu Zhang. Machine learning and natural language processing algorithms in the remote mobile medical diagnosis system of Internet hospitals. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):91:1–91:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3632172.

Fu:2024:IVL

[821] Liwei Fu and Qiang Li. Investigation of visual language landscape of tourist attractions from multimodal perspective. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (6):92:1–92:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3638049.

Han:2024:CPT

[822] Jing Han. Construction of practical teaching mode of law course based on multi-mode and low-resource language learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23 (6):93:1–93:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3622938.

Cao:2024:AIT

[823] Yejun Cao, Xiwen Yu, and Fengling Jiang. Application of 3D image technology in rural planning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):94:1–94:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3628448.

Feng:2024:IDM

[824] Qiao Feng and Tian Huang. An interaction-design method based upon a modified algorithm of Newton’s Second Law of Motion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(6):95:1–95:??, June 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3657634.

Sun:2024:TBQ

[825] Runxin Sun, Shizhu He, Jun Zhao, and Kang Liu. Towards better quantity representations for solving math word problems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):96:1–96:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3665644.

Qiu:2024:QSA

[826] Bing Qiu and Jiahao Huo. Quantitative stylistic analysis of middle Chinese texts based on the dissimilarity of evolutive core word usage. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):97:1–97:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3665794.

Kumar:2024:HHD

[827] Akshi Kumar, Abhishek Mallik, and Sanjay Kumar. HumourHindiNet: Humour detection in Hindi web series using word embedding and convolutional neural network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):98:1–98:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3661306.

Lalramhluna:2024:MMB

[828] Robert Lalramhluna, Sandeep Dash, and Dr.Partha Pakray. MizBERT: a Mizo BERT model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):99:1–99:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3666003.

Wang:2024:DLR

[829] Xiaoyi Wang, Jie Liu, Jiong Wang, Jianyong Duan, Guixia Guan, Qing Zhang, and Jianshe Zhou. Document-level relation extraction based on machine reading comprehension and hybrid pointer-sequence labeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):100:1–100:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3666042.

Xu:2024:SSC

[830] Yangyang Xu, Zhuoer Zhao, and Xiao Sun. SCBG: Semantic-constrained bidirectional generation for emotional support conversation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):101:1–101:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3666090.

Guo:2024:XPW

[831] Shih-Wei Guo and Yao-Chung Fan. X-Phishing-Writer: a framework for cross-lingual phishing e-mail generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7): 102:1–102:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3670402.

Pahari:2024:SWY

[832] Niraj Pahari and Kazutaka Shimada. Share what you already know: Cross-language-script transfer and alignment for sentiment detection in code-mixed data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):103:1–103:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3661307.

Aithal:2024:KLD

[833] Shreya R. Aithal, Muralikrishna Sn, Raghavendra Ganiga, Ashwath Rao B., and Govardhan Hegde K. KannadaLex: a lexical database with psycholinguistic information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):104:1– 104:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3670688.

Salhab:2024:AAS

[834] Mahmoud Salhab and Haidar Harmanani. AraSpot: Arabic spoken command spotting. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):105:1–105:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3674968.

Li:2024:ECE

[835] Fei Li, Kaifang Deng, Yiwen Mo, Yuanze Ji, Chong Teng, and Donghong Ji. Enhancing Chinese event extraction with event trigger structures. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):106:1–106:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3663567.

Sidhoum:2024:SMH

[836] Abdellah Hamouda Sidhoum, Mhamed Mataoui, Faouzi Sebbak, Adil Imad Eddine Hosni, and Kamel Smaili. Scoring multi-hop question decomposition using masked language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7): 107:1–107:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3665140.

Bousmaha:2024:AAS

[837] Kheira Zineb Bousmaha, Khaoula Hamadouche, Hadjer Djouabi, and Lamia Hadrich-Belguith. Automatic Algerian sarcasm detection from texts and images. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(7):108:1–108:??, July 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3670403.

He:2024:ENS

[838] Guijiao He, Yunfeng Zhou, and Yaodong Zheng. Evaluation on network social media named entity recognition model based on active learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):109:1–109:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3600055.

Dong:2024:NLP

[839] Jun Dong. Natural language processing pretraining language model for computer intelligent recognition technology. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8): 110:1–110:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3605210.

Zeng:2024:MMC

[840] Zhifa Zeng and Yuhang Li. Multi-modal Chinese text emotion metaphor computation based on mutual information and information entropy. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):111:1–111:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3605211.

Duan:2024:ISM

[841] Xingyu Duan, Chun-Nan Chen, and Mohammad Shokouhifar. Impacts of social media advertising on purchase intention and customer loyalty in e-commerce systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):112:1–112:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3613448.

Perti:2024:CHD

[842] Ashwin Perti, Amit Sinha, and Ankit Vidyarthi. Cognitive hybrid deep learning-based multi-modal sentiment analysis for online product reviews. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):113:1–113:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3615356.

Hamza:2024:MRH

[843] Ameer Hamza, Abdul Rehman Javed, Farkhund Iqbal, Amanullah Yasin, Gautam Srivastava, Dawid Połap, Thippa Reddy Gadekallu, and Zunera Jalil. Multimodal religiously hateful social media memes classification based on textual and image data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):114:1– 114:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3623396.

Yu:2024:TCB

[844] Haitao Yu, Feng Xiong, and Zuhui Chen. Text classification based on natural language processing and machine learning in multi-label corpus. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):115:1–115:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3617831.

Liu:2024:LMM

[845] Jinhui Liu and Feng Zhang. Language model method for collocation rules of parts of speech in machine translation system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):116:1–116:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3625095.

Kumar:2024:HPD

[846] Akshi Kumar, Dipika Jain, and Rohit Beniwal. HindiPersonalityNet: Personality detection in Hindi conversational data using deep learning with static embedding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):117:1–117:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3625228.

Kumar:2024:HDR

[847] Suresh Kumar, Jyoti Prakash Singh, Surya Kant, and Neha Jain. A hybrid deep ranking weighted multi-hashing recommender system. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):118:1–118:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3626195.

Manimaran:2024:EAI

[848] A. Manimaran, Mohammad Haider Syed, M. Siva Kumar, S. Selvanayaki, Gurram Sunitha, and Asmita Manna. Enhancing Asian indigenous language processing through deep learning-based handwriting recognition and optimization techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):119:1– 119:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3632173.

Ahmad:2024:ECD

[849] Pir Noman Ahmad, Yuanchao Liu, Inam Ullah, and Mohammad Shabaz. Enhancing coherence and diversity in multi-class slogan generation systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):120:1–120:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3637551.

Kusal:2024:UPA

[850] Sheetal D. Kusal, Shruti G. Patil, Jyoti Choudrie, and Ketan V. Kotecha. Understanding the performance of AI algorithms in text-based emotion detection for conversational agents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):121:1– 121:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3643133.

Ai:2024:EAI

[851] Qinghua Ai, Qingyan Ai, and Jun Wang. Exploration on advanced intelligent algorithms of artificial intelligence for verb recognition in machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):122:1–122:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3649891.

Su:2024:NLP

[852] Yawen Su. Natural language processing system for text classification corpus based on machine learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):123:1– 123:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3648361.

Naosekpam:2024:HST

[853] Veronica Naosekpam and Nilkanta Sahu. A hybrid scene text script identification network for regional Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(8):124:1–124:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3649439.

Kaur:2024:HEP

[854] Ravleen Kaur, M. P. S. Bhatia, and Akshi Kumar. Am I hurt?: Evaluating psychological pain detection in Hindi text using transformer-based models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):125:1–125:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3650206.

Tai:2024:TAM

[855] Yu Tai, Hongwei Yang, Hui He, Xinglong Wu, Yuanming Shao, Weizhe Zhang, and Arun Kumar Sangaiah. Topic-aware masked attentive network for information cascade prediction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8): 126:1–126:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3653449.

Kumar:2024:HDB

[856] Ashwini Kumar, Santosh Kumar, Kalpdrum Passi, and Aniket Mahanti. A hybrid deep BiLSTM-CNN for hate speech detection in multi-social media. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):127:1–127:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3657635.

Banik:2024:BIP

[857] Debajyoty Banik, Saneyika Das, Sheshikala Martha, and Achyut Shankar. BERT-inspired progressive stacking to enhance spelling correction in Bengali text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(8):128:1–128:??, August 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3669941.

Ma:2024:ALR

[858] Jiushun Ma, Yuxin Huang, Linqin Wang, Xiang Huang, Hao Peng, Zhengtao Yu, and Philip Yu. Augmenting low-resource cross-lingual summarization with progression-grounded training and prompting. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):129:1–129:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3675167.

Inaba:2024:TAT

[859] Michimasa Inaba, Yuya Chiba, Zhiyang Qi, Ryuichiro Higashinaka, Kazunori Komatani, Yusuke Miyao, and Takayuki Nagai. Travel agency task dialogue corpus: a multimodal dataset with age-diverse speakers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):130:1–130:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3675166.

Mishra:2024:MTL

[860] Pruthwik Mishra, Vandan Mujadia, and Dipti Misra Sharma. Multi task learning based shallow parsing for Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):131:1–131:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3664620.

Pham:2024:TVQ

[861] Quoc-Hung Pham, Huu-Loi Le, Minh Dang Nhat, Khang Tran T., Manh Tran-Tien, Viet-Hung Dang, Huy-The Vu, Minh-Tien Nguyen, and Xuan-Hieu Phan. Towards Vietnamese question and answer generation: an empirical study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):132:1–132:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3675781.

Zhao:2024:CCN

[862] Shu Zhao, Zhuoer Zhao, Yangyang Xu, and Xiao Sun. CoMix: Confronting with noisy label learning with co-training strategies on textual mislabeling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):133:1–133:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3678175.

Kaur:2024:AII

[863] Manpreet Kaur and Munish Saini. Artificial intelligence inspired method for cross-lingual cyberhate detection from low resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):134:1–134:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3677176.

Alotaibi:2024:TCD

[864] Fahd Saleh Alotaibi, Khaled Hamed Alyoubi, Ajay Mittal, Vishal Gupta, and Navdeep Kaur. TinyCheXReport: Compressed deep neural network for chest X-ray report generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):135:1– 135:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3676166.

Chowdhury:2024:AET

[865] Priyanjana Chowdhury, Nabanika Sarkar, Sanghamitra Nath, and Utpal Sharma. Analyzing the effects of transcription errors on summary generation of Bengali spoken documents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):136:1– 136:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3678005.

Lu:2024:OUS

[866] Kexin Lu, Zhihua Huang, Mingming Yin, and Ke Chen. Optimizing Uyghur speech synthesis by combining pretrained cross-lingual model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(9):137:1–137:??, September 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3675397.

Pham:2024:XMB

[867] Khang Pham, Long Nguyen, and Dien Dinh. XLIT: a method to bridge task discrepancy in machine translation pre-training. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(10):138:1–138:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3689630.

Torjmen:2024:TTD

[868] Roua Torjmen and Kais Haddar. Translation from Tunisian dialect to Modern Standard Arabic: Exploring finite-state transducers and sequence-to-sequence transformer approaches. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10):139:1–139:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3681788.

Waghmare:2024:EBB

[869] Prachi Pramod Waghmare and Ashwini Mangesh Deshpande. Enhanced BERT-based multi-head self-attention transformer for transformation of Marathi text to Marathi sign language gloss. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(10):140:1–140:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3687304.

Munaf:2024:LRS

[870] Mubashir Munaf, Hammad Afzal, Khawir Mahmood, and Naima Iltaf. Low resource summarization using pre-trained language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10):141:1–141:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3675780.

Madan:2024:CBM

[871] Anjum Madan and Devender Kumar. CNN-based models for emotion and sentiment analysis using speech data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10): 142:1–142:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3687303.

Zhao:2024:TPM

[872] Jinghua Zhao, Xiting Lyu, Haiying Rong, and Jiale Zhao. TRGCN: a prediction model for information diffusion based on transformer and relational graph convolutional network. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10): 143:1–143:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3672074.

Yang:2024:KIP

[873] Songhua Yang, Chenghao Zhang, Chenyuan He, Hongfei Xu, Hongying Zan, and Yuxiang Jia. Knowledge-injected prompt learning for Chinese biomedical entity normalization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10): 144:1–144:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3689629.

Gogoi:2024:CLR

[874] Parismita Gogoi, Priyankoo Sarmah, and S. R. M. Prasanna. Crosslinguistic rhythm analysis of Mising and Assamese. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(10):145:1–145:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3694785.

Zafar:2024:TBT

[875] Amna Zafar, Muhammad Wasim, Shaista Zulfiqar, Talha Waheed, and AbuBakar Siddique. Transformer-based topic modeling for Urdu translations of the Holy Quran. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10):146:1–146:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3694967.

Tao:2024:SDL

[876] Zekun Tao, Changjian Wang, Zhiliang Tian, Kele Xu, Yong Guo, Shanshan Li, Yanru Bai, and Da Xie. SIAT: Document-level event extraction via spatiality-augmented interaction model with adaptive thresholding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(10):147:1–147:??, October 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3698261.

Lin:2024:ISI

[877] Chi Lin, Chang Wu Yu, and Ning Wang. Introduction to the special issue on cognitive-inspired multimedia information processing and applications for low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):148:1– 148:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3676150.

Wu:2024:FFP

[878] Yirui Wu, Lilai Zhang, Hao Li, Yunfei Zhang, and Shaohua Wan. Feature fusion pyramid network for end-to-end scene text detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):149:1–149:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3582003.

Rutherford:2024:ECB

[879] Attapol Rutherford and Pawitsapak Akarajaradwong. Exploring the correlation between emojis and mood expression in Thai Twitter discourse. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):150:1–150:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3680543.

Singh:2024:CRC

[880] Niraj Singh, Komal Naaz, and Raj Aryan. Can rhyme consistency score be used as a feature in stylistics? A statistical endeavour with Hindi poetry. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):151:1–151:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3681789.

Cui:2024:DNN

[881] Xingyu Cui, Yong Li, and Lili Xu. Deep neural network with a characteristic analysis for seal stroke recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):152:1–152:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3676883.

Lv:2024:ITW

[882] Hui Lv, Hao Lv, Liu Yang, Jun Shen, La Duo, Yan Li, Qingguo Zhou, and Binbin Yong. Improved Tibetan word vectors models based on position information fusion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):153:1–153:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3681787.

Li:2024:ELR

[883] Jiajia Li, Ping Wang, Zuchao Li, Kevin Parnow, Hai Zhao, and Weiping Ding. Enhancing lyrics rewriting with weak supervision from grammatical error correction pre-training and reference knowledge fusion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):154:1–154:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3687126.

Badri:2024:AHS

[884] Nabil Badri, Ferihane Kboubi, and Anja Habacha Chaibi. Abusive and hate speech classification in Arabic text using pre-trained language models and data augmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):155:1– 155:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3679049.

Sharma:2024:MFM

[885] Richa Sharma and Arti Arya. MMHFND: Fusing modalities for multimodal multiclass Hindi fake news detection via contrastive learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):156:1–156:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3686797.

Singh:2024:LVB

[886] Shashvat Singh, Kumkum Kumari, and Ankita Vaish. Learning and vision-based approach for human fall detection and classification in naturally occurring scenes using video data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):157:1–157:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3687125.

Wang:2024:UDA

[887] Xi Wang, Ruoqing Zhao, Jing Li, and Piji Li. An unsupervised domainadaptive framework for Chinese spelling checking. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(11):158:1–158:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3689821.

Shukla:2024:SCA

[888] Shiv Shankar Prasad Shukla and Maheshwari Prasad Singh. Stacked classification approach using optimized hybrid deep learning model for early prediction of behaviour changes on social media. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):159:1–159:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3689906.

M:2024:SSM

[889] Rahul Raj M. and Dhanya S. Pankaj. Social-sum-Mal: a dataset for abstractive text summarization in Malayalam. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(11):160:1–160:??, November 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3696107.

Cao:2024:WSD

[890] Yukun Cao, Chengkun Jin, Yijia Tang, and ZiYue Wei. Word sense disambiguation combining knowledge graph and text hierarchical structure. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):161:1–161:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://

dl.acm.org/doi/10.1145/3677524.

Singh:2024:HHA

[891] Geetanjali Singh, Namita Mittal, and Satyendra Singh Chouhan. HindiSumm: a Hindi abstractive summarization benchmark dataset. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):162:1–162:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3696207.

Zhao:2024:KGG

[892] Yang Zhao, Xiaomian Kang, Yaping Zhang, Jiajun Zhang, Yu Zhou, and Chengqing Zong. Knowledge graph guided neural machine translation with dynamic reinforce-selected triples. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):163:1–163:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3696664.

Huang:2024:EDS

[893] Wen-Chieh Huang and Yu-Ling Hsueh. An emotional dialogue system using conditional generative adversarial networks with a sequence-to-sequence transformer encoder. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):164:1– 164:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3698394.

Chrismanto:2024:EBN

[894] Antonius Rachmat Chrismanto, Edi Winarko, and Yohanes Suyanto. EiAP-BC: a novel emoji aware inter-attention pair model for contextual spam comment detection based on posting text. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 23(12):165:1–165:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3696663.

Liu:2024:ICL

[895] Yijiang Liu, Fei Li, and Donghong Ji. Improving cross-lingual aspectbased sentiment analysis with sememe bridge. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):166:1–166:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3691342.

Kong:2024:RTS

[896] Chunwei Kong, Xueqiang Lv, Le Zhang, Haixing Zhao, Zangtai Cai, and Yuzhong Chen. RPEPL: Tibetan sentiment analysis based on relative position encoding and prompt learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):167:1–167:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3698575.

Prabakaran:2024:EDC

[897] Senthil Prabakaran, Navaneetha Krishnan Muthunambu, and Nagarajan Jeyaraman. Empowering digital civility with an NLP approach for detecting X (formerly known as Twitter) cyberbullying through boosted ensembles. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):168:1–168:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3695251.

Ali:2024:EPB

[898] Mubashir Ali, Anees Baqir, Hafiz Husnain Raza Sherazi, Shehzad Khalid, Phillip Smith, and Mark Lee. An extended pattern based comprehensive stemmer for the Urdu language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):169:1– 169:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3701231.

Zhang:2024:EET

[899] Guangjun Zhang, Hu Zhang, Ru Li, and Hongye Tan. EADRE: Eventtype aware dynamic representation of entities in document-level event extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):170:1–170:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3695767.

Kumar:2024:CAA

[900] Naresh Kumar, Parveen Kumar, Sushreeta Tripathy, Neelamani Samal, Debasis Gountia, Praveen Gatla, and Teekam Singh. Context-aware adversarial graph-based learning for multilingual grammatical error correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):171:1–171:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3696106.

Obiang:2024:ITR

[901] Saint Germes B. Bengono Obiang, Norbert Tsopze, Paulin Melatagia Yonta, Jean-François Bonastre, and Tania Jiménez. Improving tone recognition performance using Wav2vec 2.0-based learned representation in Yoruba, a low-resourced language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(12):172:1–172:??, December 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/10.

1145/3690384.

Singh:2025:DKM

[902] Telem Joyson Singh, Sanasam Ranbir Singh, and Priyankoo Sarmah. Distilling knowledge in machine translation of agglutinative languages with backward and morphological decoders. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1): 1:1–1:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3703455.

A:2025:SDL

[903] Madhavaraj A, Bharathi Pilar, and Ramakrishnan Angarai Ganesan. Subword dictionary learning and segmentation for expanding the vocabulary of automatic speech recognition in Tamil and Kannada. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):2:1–2:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3705312.

Bolucu:2025:SIG

[904] Necva Bölücü and Burcu Can. Semantically-informed graph neural networks for irony detection in Turkish. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):3:1– 3:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3705610.

Dang:2025:UBR

[905] Xiaochao Dang, Guozhen Ding, Xiaohui Dong, Fenfang Li, Shiwei Gao, and Yue Wang. UIE-based relational extraction task for mine hoist fault data. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):4:1–4:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3705313.

Lin:2025:SYE

[906] Nankai Lin, Meiyu Zeng, Wentao Huang, Shengyi Jiang, Lixian Xiao, and Aimin Yang. A simple yet effective corpus construction framework for Indonesian grammatical error correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1): 5:1–5:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3704264.

Acar:2025:RCS

[907] Ali Acar and Selma Tekir. Recognition of counterfactual statements in Turkish. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):6:1–6:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3706105.

Tabassam:2025:UUP

[908] Muhammad Rauf Tabassam, Hajra Waheed, Iqra Safder, Raheem Sarwar, Naif Radi Aljohani, Raheel Nawaz, Saeed-Ul Hassan, Farooq Zaman, and Ahtazaz Ahsan. UPON: Urdu poetry generation using deep learning: a novel approach and evaluation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1): 7:1–7:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-

4702 (electronic). URL https://dl.acm.org/doi/10.1145/3708535.

Wang:2025:VSG

[909] Yuanlong Wang, Qiang Ma, Ru Li, and Hu Zhang. Visual story generation model guided by multi granularity image information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):8:1–8:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3708886.

Sun:2025:SFG

[910] Teng Sun, Zhenqiu Shu, Yuxin Huang, Hongbin Wang, and Zhengtao Yu. Semantic feature graph consistency with contrastive cluster assignments for multilingual document clustering. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):9:1– 9:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3708887.

Al-Zoghby:2025:OMM

[911] Aya M. Al-Zoghby, Esraa Mohamed K. Al-Awadly, Ahmed Ismail Ebada, and Wael A. Awad. Overview of multimodal machine learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(1):10:1–10:??, January 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3701031.

Zhou:2025:GMF

[912] Dong Zhou, Qiang Ouyang, Nankai Lin, Yongmei Zhou, and Aimin Yang. GS2F: Multimodal fake news detection utilizing graph structure and guided semantic fusion. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):11:1–11:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3708536.

Boutadjine:2025:HVM

[913] Amal Boutadjine, Fouzi Harrag, and Khaled Shaalan. Human vs. machine: a comparative study on the detection of AI-generated content. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):12:1–12:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.

org/doi/10.1145/3708889.

Nasr:2025:ERS

[914] Latifa Ibn Nasr, Abir Masmoudi, and Lamia Hadrich Belguith. Emotion recognition from spontaneous Tunisian dialect speech. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):13:1–13:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3708340.

Aljeezani:2025:AAR

[915] Othman Aljeezani, Dorieh Alomari, and Irfan Ahmad. Arabic app reviews: Analysis and classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):14:1– 14:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3708987.

Sattar:2025:DER

[916] Hafsa Sattar, Mubasher Munir, Muhammad Kamran Malik, and Zara Nasar. Detecting empathy in Roman Urdu using transfer learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):15:1–15:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3711825.

Teng:2025:INM

[917] Bingtao Teng, Yuan Chen, and Juwei Zhang. Improving neural machine translation in the field of electrical engineering by using sentence backbone information. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):16:1–16:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/10.1145/3712261.

Pokhriyal:2025:SSD

[918] Himani Pokhriyal and Goonjan Jain. Supposititious sarcasm detection and sentiment analysis coping Hindi language in social networks harnessing Zipf–Mandelbrot probabilistic optimisation and perplexity entropy learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):17:1–17:??, February 2025. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3712061.

Taji:2025:PDO

[919] Mobina Taji, Arash Ghafouri, Hasan Naderi, and Behrouz Minaei-Bidgoli. PersianMHQA: a dataset for open domain Persian multi-hop question answering based on Wikipedia encyclopedia. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):18:1–18:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3711826.

Tang:2025:TTC

[920] Chao Tang, Zelin Tan, and Xiaobing Zhao. Tibetan text classification based on prompt learning and ensemble learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):19:1–19:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3711827.

Bagies:2025:CAD

[921] Taghreed Bagies, Rahaf Alsuhaimi, Miada Almasre, and Alaa Bafail. A comprehensive analysis dashboard for detecting similar Saudi Twitter accounts by using stylometric features. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(2):20:1– 20:??, February 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3705002.

Hong:2025:NMH

[922] Trung Phan Hong and Phuc Do. A novel multi-hop query answering system based on a large knowledge graph and distributed computing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):21:1–21:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3711824.

Mohan:2025:MAH

[923] Jayanth Mohan, Spandana Reddy Mekapati, Premjith B, Jyothish Lal G, and Bharathi Raja Chakravarthi. A multimodal approach for hate and offensive content detection in Tamil: From corpus creation to model development. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):22:1–22:??, March 2025. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https:

//dl.acm.org/doi/10.1145/3712260.

Alajmi:2025:AQG

[924] Anwar Alajmi, Haniah Altabaa, Sa’ed Abed, and Imtiaz Ahmad. Arabic question generation using transformers. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):23:1– 23:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3701559.

Wang:2025:SAM

[925] Haijie Wang and Jiajia Jiao. Sentiment analysis of MOOC reviews based on knowledge dependency tree. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):24:1–24:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3713073.

Sharma:2025:HSD

[926] Deepawali Sharma, Tanusree Nath, Vedika Gupta, and Vivek Kumar Singh. Hate speech detection research in South Asian languages: a survey of tasks, datasets and methods. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):25:1– 25:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702

(electronic). URL https://dl.acm.org/doi/10.1145/3711710.

Rai:2025:TSL

[927] Abigail Rai and Samarjeet Borah. Tokenization and stemming of Limbu language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):26:1–26:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3712018.

Wenbo:2025:LTC

[928] Zhang Wenbo, Dang Hongbo, Bao Zhenshan, and Song Bingyan. LAMGCN: Traditional Chinese medicine herb recommendation via LSTMs with attention mechanisms and graph convolutional networks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):27:1–27:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3708888.

Tian:2025:DPB

[929] Yu Tian, Junhui Li, Suyang Zhu, and Guodong Zhou. DialoguePFM: Prompt-based fusion model for emotion recognition in conversation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):28:1–28:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/

doi/10.1145/3714410.

Wang:2025:PPL

[930] Shiming Wang, Li-Ping Chen, Yang Ai, Yajun Hu, and Zhen-Hua Ling. PhonemeVec: a phoneme-level contextual prosody representation for speech synthesis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):29:1–29:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/10.1145/3711828.

Emiru:2025:GGE

[931] Eshete Derb Emiru and Desalegn Mamo Wendyifraw. Geéz grammar error handling using neural machine translation approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(3):30:1–30:??, March 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.acm.org/doi/

10.1145/3711829.

Chen:2025:CQE

[932] Jie Chen, Yinlong Wang, Shu Zhao, Peng Zhou, and Yanping Zhang. Contextualized quaternion embedding towards polysemy in knowledge graph for link prediction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):31:1–31:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2025:SDC

[933] Yujie Li, Hongfang Zhou, and Yinbo Xin. A DA-STCM dual-channel short text classification method based on attention mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):32:1–32:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Xu:2025:BDA

[934] Junjie Xu, Luwei Xiao, Anran Wu, Tianlong Ma, Daoguo Dong, and Liang He. Bidirectional directed acyclic graph neural network for aspectlevel sentiment classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):33:1–33:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ahmad:2025:SIC

[935] Iram Ali Ahmad, Praveen Gatla, and Rajesh Kumar Mundotiya. Sarcasm identification and classification in Hindi newspaper headlines. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):34:1–34:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Choudhury:2025:ESA

[936] Pankaj Choudhury, Prithwijit Guha, and Sukumar Nandi. Exploring semantic attributes for image caption synthesis in low-resource Assamese language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):35:1–35:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Alyousf:2025:SDS

[937] Mona Alyousf and Mohamad Firas Alhalabi. A survey of document stemming algorithms in information retrieval systems. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):36:1–36:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Song:2025:MDS

[938] Xinglong Song, Changlin Song, Haolu Yu, Yonghua Zhu, and Hong Yao. MixSong: Diverse and strictly formatted Chinese poetry generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):37:1–37:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2025:NBB

[939] Youwei Wang and Lizhou Feng. New bagging based ensemble learning algorithm distinguishing short and long texts for document classification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):38:1–38:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Connor:2025:HSR

[940] Patrick Charles Connor. A hybrid statistical and rule-based approach to extremely low-resource machine transliteration. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):39:1–39:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

K:2025:TTB

[941] Reji Rahmath K., Reghu Raj P.c, and Rafeeque P.c. TransQAM: Transformer-based question answering system in Malayalam. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(4):40:1–40:??, April 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Phan:2025:TOT

[942] Thi-Hoai Phan, Hoang-Son Bui, Tri Trung Kien Le, Thi-Ngoc-Diep Do, Thuy-Binh Nguyen, Hong-Quan Nguyen, Thanh-Hai Tran, Thanh-Thuy Pham, and Thi-Lan Le. Towards an online text-based person search in Vietnamese language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):41:1–41:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhuang:2025:TQG

[943] Yan Zhuang, Yuan Sun, Yijie Li, Sisi Liu, and Xiaobing Zhao. Tibetan question generation based on key sentence and knowledge graph. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):42:1–42:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ma:2025:DMT

[944] Xinchen Ma, Man Lan, Wenbo Hu, and Yue Lu. Dongba machine translation with transfer learning: Leveraging pre-trained ancient Chinese models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):43:1–43:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Huang:2025:LCS

[945] Yibo Huang, Li An, and Qiuyu Zhang. Low complexity speech secure hash retrieval algorithm based on KDTree nearest neighbor search. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):44:1–44:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Le:2025:LLS

[946] The Anh Le and Hai Son Le. LatVis: Large-scale task-specific language model for low-resource Vietnamese multi-document summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):45:1–45:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ahmed:2025:UWS

[947] Anil Ahmed, Degen Huang, Syed Yasser Arafat, and Khawaja Iftekhar Rashid. Urdu word sense disambiguation: Leveraging contextual stacked embedding, Siamese transformer encoder 1DCNN-BiLSTM, and gloss data augmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):46:1–46:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mazumder:2025:ICM

[948] Debajyoti Mazumder, Aakash Kumar, and Jasabanta Patro. Improving code-mixed hate detection by native sample mixing: a case study for Hindi–English code-mixed scenario. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):47:1– 47:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mortadi:2025:IIM

[949] Ahmad Mortadi, Waleed Nazih, Mohamed I. Eldesouki, and Yasser Hifny. Intelligent de-identification of medical discharge summaries using hybrid NLP techniques. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):48:1–48:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Li:2025:SAA

[950] Wei Li, Yi Li, Yanqiu Shao, and Mengxi Bi. A structure-aware approach leveraging semantic relevance graph for annotation alignment of Chinese classics. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):49:1–49:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Lora:2025:CCL

[951] Sanzana Karim Lora, M. Sohel Rahman, and Rifat Shahriyar. Con-VerSum: a contrastive learning-based approach for data-scarce solution of cross-lingual summarization beyond direct equivalents. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):50:1–50:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Ranathunga:2025:SST

[952] Surangika Ranathunga, Rumesh Sirithunga, Himashi Rathnayake, Lahiru de Silva, Thamindu Aluthwala, Saman Peramuna, and Ravi Shekhar. SiTSE: Sinhala text simplification dataset and evaluation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):51:1–51:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2025:LAT

[953] Ya Wang, Cungen Cao, Xin Huang, Ming Hui, and Wei Zheng. A linguistic approach towards automatic acquisition and classification of common Chinese affective events. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):52:1–52:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2025:FCI

[954] Xiaowen Liu, Xiaoli Zhao, Jiale Chen, and Zehui Li. FedINER: Chinese industrial named entity recognition with federated learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):53:1–53:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sulistyo:2025:PLR

[955] Danang Arbian Sulistyo, Didik Dwi Prasetya, Fadhli Almu’iini Ahda, and Aji Prasetya Wibawa. Pivoted low resource multilingual translation with NER optimization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(5):54:1–54:??, May 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Hou:2025:ABS

[956] Linlin Hou, Wenhui Tu, Ting Yu, Ting Jiang, Mohamed Bah, Zenghui Xu, Yu Zhang, Gaoming Yang, and Ji Zhang. Aspect-based sentiment analysis for COVID-19: a heterogeneous graph convolutional network approach. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):55:1–55:26, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2025:DSL

[957] Haijiang Liu, Chen Qiu, Qiyuan Li, Maofu Liu, Li Wang, and Jinguang Gu. Deep stable learning for cross-lingual dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):56:1–56:34, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Le:2025:OAG

[958] Huong Le, Ngoc Luu, Thanh Nguyen, Tuan Dao, and Sang Dinh. Optimizing answer generator in Vietnamese legal question answering systems using language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):57:1–57:17, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Boughareb:2025:ACD

[959] Djalila Boughareb, Rima Boughareb, Samir Hallaci, Mohammed Raid Elislam Boukherouba, and Hamid Seridi. Addressing the complexity of dialectal Arabic: an enhanced encoder-decoder ensemble approach for optimized sentiment analysis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):

58:1–58:21, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Dai:2025:TTT

[960] Wei Dai, Dongfang Han, Turdi Tohti, Yi Liang, Zicheng Zuo, Yuanyuan Liao, and Qingwen Yang. TFT-TL: Token-level filter training transfer learning for low-resource neural machine translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):59:1–59:19, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Sethiya:2025:ISL

[961] Nivedita Sethiya, Saanvi Nair, Puneet Walia, and Chandresh Maurya. Indic-ST: a large-scale multilingual corpus for low-resource speech-to-text translation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):60:1–60:25, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2025:OEA

[962] Lixin Liu, Aiwen Jiang, Yinuo Chen, Changhong Liu, Qi Huang, and Mingwen Wang. Open-ended autoregressive visual storytelling via parameter efficient instruction tuning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):61:1– 61:19, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Albilali:2025:AAK

[963] Eman Albilali, Nora Al-Twairesh, and Manar Hosny. AKER: Arabic knowledge-enriched reader for machine reading comprehension. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(6):62:1–62:29, June 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Minoofam:2025:DQC

[964] Seyyed Amir Hadi Minoofam, Azam Bastanfard, and Mohammad Reza Keyvanpour. DB-QM: a comparative quality measurement and its prospective on Persian/Arabic databases for OCR. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):63:1–63:25, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mehamed:2025:HAA

[965] Moges Ahmed Mehamed, Shengwu Xiong, and Awet Fesseha Aberha. Hybrid approach for automatic text summarization for low-resourced Amharic language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):64:1–64:14, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Abdulhamed:2025:ENN

[966] Ahmed Abdulhamed, Prabhat Ranjan, and Shengwu Xiong. Entity naming in NLP: Hybrid approach GPT transformer and multi-level RNN. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):65:1–65:26, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Huang:2025:ECV

[967] Yuxin Huang, Yuanlin Yang, Zhengtao Yu, Yantuan Xian, and Yan Xiang. Enhanced Chinese–Vietnamese cross-language event detection via aligned knowledge event graph. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):66:1–66:19, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Aldawsari:2025:AAE

[968] Mohammed Aldawsari and Omer Dawood. AraEventCoref: an Arabic event coreference dataset and LLM benchmarks. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):67:1–67:12, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Maitra:2025:DBS

[969] Arkadip Maitra, Shree Mitra, Siladittya Manna, Saumik Bhattacharya, and Umapada Pal. Decorrelation-based self-supervised visual representation learning for writer identification. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7): 68:1–68:17, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Trinh:2025:VKG

[970] Tam Trinh, Anh Dao, Thi Hong Nhung Hy, and Truong Son Hy. VietMedKG: Knowledge graph and benchmark for traditional Vietnamese medicine. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):69:1–69:17, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2025:IUD

[971] Yuanying Wang, Fuyong Xu, Yingzheng Zhu, Guangjin Wang, Peiyu Liu, and Ran Lu. IMDP: a unify dialogue framework with awareness and understanding for implicit personalized dialogue generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):70:1–70:21, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Yung:2025:EEP

[972] Yiu Cheong Yung, Ying-Jia Lin, and Hung-Yu Kao. Exploring the effectiveness of pre-training language models with incorporation of diglossia for Hong Kong content. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):71:1–71:16, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bawitlung:2025:MAS

[973] Andrew Bawitlung, Sandeep Kumar Dash, and Radha Mohan Pattanayak. Mizo automatic speech recognition: Leveraging Wav2vec 2.0 and XLS-R for enhanced accuracy in low-resource language processing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):72:1–72:15, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Hu:2025:TAE

[974] Weilin Hu, Wei Lin, Qizheng Li, Xiaona Yu, and Chengyan Zhu. Textile AI-enhanced translation system based on mapping probability and incontext learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(7):73:1–73:17, July 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Taylor:2025:MSA

[975] Serena Taylor and Fariza Fauzi. Multimodal sentiment analysis for the Malay language: New corpus using CNN-based framework. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):74:1–74:26, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zuhra:2025:TDN

[976] Fatima Tuz Zuhra and Khalid Saleem. Towards development of new language resource for Urdu: The large vocabulary word embeddings. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):75:1–75:14, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2025:ANI

[977] Xiangqi Wang, Yehao Kong, Luyuan Xie, Shengfang Zhai, Tairui Wang, Boyan Chen, Junkai Liang, and Xin Zhang. AdvAudio: a new information hiding method via fooling automatic speech recognition model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):76:1–76:26, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

B:2025:KRB

[978] Rashmi K B, H S Guruprasad, and Shambhavi B R. KaEnLandetector: Rule-based language annotation and transformer-based language detection for Kannada–English code-mixed text. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):77:1–77:19, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zha:2025:PER

[979] Yongfu Zha, Xuanxuan Che, Lina Sun, and Yumin Dong. Personalized explainable recommendations for self-attention collaboration. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):78:1–78:13, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Gain:2025:IVC

[980] Baban Gain, Dibyanayan Bandyopadhyay, Samrat Mukherjee, Chandranath Adak, and Asif Ekbal. Impact of visual context on noisy multimodal NMT: an empirical study for English to Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):79:1–79:27, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mahapatra:2025:MNB

[981] Sayan Mahapatra, Debtanu Datta, Shubham Soni, Adrijit Goswami, and Saptarshi Ghosh. MILPaC: a novel benchmark for evaluating translation of legal text to Indian languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):80:1–80:30, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

FathAlalim:2025:ATB

[982] Ahmed Fat’hAlalim, Yongjian Liu, Qing Xie, and Nahla Ibrahim. Advancements in transformer-based models for enhanced hate speech detection in Arabic: Addressing dialectal variations and cross-platform challenges. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):81:1–81:29, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Lata:2025:SSM

[983] Kusum Lata, Pardeep Singh, and Kamlesh Dutta. SMDDH: Singleton mention detection using deep learning in Hindi text. ACM Transactions on Asian and Low-Resource Language Information Processing (TAL-LIP), 24(8):82:1–82:24, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Chen:2025:DSL

[984] Huiyao Chen, Xin Zhang, Jing Chen, Meishan Zhang, and Min Zhang. Dependency scoring learning and corpus boosting for translation-based cross-lingual dependency parsing. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):83:1–83:18, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Dahou:2025:SDA

[985] Abdelghani Dahou, Abdelhalim Hafedh Dahou, Mohamed Amine Cheragui, Amin Abdedaiem, Mohammed A. A. Al-Qaness, Mohamed Abd Elaziz, Ahmed A. Ewees, and Zhonglong Zheng. A survey on dialect Arabic processing and analysis: Recent advances and future trends. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):84:1–84:45, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mi:2025:LIS

[986] Chenggang Mi, Shaoliang Xie, Yu Li, and Zhenghan He. Loanword identification in social media texts with extended code-switching datasets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):85:1–85:19, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mishra:2025:HMC

[987] Santosh Kumar Mishra, Soham Chakraborty, Sriparna Saha, and Pushpak Bhattacharyya. Hi-MetaCap: Configuring object relational transformer in meta-learning environment for image captioning in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):86:1–86:15, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mohammadalizadeh:2025:NBP

[988] Parman Mohammadalizadeh and Leila Safari. A novel benchmark for Persian table-to-text generation: a new dataset and baseline experiments. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):87:1–87:17, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bibi:2025:BBU

[989] Sarfraz Bibi, Sohail Asghar, and Muhammad Zubair. Breaking barriers in URDU WSD: The transfer learning enriched MAKS framework. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(8):88:1–88:12, August 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Gang:2025:CFT

[990] Liu Gang, Yang Wenli, Wang Tongli, and He Zhihao. Corpus fusion and text summarization extraction for multi-feature enhanced entity alignment. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):89:1–89:15, September 2025. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kodali:2025:HJP

[991] Prashant Kodali, Anmol Goel, Likhith Asapu, Vamshi Krishna Bonagiri, Anirudh Govil, Monojit Choudhury, Ponnurangam Kumaraguru, and Manish Shrivastava. From human judgements to predictive models: Unravelling acceptability in code-mixed sentences. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24 (9):90:1–90:31, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mughal:2025:FRD

[992] Nimra Mughal, Ghulam Mujtaba, Muhammad Hussain Mughal, Abdul Manaf, and Zainab Kamangar. Fake reviews detection on e-commerce Websites using novel user behavioral features: an experimental study. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):91:1–91:44, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhang:2025:ICI

[993] Yun Zhang, Yongguo Liu, Jiajing Zhu, Zhi Chen, Shuangqing Zhai, and Xindong Wu. Inner-character and inner-word features based representation learning for Chinese word embedding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9): 92:1–92:33, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Feng:2025:AHP

[994] Changkai Feng, Tong Xu, Shiwei Wu, Derong Xu, and Enhong Chen. Adaptive hierarchical prompt for open-vocabulary scene graph generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):93:1–93:20, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kumar:2025:TSC

[995] Senthil Kumar V and Vinoth Kumar B. Text sentiment classification using knowledge transfer-based matrix diffractive Riemannian residual triangulation topology aggregation model. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9): 94:1–94:28, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mai:2025:PSB

[996] Cheng-Cheng Mai, Yu Chen, Ziyu Gong, Hanxiang Wang, Mengchuan Qiu, Chunfeng Yuan, and Yihua Huang. PromptCNER: a segmentationbased method for few-shot Chinese NER with prompt-tuning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):95:1–95:24, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Iqbal:2025:USP

[997] Hafiz Rizwan Iqbal, Muhammad Sharjeel, Jawad Shafi, Usama Mehmood, and Agha Ali Raza. Urdu sentential paraphrased plagiarism detection using large language models. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):96:1– 96:20, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wang:2025:UPB

[998] Yu Wang, Jinmao Xu, Miao Wang, Tianrui Li, and Xuanren Qu. A unified prompt-based framework with label semantic expansion for joint multi-intent detection and slot filling. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):97:1– 97:14, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Checker:2025:HHE

[999] Saksham Checker, Madhuri Yadav, and Rahul Katarya. HiEnWrite: a Hindi–English bilingual dataset for Big Five personality detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):98:1–98:19, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Qingdaoerji:2025:MES

[1000] Ren Qingdaoerji, Yang Yang, and Wang Lele. Mongolian emotional speech synthesis based on CGAN and improved FastSpeech2. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):99:1–99:16, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Singh:2025:HCA

[1001] Sanskar Singh, Shivam Kushwaha, Avantika Singh, and Shaifu Gupta. Hinglish cross-accent model agnostic meta-learning automatic speech recognition. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):100:1–100:18, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Al-Thubaity:2025:NDA

[1002] Abdulmohsen Al-Thubaity. A novel dataset for Arabic domain specific term extraction and comparative evaluation of BERT-based models for Arabic term extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):101:1– 101:12, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Fashwan:2025:CLA

[1003] Amany Fashwan and Sameh Alansary. Computational linguistic approach to orthographic representation of Egyptian Arabic: Challenges and implications. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):102:1–102:14, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Tulajiang:2025:BLN

[1004] Paerhati Tulajiang, Yuanyuan Sun, Yuanyu Zhang, Yingying Le, Kelaiti Xiao, and Hongfei Lin. A bilingual legal NER dataset and semanticsaware cross-lingual label transfer method for low-resource languages. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(9):103:1–103:21, September 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Bouchal:2025:TAR

[1005] Hakim Bouchal, Ahror Belaid, and Farid Meziane. Towards accurate recognition of historical Arabic manuscripts: a novel dataset and a generalizable pipeline. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):104:1–104:30, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Babayigit:2025:BBM

[1006] Bilal Babayigit, Hamza Sattuf, and Mohammed Abubaker. BERT-based models for keyword extraction from Arabic scientific articles. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):105:1–105:25, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Shukla:2025:CMR

[1007] Shubham Shukla, Sushama Nagpal, and Sangeeta Sabharwal. Codemixed romanized Hindi hate speech identification: Leveraging BERT embeddings and particle swarm optimization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24 (10):106:1–106:24, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Jose:2025:OMH

[1008] Bineesh Jose and K. P. Pushpalatha. Optimized Malayalam handwritten character recognition model using a novel DSC and stacked bi-LSTM with data augmentation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):107:1–107:30, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhou:2025:MHT

[1009] Yifei Zhou and Yves Lepage. Mixup helps translation, but do the coefficients and the selection strategy influence translation quality? ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):108:1–108:20, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Wu:2025:EPM

[1010] Ning-Qian Wu and Zhen-Hua Ling. Enhanced prosody modeling and character voice controlling for audiobook speech synthesis. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):109:1–109:24, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Liu:2025:DDM

[1011] Cencen Liu, Wen Yin, Yi Xu, Qiugang Zhan, Dongyang Zhang, and Raza Ahmad. DDM4TST: Diffusion model for fine-grained text style transfer by disentangled representation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10): 110:1–110:17, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Faheem:2025:ASU

[1012] Ali Faheem, Faizad Ullah, Muhammad Sohaib Ayub, and Asim Karim. Abstractive summarization for Urdu video description generation. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):111:1–111:21, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Kboubi:2025:FGA

[1013] Ferihane Kboubi and Anja Habacha Chaibi. Fine-grained Arabic dialect identification: Investigating various approaches across multiple datasets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):112:1–112:33, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Aloraini:2025:SCZ

[1014] Abdulrahman Aloraini, Juntao Yu, Wateen Aliady, and Massimo Poesio. A survey of coreference and zeros resolution for Arabic. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):113:1–113:24, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zhao:2025:CAK

[1015] Shuo Zhao, Jianyong Duan, Li He, Hao Wang, Qing Zhang, and Jie Liu. Continuous adaptive knowledge distillation for few-shot relation extraction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):114:1–114:13, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Hashmi:2025:RLA

[1016] Ehtesham Hashmi, Hasnain Ahmad, Muhammad Tayyab Mazhar, Sule Yildirim Yayilgan, Mehtab Afzal, and Sarang Shaikh. A robust and linguistically-aware hate speech detection system for Roman Urdu. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):115:1–115:22, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Lin:2025:DAR

[1017] Luyang Lin, Jing Li, and Kam-Fai Wong. Data-augmented and retrievalaugmented context enrichment in Chinese media bias detection. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):116:1–116:24, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Guo:2025:MAG

[1018] Hai Guo, Dawei Zhu, Jingying Zhao, and Lingling Tong. Multi-attention ghost pyramid fusion network for script identification of Chinese ancient document images. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24(10):117:1–117:16, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Mishra:2025:CLI

[1019] Santosh Kumar Mishra, Sriparna Saha, and Pushpak Bhattacharyya. Continual learning for image captioning in Hindi. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 24 (10):118:1–118:19, October 2025. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).